### v1

# Direct Bias-Correction Term Estimation for Propensity Scores and Average Treatment Effect Estimation

Abstract
This study considers the estimation of the average treatment effect (ATE). For ATE estimation, we estimate the propensity score through direct bias-correction term estimation. Let $\{(X_{i},D_{i},Y_{i})\}_{i=1}^{n}$ be the observations, where $X_{i}\in{\mathbb{R}}^{K}$ denotes $K$-dimensional covariates, $D_{i}\in\{0,1\}$ denotes a binary treatment assignment indicator, and $Y_{i}\in{\mathbb{R}}$ is an outcome. In ATE estimation, the bias-correction term $h_{0}(X_{i},D_{i})\coloneqq\frac{\mathbbm{1}[D_{i}=1]}{e_{0}(X_{i})}-\frac{\mathbbm{1}[D_{i}=0]}{1-e_{0}(X_{i})}$ plays an important role, where $e_{0}(X_{i})$ is the propensity score, the probability of being assigned treatment $1$. In this study, we propose estimating $h_{0}$ (or equivalently $e_{0}$) by directly minimizing the prediction error for $h_{0}$ without knowing $h_{0}$ in advance. After showing a basic result with least squares, we present a general framework for this direct bias-correction term estimation approach from the perspective of Bregman divergence minimization, which also generalizes the Riesz regression and covariate balancing.
1 Introduction
We consider the problem of estimating the average treatment effect (ATE) in causal inference (Imbens & Rubin, 2015). Methods for estimating ATEs are typically designed to eliminate bias arising from treatment assignment and the estimation of nuisance parameters, aiming for (asymptotic) unbiasedness and efficiency.
In ATE estimation, the propensity score, the probability of treatment assignment given covariates, plays a crucial role (Rosenbaum & Rubin, 1983). For example, the inverse probability weighting (IPW) estimator, also known as the Horvitz-Thompson estimator (Horvitz & Thompson, 1952), estimates the ATE by computing a weighted average of outcomes using the inverse of the propensity score. In the one-step bias correction method (van der Vaart, 2002), a bias-correction term depending on the propensity score is added to an initial estimator.
In this study, we focus on the role of the propensity score as a bias-correction component and propose a novel method to directly estimate either the propensity score or the bias-correction term itself. The bias-correction term is fundamental to IPW-type estimators and the one-step bias correction method, and its accurate estimation can substantially improve ATE estimation. Importantly, our objective is not to estimate the propensity score, but rather the bias-correction term, in which the propensity score appears inversely and is weighted by the treatment indicator. Since estimating the propensity score itself is not the target, we hypothesize that directly estimating the bias-correction term will lead to improved performance. To this end, we propose estimating the bias-correction term by minimizing the empirical risk targeted for the true bias-correction term.
The technical challenge is that the target variable remains unobserved, even when we directly estimate the bias-correction term or the propensity score. To address this issue, we employ techniques developed in the direct density-ratio estimation (DRE) literature (Sugiyama et al., 2012). In direct DRE, the goal is to minimize the empirical risk between the true density ratio and its model, even though the true density ratio is unknown. It is known that empirical risk minimization is feasible even without knowledge of the true propensity score. Since the inverse propensity score can be viewed as a density ratio, we can extend these existing methods to our setting.
Our motivation is closely aligned with studies on covariate balancing (Imai & Ratkovic, 2013) and Riesz regression (Chernozhukov et al., 2022a), which also aim to improve ATE estimation by appropriately estimating the propensity score or the bias-correction term. Studies in covariate balancing focus on the balancing property of propensity score estimator and estimate them using the property. Chernozhukov et al. (2022a) proposes Riesz regression which represents the bias-correction term as the Riesz representer. Although the derivation process is different, we derive the objective function that is the same as Chernozhukov et al. (2022a) by using the DRE techniques. Further, we generalize our objective by using the Bregman divergence as well as DRE in Sugiyama et al. (2011). From this generalization, we connect our approach to the covariate balancing by showing the equivalence between our objective and empirical balancing (Chan et al., 2015).
1.1 ATE estimators and bias correction
We begin by formulating the problem. There are two treatments, denoted by $1$ and 0.111In some cases, only treatment $1$ is referred to as the treatment, while treatment 0 is referred to as the control. For simplicity, we refer to them as treatment $1$ and treatment 0 throughout this study. For each treatment $d\in\{1,0\}$, let $Y(d)\in{\mathbb{R}}$ denote the potential outcome under treatment $d$. The treatment assignment indicator is denoted by $D\in\{1,0\}$, and the observed outcome is given by $Y=\mathbbm{1}[D=1]Y(1)+\mathbbm{1}[D=0]Y(0)$, meaning that we observe $Y(d)$ only if the unit is actually assigned to treatment $d$. Each unit is characterized by $K$-dimensional covariates $X\in\mathcal{X}\subset{\mathbb{R}}^{K}$, where $\mathcal{X}$ denotes the covariate space. For $n$ units indexed by $1,2,\dots,n$, let $\mathcal{D}\coloneqq\{(X_{i},D_{i},Y_{i})\}_{i=1}^{n}$ denote the observed data, where each $(X_{i},D_{i},Y_{i})$ is an i.i.d. copy of $(X,D,Y)$ generated from an underlying distribution $P_{0}$. Our goal is to estimate the ATE, defined as
$\tau_{0}\coloneqq{\mathbb{E}}\big[Y(1)-Y(0)\big],$
where the expectation is taken over the distribution $P_{0}$. Note that we can also apply our method for the ATE for the treated group (ATT). For the details about ATT estimation, see Appendix C.
Let $e_{0}(X)=P_{0}(D=1\mid X)$ denote the probability of assigning treatment $1$ given covariates $X$, which is known as the propensity score. Throughout this study, we impose the following condition, commonly referred to as the overlap assumption:
Assumption 1.1.
There exists a constant $C>0$ independent of $n$ such that $C<e_{0}(x)<1-C$ for all $x\in\mathcal{X}$.
When $e_{0}(x)$ is not constant, a distributional shift arises between the observed outcomes in the treatment and control groups, denoted by $\mathcal{G}_{1}$ and $\mathcal{G}_{0}$, respectively, where $\mathcal{G}_{d}\coloneqq\{i\in\{1,2,\dots,n\}\colon D_{i}=d\}$. This shift induces bias in the sample mean, $\frac{1}{|\mathcal{G}_{d}|}\sum_{i\in\mathcal{G}_{d}}Y_{i}=\frac{1}{|\mathcal{G}_{d}|}\sum_{i\in\mathcal{G}_{d}}Y_{i}(d)$, which deviates from ${\mathbb{E}}\left[Y(d)\right]$ and thus prevents the sample mean difference, $\frac{1}{|\mathcal{G}_{1}|}\sum_{i\in\mathcal{G}_{1}}Y_{i}-\frac{1}{|\mathcal{G}_{0}|}\sum_{i\in\mathcal{G}_{0}}Y_{i}$, from being an unbiased estimator of the ATE.
To address this issue, several debiased estimators have been proposed under standard regularity conditions. In this section, we introduce two representative estimators, the inverse probability weighting (IPW) estimator and the augmented IPW (AIPW) estimator, as follows:
IPW estimator.
$\widetilde{\tau}^{\mathrm{IPW}}\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\frac{\mathbbm{1}[D_{i}=1]Y_{i}}{e_{0}(X_{i})}-\frac{\mathbbm{1}[D_{i}=0]Y_{i}}{1-e_{0}(X_{i})}\right)=\frac{1}{n}\sum^{n}_{i=1}\left(\frac{\mathbbm{1}[D_{i}=1]}{e_{0}(X_{i})}-\frac{\mathbbm{1}[D_{i}=0]}{1-e_{0}(X_{i})}\right)Y_{i}$.
AIPW estimator.
$\widetilde{\tau}^{\mathrm{AIPW}}\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\left(\frac{\mathbbm{1}[D_{i}=1]}{e_{0}(X_{i})}-\frac{\mathbbm{1}[D_{i}=0]}{1-e_{0}(X_{i})}\right)\left(Y_{i}-\mu_{0}(D_{i},X_{i})\right)+\mu_{0}(1,X_{i})-\mu_{0}(0,X_{i})\right)$, where $\mu_{0}(d,X)$ is the expected conditional outcome ${\mathbb{E}}\left[Y(d)\mid X\right]$ of treatment $d$ given $X$. The AIPW estimator is also known as the doubly robust (DR) estimator (Bang & Robins, 2005).
Bias-correction term. In both estimators, the term
$h_{0}(D,X)\coloneqq h(D,X;1/e_{0})\coloneqq\frac{\mathbbm{1}[D=1]}{e_{0}(X)}-\frac{\mathbbm{1}[D=0]}{1-e_{0}(X)}$
is crucial. This term, referred to as the bias-correction term, is central to ATE estimation (Schuler & van der Laan, 2024). A common approach is to estimate $e_{0}$ using logistic regression and then plug the resulting estimate $\widehat{e}^{{\mathrm{L}}}_{n}$ into $h$. Note that in automatic debiased machine learning, the term is also referred to as the Riesz representer (Chernozhukov et al., 2022b).
For example, in a typical one-step bias correction, we first construct an ATE estimator as $\widehat{\tau}^{\mathrm{DM}}_{n}\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\widehat{\mu}_{n}(1,X)-\widehat{\mu}_{n}(0,X)\right)$, where $\widehat{\mu}_{n}$ is an estimator of $\mu_{0}$. This estimator is known as the direct method (DM) or naive plug-in estimator. To obtain an efficient estimator, we add the bias-correction term $\frac{1}{n}\sum^{n}_{i=1}h_{0}(X_{i},D_{i})$ to the first-stage DM estimator $\widehat{\tau}^{\mathrm{DM}}_{n}$, yielding the AIPW estimator.
In this process, estimating the propensity score $e_{0}$ itself is a challenging problem, independent of ATE estimation. The well-known Vapnik principle states that “when solving a problem of interest, do not solve a more general problem as an intermediate step” (Vapnik, 1998). Following this principle, this study aims to estimate $h_{0}(D,X)$, or equivalently $e_{0}$, by directly minimizing the estimation error of $h_{0}(D,X)$. We emphasize that while $e_{0}$ is estimated in our approach, the primary estimation target is not $e_{0}$ but $h_{0}$.
1.2 Our contributions
Our first contribution is the proposal of a framework for direct bias-correction term estimation. We estimate the bias-correction term by directly minimizing the estimation error for the true bias-correction term $h_{0}$. To model the bias-correction term $h_{0}$, we define the inverse propensity score as $r_{0}(1,X)=\frac{1}{e_{0}(X)}$ and $r_{0}(0,X)=\frac{1}{1-e_{0}(X)}$. Let $\mathcal{R}$ be a model (hypothesis class) for $r_{0}$, and consider approximating $r_{0}$ using some $r\in\mathcal{R}$. We approximate $r_{0}$ by the minimizer $r^{*}\in\mathcal{R}$ of the mean squared error (MSE) ${\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}\right]$. Here, we estimate $h_{0}(D,X)=h(D,X;r_{0})$ using a model $h(D,X;r)$ that depends on $r$. We emphasize again that although we bypass the estimation of $r_{0}$, our primary estimation target is $h_{0}$, not $r_{0}$.
Since this expected squared error involves the unknown function $r_{0}$, direct optimization is infeasible. However, we show that minimizing this expected squared error with respect to $r$ is equivalent to minimizing ${\mathbb{E}}\left[-2\left(r(1,X)+r(0,X)\right)+h(D,X;r)^{2}\right]$, which does not depend on the unknown function. That is, we establish the equivalence: $r^{*}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}{\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}\right]=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}{\mathbb{E}}\left[-2\left(r(1,X)+r(0,X)\right)+h(D,X;r)^{2}\right]$. The obtained squared error can then be approximated using a sample-based empirical risk function.
Our second main contribution is the theoretical analysis of the estimator obtained via direct bias-correction term estimation. Since we estimate $r_{0}$ using empirical risk minimization, we establish bounds on the estimation error using empirical process theory. Furthermore, we present examples of ATE estimators that incorporate the bias-correction term estimated using our framework and conduct simulation studies. Using standard ATE estimation techniques, we demonstrate that our method yields a $\sqrt{n}$-consistent ATE estimator.
Our third main contribution is the generalization of our framework. From the perspective of Bregman divergence minimization, we extend our framework to provide a more general methodology for direct bias-correction term estimation. Under this framework, various estimation strategies can be incorporated to enhance bias-correction term estimation.
Our fourth contribution is the unification of the existing literature: Riesz regression and covariate balancing. If we use the squared loss function in the Bregman divergence minimization, we obtain the same loss as the one used in Riesz regression. If we use a different function in the Bregman divergence, we recover a special case of the tailored loss for covariate balancing proposed in Zhao (2019), which is also equivalent to empirical balancing (Chan et al., 2015; Hainmueller, 2012).
2 Direct bias-correction term estimation
In this study, we consider estimating $h_{0}$ by minimizing the empirical risk associated with the MSE ${\mathbb{E}}\left[\big(h_{0}(D,X)-h(D,X)\big)^{2}\right]$ over $h:\{1,0\}\times\mathcal{X}\to{\mathbb{R}}$. Here, since $h_{0}$ contains the unknown function $r_{0}$, this minimization problem is inherently intractable. However, we demonstrate that empirical risk minimization remains feasible without explicit knowledge of $r_{0}$.
2.1 Least squares
Rather than directly modeling $h_{0}$, we model either $r_{0}$ or $e_{0}$ within $h_{0}$. That is, we model $h_{0}$ via a model of $r_{0}$ as $h(D,X;r)$, while $h_{0}(D,X)$ is given as $h_{0}(D,X;r_{0})$. We then aim to estimate $r_{0}$ or $e_{0}$. Given a set $\mathcal{R}$ of functions $r\colon{\mathcal{X}}\to(1,\infty)$, we approximate the bias-correction term by solving
$r^{*}\coloneqq\operatorname*{arg\,min}_{r\in\mathcal{R}}{\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}\right].$
If $r_{0}\in\mathcal{R}$, then it follows that $r_{0}(\cdot,\cdot)=r^{*}(\cdot)$ and $h(\cdot,\cdot;r^{*})=h_{0}(\cdot,\cdot)$. This MSE is associated with the MSE between the ATE $\tau_{0}$ and some estimator using the bias-correction term. Since $e_{0}(x)\in(0,1)$, a model $r(x)$ takes a value in $(1,\infty)$.
Example. For example, for $\widehat{\tau}^{\mathrm{IPW}}(r)\coloneqq\frac{1}{n}\sum^{n}_{i=1}h(D_{i},X_{i};r)Y_{i}$, we have ${\mathbb{E}}\left[\left(\tau_{0}-\widehat{\tau}^{\mathrm{IPW}}(r)\right)^{2}\right]={\mathbb{E}}\left[\left(\tau_{0}-\widetilde{\tau}^{\mathrm{IPW}}+\widetilde{\tau}^{\mathrm{IPW}}-\widehat{\tau}^{\mathrm{IPW}}(r)\right)^{2}\right]={\mathbb{E}}\left[\left(\tau_{0}-\widetilde{\tau}^{\mathrm{IPW}}\right)^{2}\right]+{\mathbb{E}}\left[\left(\widetilde{\tau}^{\mathrm{IPW}}-\widehat{\tau}^{\mathrm{IPW}}(r)\right)^{2}\right]$, where recall that we defined $\widetilde{\tau}^{\mathrm{IPW}}$ in the Introduction. Here, we have ${\mathbb{E}}\left[\left(\widetilde{\tau}^{\mathrm{IPW}}-\widehat{\tau}^{\mathrm{IPW}}(r)\right)^{2}\right]=\frac{1}{n}{\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}{\mathbb{E}}\left[Y^{2}\mid D,X\right]\right]$. Thus, the MSE ${\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}\right]$ is closely connected to the MSE ${\mathbb{E}}\left[\left(\tau_{0}-\widehat{\tau}^{\mathrm{IPW}}(r)\right)^{2}\right]$.
Direct estimation without $r_{0}$. The essential problem in this approach is that the target $h_{0}$ or $r_{0}$ is unknown; therefore, this optimization looks infeasible. However, even if we do not know $h_{0}$, we can reformulate the optimization problem into an equivalent form that does not involve $h_{0}$ as
$r^{*}=\operatorname*{arg\,min}_{r\in\mathcal{R}}{\mathbb{E}}\Big[-2r(1,X)-2r(0,X)+\mathbbm{1}[D=1]r(1,X)^{2}+\mathbbm{1}[D=0]r(0,X)^{2}\Big].$
This reformulation is derived as follows: $\min_{r\in\mathcal{R}}{\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}\right]=\min_{r\in\mathcal{R}}{\mathbb{E}}\Big[h(D,X;r_{0})^{2}-2h(D,X;r_{0})h(D,X;r)+h(D,X;r)^{2}\Big]=\min_{r\in\mathcal{R}}{\mathbb{E}}\left[-2h(D,X;r_{0})h(D,X;r)+h(D,X;r)^{2}\right]=\min_{r\in\mathcal{R}}{\mathbb{E}}\left[-2\left(r(1,X)+r(0,X)\right)+h(D,X;r)^{2}\right]$. From the second to the third line, we omit the term $h_{0}(D,X)^{2}$ since it does not affect the optimization. From the third to the fourth line, we use the following for the first term: ${\mathbb{E}}\left[h(D,X;r_{0})h(D,X)\right]={\mathbb{E}}\Bigg[\left(\frac{\mathbbm{1}[D=1]}{e_{0}(X)}-\frac{\mathbbm{1}[D=0]}{1-e_{0}(X)}\right)\Big(\mathbbm{1}[D=1]r(1,X)-\mathbbm{1}[D=0]r(0,X)\Big)\Bigg]={\mathbb{E}}\left[\frac{\mathbbm{1}[D=1]}{e_{0}(X)}r(1,X)+\frac{\mathbbm{1}[D=0]}{1-e_{0}(X)}r(0,X)\right]={\mathbb{E}}\left[\frac{e_{0}(X)}{e_{0}(X)}r(1,X)+\frac{1-e_{0}(X)}{1-e_{0}(X)}r(0,X)\right]={\mathbb{E}}\left[r(1,X)+r(0,X)\right]$.
Thus, surprisingly, we demonstrate that the least squares estimate for the unknown true bias-correction term $h_{0}$ can be defined by an objective function that does not explicitly include $h_{0}$ (or the unknown $r_{0}$) itself. As discussed in the following subsection, this objective function can be easily approximated using observations.
2.2 Empirical risk minimization
We then estimate $r_{0}$ by solving the following empirical risk minimization:
$\widehat{r}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\widehat{{\mathcal{L}}}_{n}(r)+\lambda J(r),$
where the empirical risk is given as
$\widehat{{\mathcal{L}}}_{n}(r)\coloneqq\frac{1}{n}\sum^{n}_{i=1}\Big(-2r(1,X_{i})-2r(0,X_{i})+\mathbbm{1}[D_{i}=1]r(1,X_{i})^{2}+\mathbbm{1}[D_{i}=0]r(0,X_{i})^{2}\Big),$
and $J\colon{\mathcal{R}}\to{\mathbb{R}}^{+}$ is a regularization term with a penalty coefficient $\lambda>0$.
This type of estimation method is referred to as least-squares importance fitting (LSIF, Kanamori et al., 2009) in the literature on density-ratio estimation (DRE).
3 Estimation error analysis
This section provides an estimation error analysis for $r_{0}$ estimated by the direct bias-correction term estimation method. For simplicity, throughout the analysis, we use the following model for $e_{0}$: $e(x)=\frac{1}{1+\exp(-f(x))}$, where $f\colon{\mathcal{X}}\to{\mathbb{R}}$ is a function belonging to a set $\mathcal{F}$}. We use various models for $\mathcal{F}$}, including linear models, RKHS, and neural networks. Then, the model of $r_{0}$ is given as $r(1,x)=1+\exp(-f(x)),\quad r(0,x)=\exp(f(x))$. Let $f_{0}$ be the true function for $e_{0}$; that is, $e_{0}(X)=\frac{1}{1+\exp(-f_{0}(x))}$.
3.1 Linear models
First, we consider the case where $f_{0}$ belongs to a linear model. There exists $\theta_{0}\in\Theta^{K}$ such that $e_{0}(x)=\frac{1}{1+\exp\big(-x^{\top}\theta_{0}\big)}$, where $\Theta$ is a compact parameter space. Under this assumption, we have $r_{0}(1,X)=1+\exp\big(-x^{\top}\theta_{0}\big)$ and $r_{0}(0,X)=1+\exp\big(x^{\top}\theta_{0}\big)$. Let us use $r_{\theta}(1,X)=1+\exp\big(-x^{\top}\theta\big)$ and $r_{\theta}(0,X)=1+\exp\big(x^{\top}\theta\big)$ as a model for $r_{0}$. Then, we define the estimator as $\widehat{r}^{\mathrm{Lin}}_{n}=r_{\widehat{\theta}_{n}}$, where $\widehat{\theta}_{n}\coloneqq\operatorname*{arg\,min}_{\theta\in\Theta}\widehat{{\mathcal{L}}}_{n}(r_{\theta})$. Here, we set $\lambda=0$.
We show the asymptotic normality of this estimator.
Assumption 3.1.
We assume the following: (i) $\theta_{0}$ is in the interior of $\Theta$; (ii) $\frac{\partial}{\partial\theta}\widehat{{\mathcal{L}}}_{n}(r_{\theta})$ is twice continuously differentiable on some neighborhood $\mathcal{M}$} of $\theta_{0}$ (with probability one); (iii) $\sqrt{n}\frac{\partial}{\partial\theta}\widehat{{\mathcal{L}}}_{n}(r_{\theta})\xrightarrow{d}{\mathcal{N}}(0,\Sigma_{0})$; (iv) for any sequence $\tilde{\theta}_{n}\xrightarrow{p}\theta_{0}$, it holds that $\frac{\partial^{2}}{\partial\theta\partial\theta^{T}}\widehat{{\mathcal{L}}}_{n}(r_{\theta})-B_{0}\xrightarrow{p}0$, for some non-stochastic $K\times K$ matrix $B_{0}$ that is nonsingular.
Then, from the asymptotic theory of extremum estimators, we obtain the following result:
Lemma 3.2.
Suppose that Assumptions 1.1 and 3.1 hold. Then, it holds that $\sqrt{n}\left(\widehat{\theta}_{n}-\theta_{0}\right)\xrightarrow{{\mathrm{d}}}{\mathcal{N}}(0,B_{0}^{-1}\Sigma_{0}B_{0}^{-1})$.
From the Taylor expansion of $r_{\theta}(x)$ around $\theta=\theta_{0}$, Lemma 3.2 immediately yields the following result:
Theorem $3.$
Suppose that Assumptions 1.1 and 3.1 hold. Then, it holds that $\sqrt{n}\left(\widehat{r}^{\mathrm{Lin}}_{n}(x)-r_{0}(x)\right)\xrightarrow{{\mathrm{d}}}{\mathcal{N}}(0,\kappa(x))$, where $\kappa(x)\coloneqq\left(\frac{\partial}{\partial\theta}r_{\theta_{0}}(x)\right)^{\top}B_{0}^{-1}\Sigma_{0}B_{0}^{-1}\left(\frac{\partial}{\partial\theta}r_{\theta_{0}}(x)\right)$.
3.2 RKHS
Next, we investigate the case with RKHS regression. Let ${\mathcal{R}}^{\mathrm{RKHS}}$ be a class of RKHS functions, and define the estimator as $\widehat{r}^{\mathrm{RKHS}}_{n}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}^{\mathrm{RKHS}}}\widehat{{\mathcal{L}}}_{n}(r)+\lambda\|r\|^{2}_{{\mathcal{H}}}$, where $\|\cdot\|^{2}_{{\mathcal{H}}}$ is the RKHS norm. We analyze the estimation error by employing the results in Kanamori et al. (2012), which study RKHS-based LSIF in DRE.
We define the following localized class of RKHS functions as a technical device: ${\mathcal{R}}^{\text{RKHS}}_{M}\coloneqq\big\{r\in{\mathcal{R}}^{\text{RKHS}}\colon I(r)\leq M\big\}$. We then make the following assumption using this localized class.
Assumption 3.4.
There exist constants $0<\gamma<2$, $0\leq\beta\leq 1$, $c_{0}>0$, and $A>0$ such that for all $M\geq 1$, it holds that $H_{B}(\delta,{\mathcal{R}}^{\text{RKHS}}_{M},P_{0})\leq A\left(\frac{M}{\delta}\right)^{\gamma}$, where $H_{B}(\delta,{\mathcal{R}}^{\text{RKHS}}_{M},P_{0})$ is the bracketing entropy with radius $\delta>0$ for the function class ${\mathcal{R}}^{\text{RKHS}}_{M}$ and the distribution $P_{0}$.
For the details of the definition of the bracketing entropy, see Appendix F and Definition 2.2 in van de Geer (2000).
Under these preparations, we establish an estimation error bound.
Theorem 3.5 ($L_{2}$-norm estimation error bound).
Suppose that Assumptions 1.1 and 3.4 hold. Set the regularization parameter $\lambda=\lambda_{n}$ so that $\lim_{n\to\infty}\lambda_{n}=0$ and $\lambda^{-1}_{n}=O(n^{1-\delta})$ ($n\to\infty$). If $r_{0}\in{\mathcal{R}}^{\mathrm{RKHS}}$, then we have $\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}=\left\|e_{0}(X)\left(\widehat{r}^{\mathrm{RKHS}}_{n}(1,X)-r_{0}(1,X)\right)\right\|^{2}_{P_{0}}+\left\|\big(1-e_{0}(X)\big)\left(\widehat{r}^{\mathrm{RKHS}}_{n}(0,X)-r_{0}(0,X)\right)\right\|^{2}_{P_{0}}=O_{P_{0}}\left(\lambda^{1/2}\right)$.
The proof is provided in Appendix F, following the approach of Kanamori et al. (2012). The parameter $\gamma$ is determined by the function class to which $f_{0}$ belongs.
3.3 Neural networks
This section provides an estimation error analysis when we use neural networks for $\mathcal{R}$}. Our analysis is mostly based on Kato & Teshima (2021) and Zheng et al. (2022).
Feedforward neural networks (FNNs). We define FNNs as follows:
Definition 3.6 (Feedforward neural networks. From Zheng et al. (2022)).
Let $\mathcal{D}$}, $\mathcal{W}$}, $\mathcal{U}$}, and ${\mathcal{S}}\in(0,\infty)$ be parameters that can depend on $n$. Let ${\mathcal{F}}^{\mathrm{FNN}}\coloneqq{\mathcal{F}}^{\mathrm{FNN}}_{M,{\mathcal{D}},{\mathcal{W}},{\mathcal{U}},{\mathcal{S}}}$ be a class of ReLU-activated FNNs $r_{\theta}(1,x)=1/e_{\theta}(x)$ and $r_{\theta}(0,x)=1/(1-e_{\theta}(x))$, where $e_{\theta}\colon{\mathbb{R}}^{K}\to{\mathbb{R}}$ with parameter $\theta$, depth $\mathcal{D}$}, width $\mathcal{W}$}, size $\mathcal{S}$}, number of neurons $\mathcal{U}$}, satisfies the following conditions: (i) the number of hidden layers is $\mathcal{D}$}; (ii) the maximum width of the hidden layers is $\mathcal{W}$}; (iii) the number of neurons in $e_{\theta}$ is $\mathcal{U}$}; (iv) the total number of parameters in $e_{\theta}$ is $\mathcal{S}$}.
Let $\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})$ be the pseudo-dimension of ${\mathcal{F}}^{\mathrm{FNN}}$. For the definition, see Anthony & Bartlett (1999) and Definition 3 in Zheng et al. (2022).
We model $r_{0}$ by $r(1,X)=\frac{1}{e(x)}$ and $r(0,X)=\frac{1}{1-e(x)}$, where $e(x)=\frac{1}{1+\exp(-f(x))}$ for $f\in{\mathcal{F}}^{\mathrm{FNN}}$. Let us denote such an estimator by $r_{f}$ for $f\in{\mathcal{F}}^{\mathrm{FNN}}$. Let $f_{0}$ denote the true function such that $r_{f_{0}}=1+\exp(-f_{0}(x))$ holds. Then, the estimator is denoted as $\widehat{f}^{\mathrm{FNN}}_{n}\coloneqq\operatorname*{arg\,min}_{f\in{\mathcal{F}}^{\mathrm{FNN}}}\widehat{{\mathcal{L}}}_{n}(r_{f})$.
Estimation error analysis. For the estimator, we can prove an estimation error bound. Let us make the following assumption.
Assumption 3.7.
There exists a constant $0<M<\infty$ such that $\|f_{0}\|_{\infty}<M$, and $\|f\|_{\infty}\leq M$ for any $f\in{\mathcal{F}}^{\mathrm{FNN}}$.
Then, we prove the following estimation error bound:
Theorem 3.8 (Estimation error bound for neural networks).
Suppose that Assumption 3.7 holds. Also assume $r_{0}\in\Sigma(\beta,M,[0,1]^{d})$ with $\beta=k+a$, where $k\in{\mathbb{N}}^{+}$ and $a\in(0,1]$, and ${\mathcal{F}}^{\mathrm{FNN}}$ has width $\mathcal{W}$} and depth $\mathcal{D}$} such that ${\mathcal{W}}=38\big(\lfloor\beta\rfloor+1\big)^{2}d^{\lfloor\beta\rfloor+1}$ and ${\mathcal{D}}=21\big(\lfloor\beta\rfloor+1\big)^{2}\lceil n^{\frac{d}{2(d+2\beta)}}\log_{2}\left(8n^{\frac{d}{2(d+2\beta)}}\right)\rceil$. Then, for $M\geq 1$ and $n\leq\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})$, it holds that $\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}=\left\|e_{0}(X)\left(\widehat{r}^{\mathrm{FNN}}_{n}(1,X)-r_{0}(1,X)\right)\right\|^{2}_{P_{0}}+\left\|(1-e_{0}(X))\left(\widehat{r}^{\mathrm{FNN}}_{n}(0,X)-r_{0}(0,X)\right)\right\|^{2}_{P_{0}}=C_{0}\big(\lfloor\beta\rfloor+1\big)^{9}d^{2\lfloor\beta\rfloor+(\beta\land 3)}n^{-\frac{2\beta}{d+2\beta}}\log^{3}n$, where $C_{0}>0$ is a constant independent of $n$.
The proof is provided in Appendix G, following the approach of Zheng et al. (2022). This result directly implies the minimax optimality of the proposed method when $f_{0}$ belongs to a Hölder class.
4 Example about the AIPW estimator
This section introduces the AIPW estimator with nuisance parameters estimated using our proposed direct bias-correction term estimation. We prove that under certain conditions, the proposed estimator is asymptotically normal. Note that this result is well known in the literature except for the use of nuisance parameters estimated via our direct bias-correction term estimation. The purpose of this section is not to provide novel methodological or theoretical results but to present an application of our proposed method.
We analyze the AIPW estimator with an estimated propensity score. Recall that the AIPW estimator is defined as $\widetilde{\tau}^{\mathrm{AIPW}}_{n}=\frac{1}{n}\sum^{n}_{i=1}\Big(h(X_{i},D_{i};\widehat{r}_{n})\left(Y_{i}-\widehat{\mu}_{n}(D_{i},X_{i})\right)+\widehat{\mu}_{n}(1,X_{i})-\widehat{\mu}_{n}(0,X_{i})\Big)$, which is also called the DR estimator.
We first make the following assumption.
Assumption 4.1 (Donsker condition).
The hypothesis classes $\mathcal{R}$} and $\mathcal{M}$} belong to the Donsker class.
For example, this assumption holds when the bracketing entropy is finite. In contrast, it is violated in high-dimensional regression or series regression settings where the model complexity diverges as $n\to\infty$. For neural networks, the assumption holds if both the number of layers and the width are finite. However, if these quantities grow with the sample size, the assumption is no longer valid.
Even if this assumption does not hold, we can still establish asymptotic normality by employing sample splitting (Klaassen, 1987). There are various ways to implement sample splitting, and one of the most well-known is cross-fitting, used in double machine learning (DML Chernozhukov et al., 2018, ,). In DML, the dataset is split into several folds, and the nuisance parameters are estimated using only a subset of the folds. This ensures that in $h(X_{i},D_{i};\widehat{r}_{n})\left(Y_{i}-\widehat{\mu}_{n}(D_{i},X_{i})\right)+\widehat{\mu}_{n}(1,X_{i})-\widehat{\mu}_{n}(0,X_{i})$, the observations $(X_{i},D_{i},Y_{i})$ are not used to construct $\widehat{\mu}_{n}$ and $\widehat{r}_{n}$. For more details, see Chernozhukov et al. (2018).
Assumption 4.2 (Convergence rate).
$\big\|\widehat{r}-r_{0}\big\|_{2}=o_{p}(1)$, $\big\|\widehat{\mu}-\mu_{0}\big\|_{2}=o_{p}(1)$, and $\big\|\widehat{r}-r_{0}\big\|_{2}\big\|\widehat{\mu}-\mu_{0}\big\|_{2}=o_{p}(1/\sqrt{n})$.
Under these assumptions, we show the asymptotic normality of $\widetilde{\tau}^{\mathrm{AIPW}}_{n}$. We omit the proof. For details, see Schuler & van der Laan (2024), for example.
Theorem 4.3 (Asymptotic normality).
Suppose that Assumptions 1.1, and 4.14.2 hold. Then, the AIPW estimator converges in distribution to a normal distribution as $\sqrt{n}\Big(\widetilde{\tau}^{\mathrm{AIPW}}_{n}-\tau_{0}\Big)\xrightarrow{{\mathrm{d}}}{\mathcal{N}}(0,V^{*})$, where $V^{*}$ is the efficiency bound defined as $V^{*}\coloneqq{\mathbb{E}}\left[\frac{\sigma^{2}(1,X)}{e_{0}(X)}+\frac{\sigma^{2}(0,X)}{1-e_{0}(X)}+\big(\tau_{0}(X)-\tau_{0}\big)^{2}\right]$ and $\tau_{0}(X)\coloneqq{\mathbb{E}}[Y(1)-Y(0)\mid X]$.
Here, $V^{*}$ is the efficiency bound given by the variance of the efficient influence function; that is, $V^{*}={\mathbb{E}}\left[\Psi^{*}(X,D,Y)^{2}\right]$ holds (van der Vaart, 1998). Thus, this estimator is efficient.
5 Generalization via the Bregman divergence minimization
We can further generalize our direct bias-correction term estimation from the viewpoint of Bregman divergence minimization and point out the connection to Riesz regression and covariate balancing.
5.1 Riesz regression and covariate balancing
Our resulting objective function is the same as the one used in Riesz regression (Chernozhukov et al., 2022a, 2024). Although the motivation and derivation differ, we point out that Riesz regression can be interpreted as a specific instance of DRE.
Building on this perspective, we generalize the objective function using Bregman divergence, following Sugiyama et al. (2011). Bregman divergence is a measure of discrepancy between two points, defined in terms of a strictly convex function (Bregman, 1967). In the context of DRE, Sugiyama et al. (2011) demonstrates that various existing methods can be formulated as Bregman divergence minimization problems. Inspired by this idea, we also extend the direct estimation method for the bias-correction term within the Bregman divergence minimization framework.
This generalization further allows us to derive the empirical balancing method proposed in Chan et al. (2015), which aims to achieve covariate balance and known as an instance of tailored loss function (Zhao, 2019). These results suggest that Riesz regression and covariate balancing can be unified under the DRE framework.
5.2 Bregman divergence minimization
Let $\mathcal{G}$} be a set of functions $g\colon{\mathbb{R}}\to{\mathbb{R}}$ that is differentiable and strictly convex. Given $d\in\{1,0\}$, we define the Bregman divergence between $r(d,\cdot),r(d,\cdot)\colon{\mathcal{X}}\to(1,\infty)$ as
$\mathrm{br}^{\dagger}_{g}\big(r_{0}(d,x)\mid r(d,x)\big)\coloneqq g(r_{0}(d,x))-g(r(d,x))-\partial g(r(d,x))(r_{0}(d,x)-r(d,x)),$
where $\partial g$ denotes the derivative of $g$. Then, we define the average Bregman divergence as
$\mathrm{BR}^{\dagger}_{g}\big(r_{0}\mid r\big)\coloneqq\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(g(r_{0}(d,X))-g(r(d,X))-\partial g(r(d,x))\Big(r_{0}(d,X)-r(d,X)\Big)\right)\Big].$
We estimate $r_{0}$ by $r^{*}=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}^{\dagger}_{g}\big(r_{0}\mid r\big)$. By dropping the term that is irrelevant to learning, we have
$r^{*}=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}_{g}\big(r\big),$
where $\mathrm{BR}_{g}\big(r\big)\coloneqq\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\Big(-g(r(d,X))+\partial g(r(d,X))r(d,X)\Big)-\partial g(r(d,X))\Big]$. For the derivation, see Appendix B.
Then, we estimate the bias-correction term $h_{0}$ (or $r_{0}$) by minimizing an empirical Bregman divergence as
$\widehat{r}_{n}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\widehat{\mathrm{BR}}_{g}\big(r\big)+\lambda J(r),$
where $\widehat{\mathrm{BR}}_{g}(r)\coloneqq\sum_{d\in\{1,0\}}\frac{1}{n}\sum^{n}_{i=1}\left(\mathbbm{1}[D_{i}=d]\Big(-g(r(d,X_{i}))+\partial g(r(d,X_{i}))r(d,X_{i})\Big)-\partial g(r(d,X_{i}))\right).$
5.3 Least squares
Our least squares method for direct bias-correction term estimation can be obtained by using $g(r)=r^{2}$. Under this choice of $g$, we obtain $\mathrm{BR}_{g}\big(r\big)=\sum_{d\in\{1,0\}}{\mathbb{E}}\left[-2r(d,X)+\mathbbm{1}[D=d]r(d,X)^{2}\right]$. This gives an objective that is the same as the one used in Chernozhukov et al. (2022a).
5.4 Constrained maximum likelihood estimation
By using different choices of $g$, we can derive various objective functions for direct bias-correction term estimation. As a specific case of Bregman divergence minimization, we introduce constrained maximum likelihood estimation.
Consider $g^{\mathrm{L}}(r)=r\log r-r$, which is a convex function. By substituting this function into the Bregman divergence, we obtain
$r^{*}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}_{g^{\mathrm{L}}}\big(r\big).$ (1)
where $\mathrm{BR}_{g^{\mathrm{L}}}\big(r\big)\coloneqq{\mathbb{E}}\big[-\log(r(1,X))-\log(r(0,X))+\mathbbm{1}[D_{i}=1]r(1,X_{i})+\mathbbm{1}[D_{i}=0]r(0,X_{i})\big]$.
Then, we estimate the bias-correction term as $\widehat{r}_{n}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\widehat{\mathrm{BR}}_{g^{\mathrm{L}}}\big(r\big)+\lambda J(r)$, where $\widehat{\mathrm{BR}}_{g^{\mathrm{L}}}(r)=\frac{1}{n}\sum^{n}_{i=1}\big(-\log\big(r(1,X_{i})\big)-\log\big(r(0,X_{i})\big)+\mathbbm{1}[D_{i}=1]r(1,X_{i})+\mathbbm{1}[D_{i}=0]r(0,X_{i})\big)$. This estimation method corresponds to unnormalized Kullback–Leibler (UKL) minimization in DRE (Sugiyama et al., 2011), which generalizes the KL importance estimation procedure (KLIEP).
Note that solving (1) is equal to
$r^{*}=\operatorname*{arg\,max}_{r\in{\mathcal{R}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\left[\log r(d,X)\right]\quad\mathrm{s.t.}~~~{\mathbb{E}}\big[\mathbbm{1}[D=1]r(1,X_{i})\big]={\mathbb{E}}\big[\mathbbm{1}[D=0]r(0,X_{i})\big]=1.$
This technique is known as Silverman’s trick (Silverman, 1982). For details, see Theorem 3.3 in Kato et al. (2023). We can replace the expected values with the sample means and define the estimation problem as $\widehat{r}_{n}=\operatorname*{arg\,max}_{r\in{\mathcal{R}}}\frac{1}{n}\sum^{n}_{i=1}\sum_{d\in\{1,0\}}\log r(d,X_{i})\quad\mathrm{s.t.}~~~\frac{1}{n}\sum^{n}_{i=1}\mathbbm{1}[D_{i}=1]r(1,X_{i})=\frac{1}{n}\sum^{n}_{i=1}\mathbbm{1}[D_{i}=0]r(0,X_{i})=1$.
5.5 Empirical balancing
Next, we derive empirical balancing as a special case of Bregman divergence minimization (Chan et al., 2015). Empirical balancing is known as a specific form of covariate balancing and can be derived from a tailored loss function (Zhao, 2019).
Let us consider $g^{\mathrm{E}}(r)=(r-1)\log\left(r-1\right)-r$, which is also convex for $r\in(1,\infty)$ and its derivative is $\partial g^{\mathrm{E}}(r)=\log\left(r-1\right)$. By substituting this function, we obtain $\mathrm{BR}_{g^{\mathrm{E}}}\big(r\big)\coloneqq\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\Big(\log\left(r-1\right)+r(d,X)\Big)-\log\left(r-1\right)\Big]$. Note that it holds that $\mathrm{BR}_{g^{\mathrm{E}}}\big(r\big)=$
${\mathbb{E}}\Big[-\mathbbm{1}[D=0]\log\left(r(1,X)-1\right)-\mathbbm{1}[D=1]\log\left(r(0,X)-1\right)+\mathbbm{1}[D=1]r(1,X)+\mathbbm{1}[D=0]r(0,X)\Big].$
Then, we estimate the bias-correction term as $\widehat{r}_{n}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\widehat{\mathrm{BR}}_{g^{\mathrm{E}}}\big(r\big)$, where the empirical Bregman divergence becomes $\widehat{\mathrm{BR}}_{g^{\mathrm{E}}}(r)=\frac{1}{n}\sum^{n}_{i=1}\big(\mathbbm{1}[D_{i}=0]\log\big(r(1,X_{i})-1\big)+\mathbbm{1}[D_{i}=1]\log\big(r(0,X_{i})-1\big)+\mathbbm{1}[D_{i}=1]r(1,X_{i})+\mathbbm{1}[D_{i}=0]r(0,X_{i})\big)$. If we model $r$ as $r(1,x)=1/e(x)$ and r(0,x)=1/(1e(x))r(0,x)=1/(1-e(x)), we have r(1,x)1=1/(r(0,x)1)r(1,x)-1=1/(r(0,x)-1). Therefore, we have $\widehat{\mathrm{BR}}_{g^{\mathrm{E}}}(r)=$
$\frac{1}{n}\sum^{n}_{i=1}\left(\mathbbm{1}[D_{i}=1]\left(-\log\left(\frac{1}{r(1,X_{i})-1}\right)+r(1,X_{i})\right)+\mathbbm{1}[D_{i}=0]\left(-\log\left(\frac{1}{r(0,X_{i})-1}\right)+r(0,X_{i})\right)\right).$
This objective function is equivalent to the one with the tailored loss proposed in Zhao (2019). From this objective, we can also derive empirical balancing (Chan et al., 2015).
6 Discussion and related work
6.1 Related work
The estimation of the bias-correction term or the propensity score has been a core interest in causal inference. The bias-correction term can be interpreted as a gradient in a one-step estimator (van der Vaart, 2002). This idea was refined by van der Laan (2006) as targeted maximum likelihood estimation (TMLE). Chernozhukov et al. (2024) also explore a related topic from the viewpoint of Riesz representers and double machine learning (DML) (Chernozhukov et al., 2018). They propose automatic DML along with its implementation (Chernozhukov et al., 2022a).
This topic has been addressed from various perspectives. One such approach focuses on estimating the propensity score by matching the distributions of the treatment and control groups (Chan et al., 2015; Deville & Särndal, 1992; Graham et al., 2012; Bryan S. Graham & Egel, 2016; Hellerstein & Imbens, 1999). For example, Imai & Ratkovic (2013) introduces the Covariate Balancing Propensity Score (CBPS), which estimates propensity scores by explicitly balancing covariate means. In parallel with, or subsequent to, Imai & Ratkovic (2013), several methods related to covariate balancing have been proposed (Hainmueller, 2012; Zubizarreta, 2015; Athey et al., 2018).
From a methodological perspective, our study is inspired by DRE (Sugiyama et al., 2012). We review the literature of DRE in Appendix A.
Table 1: Experimental results. We report the empirical MSE and Bias of each method.
Data Dimension DM DBC (LS) DBC (UKL) Logistic CBPS RieszNet DM
Three-layer perceptron Dragonnet Linear model
IPW DR IPW DR IPW DR IPW DR IPW DM DR
Model 1 $K=3$ MSE 0.006 0.392 0.005 0.374 0.005 0.330 0.004 1.429 0.006 0.017 0.021 0.040 2.781
$K=3$ Bias -0.037 -0.299 -0.024 -0.316 -0.023 -0.257 -0.022 -0.747 -0.037 -0.027 -0.025 -0.053 -0.197
$K=3$ MSE 0.521 1.956 0.481 2.779 0.478 6.510 0.507 3.570 0.515 0.464 0.510 0.379 7.511
$K=10$ Bias 0.094 -0.930 0.086 -0.822 0.088 -0.268 0.091 -1.422 0.089 -0.093 -0.106 -0.017 0.101
Model 2 $K=3$ MSE 0.048 0.343 0.033 0.819 0.037 2.838 0.045 1.848 0.044 0.030 0.034 0.051 2.866
Bias $K=3$ -0.009 -0.275 -0.011 -0.382 -0.010 -0.403 -0.011 -0.781 -0.012 -0.022 -0.020 -0.057 -0.214
$K=3$ MSE 0.517 2.006 0.474 2.980 0.477 6.517 0.507 3.816 0.512 0.407 0.446 0.424 7.482
$K=10$ Bias 0.085 -0.944 0.082 -0.823 0.085 -0.269 0.089 -1.410 0.084 -0.087 -0.096 -0.012 0.093
6.2 Comparison with the standard DRE approaches
If we follow the standard DRE approach, we may formulate the problem as the direct estimation of $r_{0}(1,X)$. For example, when using LSIF, the risk is given by ${\mathbb{E}}\big[-2r(1,X)\big]+{\mathbb{E}}\big[\mathbbm{1}[D=1]r(1,X)^{2}\big]$, which corresponds to a part of our risk: ${\mathbb{E}}\Big[-2r(1,X)-2r(0,X)+\mathbbm{1}[D=1]r(1,X)^{2}+\mathbbm{1}[D=0]r(0,X)^{2}\Big]$. Thus, our proposed method is closely connected to LSIF. However, the standard DRE approach does not address whether it is suitable for bias-correction term estimation. In fact, we can estimate $r_{0}$ by minimizing the LSIF risk, but our proposed method adopts a different risk: the sum of ${\mathbb{E}}\big[-2r(1,X)\big]+{\mathbb{E}}\big[\mathbbm{1}[D=1]r(1,X)^{2}\big]$ and ${\mathbb{E}}\big[-2r(0,X)\big]+{\mathbb{E}}\big[\mathbbm{1}[D=0]r(0,X)^{2}\big]$, which is directly related to the bias-correction term.
7 Simulation studies
We assess the empirical performance of our method through simulation studies, evaluating ATE estimation error across a range of scenarios. We compare our approach with CBPS (Imai & Ratkovic, 2013) and RieszNet (Chernozhukov et al., 2022a). Because our least squares is equivalent to Resz regression, we include RieszNet primarily as a numerical check of equivalence, noting architectural differences.
We consider two different dimensions for $X$, setting $K=3$ and $K=10$, and two different outcome models. This results in a total of four experimental settings. In all cases, the true ATE is fixed at $\tau_{0}=5.0$. To generate synthetic data, we first sample covariates $X_{i}$ from a multivariate normal distribution $\mathcal{N}(0,I_{K})$, where $I_{K}$ denotes the $K\times K$ identity matrix. The propensity score is then defined as $e_{0}(X_{i})=\frac{1}{1+\exp\bigl(-h(X_{i})\bigr)}$, where $h(X_{i})=\sum_{j=1}^{3}\alpha_{j}X_{i,j}+\sum_{j=1}^{3}\beta_{j}X_{i,j}^{2}+\gamma_{1}X_{i,1}X_{i,2}+\gamma_{2}X_{i,2}X_{i,3}+\gamma_{3}X_{i,1}X_{i,3}$. The coefficients $\alpha_{j}$, $\beta_{j}$, and $\gamma_{j}$ are independently drawn from $\mathcal{N}(0,0.5)$. Given these propensity scores, the treatment assignment $D$ is sampled accordingly. The outcome is then generated under two models, referred to as Model 1 and Model 2. In Model 1, we specify $Y_{i}=\left(X^{\top}_{i}\beta\right)^{2}+1.1+\tau_{0}D_{i}+\varepsilon_{i}$, where $\varepsilon_{i}\sim\mathcal{N}(0,1)$ and $\tau_{0}=5.0$. In Model 2, the outcome is generated as $Y_{i}=X^{\top}_{i}\beta+\left(X^{\top}_{i}\beta\right)^{2}+3\sin(X_{i,1})+1.1+\tau_{0}D_{i}+\varepsilon_{i}$.
For propensity score estimation, we employ a three-layer neural network with an Exponential Linear Unit (ELU) activation function for each hidden layer ($100$ nodes per layer). The final output layer applies a sigmoid function to ensure that the estimated propensity scores remain in $(0,1)$. We use this model for our method, logistic regression, and CBPS. For RieszNet, we adopt the DragonNet architecture proposed in Shi et al. (2019), following the original implementation by Chernozhukov et al. (2022a). For each propensity score estimation method, including ours, we compute both the IPW and AIPW estimators using the estimated scores. Additionally, we include the direct method (DM) estimator with neural networks for comparison. In each case, the expected conditional outcomes are estimated using a three-layer neural network (100 nodes per hidden layer, with ELU activation). As a baseline, we also consider the DM estimator with linear models.
The sample size is fixed at $n=3000$. As noted earlier, we evaluate two values of $K$ ($K=3$ and $K=10$) and two outcome-model specifications (Model 1 and Model 2), resulting in four experimental configurations. Each setting is repeated 500 times. We report the MSEs and biases of the resulting ATE estimates in Table 1 for $n=3000$. Overall, the results indicate that our direct bias-correction approach achieves competitive or superior estimation accuracy compared with logistic regression and CBPS, highlighting the benefits of explicitly estimating the bias-correction term in the ATE context. RieszNet tends to outperform our method, but we consider this to be partly due to differences in the regression models. While RieszNet employs DragonNet, we use a simpler implementation. We do not employ such models, as model complexity is not our primary focus. Nevertheless, we emphasize that our method outperforms most existing approaches while exhibiting comparable performance to RieszNet.
8 Conclusion
This study proposed direct bias-correction term estimation in ATE estimation. Instead of focusing on estimating the propensity score itself, our approach directly minimizes the estimation error of the bias-correction term, leveraging empirical risk minimization techniques. We demonstrated that this direct approach enhances estimation accuracy by avoiding the intermediate step of propensity score estimation. Additionally, our method was analyzed through the lens of Bregman divergence minimization, providing a generalized framework.
Appendix A Density-Ratio Estimation (DRE)
Given two probability distributions $P$ and $Q$ over a common space $\mathcal{X}$, the density ratio function is defined as
$r_{0}(x)\coloneqq\frac{p(x)}{q(x)},$
where $p(x)$ and $q(x)$ denote the density functions of $P$ and $Q$, respectively. DRE is a fundamental problem in statistical learning, with applications in importance sampling, anomaly detection, and covariate shift adaptation.
In DRE, estimating the two densities separately can magnify estimation errors, whereas directly modeling and estimating the density ratio can lead to improved accuracy. Thus, the aim of DRE is to estimate the density ratio in an end-to-end manner by directly optimizing a single objective. Various methods for DRE have been proposed (Huang et al., 2007; Gretton et al., 2009; Qin, 1998; Cheng & Chu, 2004; Nguyen et al., 2010; Kato et al., 2019), many of which can be generalized as instances of Bregman divergence minimization (Sugiyama et al., 2011; Kato & Teshima, 2021).
Let $\mathcal{R}$ be a hypothesis class for $r_{0}$, consisting of functions $r\colon\mathcal{X}\to{\mathbb{R}}$. The goal of direct DRE is to find an optimal function $r^{*}\in\mathcal{R}$ that best approximates $r_{0}$. A natural approach is to minimize the expected squared error:
${\mathbb{E}}_{P}\left[\big(r_{0}(X)-r(X)\big)^{2}\right].$
However, since $r_{0}(x)$ is unknown, direct minimization of this objective is infeasible.
Instead, we derive an equivalent formulation that does not require knowledge of $r_{0}$. Specifically, we show that minimizing the expected squared error is equivalent to minimizing the following alternative objective:
$2{\mathbb{E}}_{Q}\left[r(X)\right]+{\mathbb{E}}_{P}\left[r(X)^{2}\right].$
This transformation enables empirical risk minimization without explicit access to the true density ratio.
Furthermore, we extend this framework by providing theoretical guarantees on the estimation error using tools from empirical process theory. From the perspective of Bregman divergence minimization, we establish a generalized methodology for DRE that accommodates various estimation strategies.
Finally, we present numerical experiments that demonstrate the effectiveness of our approach in practical scenarios, including importance weighting and outlier detection.
Appendix B Bregman divergence
Here, we show the equivalence between
$r^{*}=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}^{\dagger}_{g}\big(r_{0}\mid r\big),$
where $\mathrm{BR}^{\dagger}_{g}\big(r_{0}\mid r\big)\coloneqq\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(g(r_{0}(d,X))-g(r(d,X))-\partial g(r(d,x))\Big(r_{0}(d,X)-r(d,X)\Big)\right)\Big]$, and
$r^{*}=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}_{g}\big(r\big),$
where $\mathrm{BR}_{g}\big(r\big)=\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\Big(-g(r(d,X))+\partial g(r(d,X))r(d,X)\Big)-\partial g(r(d,X))\Big]$.
This can be shown as follows:
$r^{*}=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}^{\dagger}_{g}\big(r_{0}\mid r\big)$
$=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(g(r_{0}(d,X))-g(r(d,X))-\partial g(r(d,x))\Big(r_{0}(d,X)-r(d,X)\Big)\right)\Big]$
$=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(-g(r(d,X))-\partial g(r(d,x))\Big(r_{0}(d,X)-r(d,X)\Big)\right)\Big]$
$=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(-g(r(d,X))-\partial g(r(d,x))r(d,X)\right)\Big]-{\mathbb{E}}\Big[\mathbbm{1}[D=d]\partial g(r(d,x))r_{0}(d,X)\Big]$
$=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(-g(r(d,X))-\partial g(r(d,x))r(d,X)\right)\Big]-{\mathbb{E}}\Big[\partial g(r(d,x))\Big].$
Here, we used
${\mathbb{E}}[\mathbbm{1}[D=1]r_{0}(1,X)\mid X]={\mathbb{E}}[e_{0}(X)r_{0}(1,X)\mid X]=1.$
We can show the same result for the case with $D=0$.
Appendix C Estimation of the average treatment effect for the treated (ATT)
Our method can also be applied to other estimands, such as the ATT, which is defined as
$\alpha_{0}\coloneqq{\mathbb{E}}\big[Y(1)-Y(0)\mid D=1\big].$
The IPW and AIPW estimators designed for the ATT are given by
IPW estimator.
$\widetilde{\alpha}^{\mathrm{IPW}}\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\frac{\mathbbm{1}[D_{i}=1]Y_{i}}{\pi_{0}}-\frac{e_{0}(X_{i})\mathbbm{1}[D_{i}=0]Y_{i}}{\pi_{0}(1-e_{0}(X_{i}))}\right)=\frac{1}{n}\sum^{n}_{i=1}\left(\frac{\mathbbm{1}[D=1]}{\pi_{0}}-\frac{e_{0}(X)\mathbbm{1}[D=0]}{\pi_{0}(1-e_{0}(X))}\right)Y_{i}$.
AIPW estimator.
$\widetilde{\alpha}^{\mathrm{AIPW}}\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\frac{\mathbbm{1}[D=1]}{\pi_{0}}-\frac{e_{0}(X)\mathbbm{1}[D=0]}{\pi_{0}(1-e_{0}(X))}\right)\left(Y_{i}-\mu_{0}(0,X_{i})\right)$,
where $\pi_{0}={\mathbb{E}}[\mathbbm{1}[D=1]]$.
Thus, the bias-correction term for ATT estimation is given as
$\widetilde{h}_{0}(D,X,e_{0},\pi_{0})\coloneqq\frac{\mathbbm{1}[D=1]}{\pi_{0}}-\frac{e_{0}(X)\mathbbm{1}[D=0]}{\pi_{0}(1-e_{0}(X))},$
where $\pi_{0}={\mathbb{E}}[\mathbbm{1}[D=1]]$.
Let $w_{0}(x)\coloneqq\frac{e_{0}(X)\mathbbm{1}[D=0]}{(1-e_{0}(X))}$. Then, we denote the bias-correction term as
$\widetilde{h}_{0}(D,X,w_{0},\pi_{0})\coloneqq\frac{\mathbbm{1}[D=1]}{\pi_{0}}-\frac{w_{0}(X)\mathbbm{1}[D=0]}{\pi_{0}}.$
Let $\mathcal{W}$} be a set of functions $w\colon{\mathcal{X}}\to{\mathbb{R}}_{+}$. Then, we define the following least squares:
$w^{*}\coloneqq\operatorname*{arg\,min}_{r\in\mathcal{R}}{\mathbb{E}}\left[\big(\widetilde{h}(D,X;r_{0},\pi_{0})-\widetilde{h}(D,X;r,\pi_{0})\big)^{2}\right].$
Note that we use $\pi_{0}$ itself. We can show that this least squares is equivalent to
$w^{*}=\operatorname*{arg\,min}_{r\in\mathcal{R}}\left\{-2{\mathbb{E}}_{1}\left[w(X)\right]+{\mathbb{E}}\left[w(X)^{2}\mathbbm{1}[D=0]\right]\right\},$
where ${\mathbb{E}}_{1}$ is expectation over the treated group ($p(x\mid d=1)$). The empirical version of this risk is given as
$\widehat{w}\coloneqq\operatorname*{arg\,min}_{r\in\mathcal{R}}\left\{-2\frac{1}{\sum^{n}_{i=1}\mathbbm{1}[D_{i}=1]}\sum^{n}_{i=1}\mathbbm{1}[D_{i}=1]w(X_{i})+\frac{1}{n}\sum^{n}_{i=1}w(X_{i})^{2}\right\},$
We can demonstrate the equivalence between the two least-squares formulations as follows:
$w^{*}$ $=\operatorname*{arg\,min}_{r\in\mathcal{R}}{\mathbb{E}}\left[\big(\widetilde{h}(D,X;r_{0},\pi_{0})-\widetilde{h}(D,X;r,\pi_{0})\big)^{2}\right]$
$=\operatorname*{arg\,min}_{r\in\mathcal{R}}{\mathbb{E}}\left[\big(w_{0}(X)\mathbbm{1}[D=0]-w(X)\mathbbm{1}[D=0]\big)^{2}\right]$
$=\operatorname*{arg\,min}_{r\in\mathcal{R}}{\mathbb{E}}\left[-2w_{0}(X)w(X)\mathbbm{1}[D=0]+w(X)^{2}\mathbbm{1}[D=0]\right].$
To see this equivalence, consider
${\mathbb{E}}\left[w_{0}(X)w(X)\mathbbm{1}[D=0]\right]$
$={\mathbb{E}}\left[{\mathbb{E}}\left[w_{0}(X)w(X)(1-e_{0}(X))\right]\right]$
$={\mathbb{E}}\left[e_{0}(X)w(X)/\pi_{0}\right]$
$=\int\frac{1}{\pi_{0}}e_{0}(x)w(x)\pi_{0}(x){\mathrm{d}}x$
$=\int\frac{1}{\pi_{0}}\frac{\pi_{0}p_{0}(x\mid d=1)}{p_{0}(x)}w(x)p_{0}(x){\mathrm{d}}x$
$=\int p_{0}(x\mid d=1)w(x){\mathrm{d}}x.$
This confirms the equivalence between the two least-squares objectives.
Appendix D Preliminary
This section introduces notions that are useful for the theoretical analysis.
D.1 Rademacher complexity
Let $\sigma_{1},\dots,\sigma_{n}$ be $n$ independent Rademacher random variables; that is, independent random variables for which $P(\sigma_{i}=1)=P(\sigma_{i}=-1)=1/2$. Let us define
$\mathfrak{R}_{n}f\coloneqq\frac{1}{n}\sum^{n}_{i=1}\sigma_{i}f(X_{i}).$
Additionally, given a class $\mathcal{F}$}, we define
$\mathfrak{R}_{n}{\mathcal{F}}\coloneqq\sup_{f\in{\mathcal{F}}}\mathfrak{R}_{n}f.$
Then, we define the Rademacher average as ${\mathbb{E}}[\mathfrak{R}_{n}{\mathcal{F}}]$ and the empirical Rademacher average as ${\mathbb{E}}_{\sigma}[\mathfrak{R}_{n}{\mathcal{F}}\mid X_{1},\dots,X_{n}]$.
D.2 Local Rademacher complexity bound
Let $\mathcal{F}$} be a class of functions that map $\mathcal{X}$} into $[a,b]$. For $f\in{\mathcal{F}}$, let us define
$Pf\coloneqq{\mathbb{E}}[f(X)],$
$P_{n}f\coloneqq\frac{1}{n}\sum^{n}_{i=1}f(X_{i}).$
We introduce the following result about the Rademacher complexity.
Proposition D.1 (From Theorem 2.1 in Bartlett et al. (2005)).
Let $\mathcal{F}$} be a class of functions that map $\mathcal{X}$} into $[a,b]$. Assume that there is some $r>0$ such that for every $f\in{\mathcal{F}}$, $\mathrm{Var}(f(X))\leq r$. Then, for every $z>0$, with probability at least $1-\exp(-z)$, it holds that
$\sup_{f\in{\mathcal{F}}}\Big(Pf-P_{n}f\Big)\leq\inf_{\alpha>0}\left\{2(1+\alpha){\mathbb{E}}[\mathfrak{R}_{n}f]+\sqrt{\frac{2rx}{n}}+(b-a)\left(\frac{1}{3}+\frac{1}{\alpha}\right)\frac{z}{n}\right\}.$
D.3 Bracketing entropy
We define the bracketing entropy. For a more detailed definition, see Definition 2.2 in van de Geer (2000).
Definition D.2.
Bracketing entropy. Given a class of functions $\mathcal{F}$}, the logarithm of the smallest number of balls in a norm $\|\cdot\|_{2,P}$ of radius $\delta>0$ needed to cover $\mathcal{F}$} is called the $\delta$-entropy with bracketing of $\mathcal{F}$} under the $L_{2}(P)$ metric, denoted by $H_{B}(\delta,{\mathcal{F}},P)$.
D.4 Talagrand’s concentration inequality
We introduce Talagrand’s lemma.
Proposition D.3 (Talagrand’s Lemma).
Let $\phi\colon{\mathbb{R}}\to{\mathbb{R}}$ be a Lipschitz continuous function with a Lipschitz constant $L>0$. Then, it holds that
$\mathfrak{R}_{n}(\phi\circ{\mathcal{F}})\leq L\mathfrak{R}_{n}({\mathcal{F}}).$
Appendix E Basic inequalities
Recall that we have defined an estimator $\widehat{r}$ as follows:
$\widehat{r}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\widehat{{\mathcal{L}}}_{n}(r)+\lambda J(r),$
where $J(r)$ is some regularization term.
Throughout the proof, we use the following basic inequalities that hold for $\widehat{r}$.
Proposition E.1.
The estimator $\widehat{r}$ satisfies the following inequality:
[…]
$\leq\frac{1}{n}\sum^{n}_{i=1}\Big(-2r_{0}(1,X_{i})-2r_{0}(0,X_{i})+\mathbbm{1}[D_{i}=1]r_{0}(1,X_{i})^{2}+\mathbbm{1}[D_{i}=0]r_{0}(0,X_{i})^{2}\Big)+\lambda J(r_{0}).$
For a function $f\colon{\mathcal{X}}\to{\mathbb{R}}$ and $X$ following $P$, let us denote the sample mean and expectation as
$\int f{\mathrm{d}}P_{n}$ $\coloneqq\int f(\cdot){\mathrm{d}}P_{n}\coloneqq\frac{1}{n}\sum^{n}_{i=1}f(X_{i}),$
$\int f{\mathrm{d}}P$ $\coloneqq\int f(\cdot){\mathrm{d}}P\coloneqq{\mathbb{E}}[f(X)].$
Following this notation, we also express this inequality as
$\int-2\big(\widehat{r}_{n}(1,\cdot)-r_{0}(1,\cdot)\big){\mathrm{d}}P_{n}+\int-2\big(\widehat{r}_{n}(0,\cdot)-r_{0}(0,\cdot)\big){\mathrm{d}}P_{n}$
$\ \ \ +\int\mathbbm{1}[D_{i}=1]\Big(\widehat{r}_{n}(1,\cdot)^{2}-r_{0}(1,\cdot)^{2}\Big){\mathrm{d}}P_{n}+\int\mathbbm{1}[D_{i}=0]\Big(\widehat{r}_{n}(0,\cdot)^{2}-r_{0}(0,X)^{2}\Big){\mathrm{d}}P_{n}$
$\ \ \ +\lambda J(\widehat{r})-\lambda J(r_{0})\leq 0.$
Proposition E.2.
The estimator $\widehat{r}$ satisfies the following inequality:
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}$
$\leq\sum_{d\in\{1,0\}}\Bigg(\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\int e_{0}(d,\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\Bigg)$
$\ \ \ +\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{n}+\int e_{0}(d\mid X)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}\Bigg)$
[…]
Proof of Proposition E.1 is trivial. We prove Proposition E.2 below.
Proof.
Let $e_{0}(1\mid x)=e_{0}(x)$ and $e_{0}(0\mid x)=1-e_{0}(x)$. Then, the following holds for the $L_{2}$ estimation error:
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}$
$=\left\|\mathbbm{1}[D=1]\big(\widehat{r}(1,X)-r_{0}(1,X)\big)-\mathbbm{1}[D=0]\big(\widehat{r}(0,X)-r_{0}(0,X)\big)\right\|^{2}_{P_{0}}$
$=\left\|e_{0}(1\mid\cdot)\big(\widehat{r}(1,X)-r_{0}(1,X)\big)\right\|^{2}_{P_{0}}+\left\|e_{0}(1\mid 0)\big(\widehat{r}(0,X)-r_{0}(0,X)\big)\right\|^{2}_{P_{0}}$
$=\int-2\big(\widehat{r}_{n}(1,\cdot)-r_{0}(1,\cdot)\big){\mathrm{d}}P_{0}+\int-2\big(\widehat{r}_{n}(0,\cdot)-r_{0}(0,\cdot)\big){\mathrm{d}}P_{0}$
$\ \ \ +\int e_{0}(1\mid\cdot)\Big(\widehat{r}_{n}(1,\cdot)^{2}-r_{0}(1,\cdot)^{2}\Big){\mathrm{d}}P_{0}+\int e_{0}(0,\cdot)\Big(\widehat{r}_{n}(0,\cdot)^{2}-r_{0}(0,\cdot)^{2}\Big){\mathrm{d}}P_{0}.$
From Proposition E.1, we have
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}$
$\leq\int-2\big(\widehat{r}_{n}(1,\cdot)-r_{0}(1,\cdot)\big){\mathrm{d}}P_{0}+\int-2\big(\widehat{r}_{n}(0,\cdot)-r_{0}(0,\cdot)\big){\mathrm{d}}P_{0}$
$\ \ \ +\int e_{0}(1\mid\cdot)\Big(\widehat{r}_{n}(1,\cdot)^{2}-r_{0}(1,\cdot)^{2}\Big){\mathrm{d}}P_{0}+\int e_{0}(0,\cdot)\Big(\widehat{r}_{n}(0,\cdot)^{2}-r_{0}(0,\cdot)^{2}\Big){\mathrm{d}}P_{0}$
$\ \ \ -\int-2\big(\widehat{r}_{n}(1,\cdot)-r^{*}(1,\cdot)\big){\mathrm{d}}P_{n}-\int-2\big(\widehat{r}_{n}(0,\cdot)-r^{*}(0,\cdot)\big){\mathrm{d}}P_{n}$
$\ \ \ -\int e_{0}(1\mid\cdot)\Big(\widehat{r}_{n}(1,\cdot)^{2}-r^{*}(1,\cdot)^{2}\Big){\mathrm{d}}P_{n}-\int e_{0}(0,\cdot)\Big(\widehat{r}_{n}(0,\cdot)^{2}-r^{*}(0,\cdot)^{2}\Big){\mathrm{d}}P_{n}$
$\ \ \ -\lambda J(\widehat{r})+\lambda J(r_{0})$
$=\int-2\big(\widehat{r}_{n}(1,\cdot)-r_{0}(1,\cdot)\big){\mathrm{d}}P_{0}+\int-2\big(\widehat{r}_{n}(0,\cdot)-r_{0}(0,\cdot)\big){\mathrm{d}}P_{0}$
$\ \ \ +\int e_{0}(1\mid\cdot)\Big(\widehat{r}_{n}(1,\cdot)^{2}-r_{0}(1,\cdot)^{2}\Big){\mathrm{d}}P_{0}+\int e_{0}(0,\cdot)\Big(\widehat{r}_{n}(0,\cdot)^{2}-r_{0}(0,\cdot)^{2}\Big){\mathrm{d}}P_{0}$
$\ \ \ -\int-2\big(\widehat{r}_{n}(1,\cdot)-r^{*}(1,\cdot)\big){\mathrm{d}}P_{n}-\int-2\big(\widehat{r}_{n}(0,\cdot)-r^{*}(0,\cdot)\big){\mathrm{d}}P_{n}$
$\ \ \ -\int e_{0}(1\mid\cdot)\Big(\widehat{r}_{n}(1,\cdot)^{2}-r^{*}(1,\cdot)^{2}\Big){\mathrm{d}}P_{n}-\int e_{0}(0,\cdot)\Big(\widehat{r}_{n}(0,\cdot)^{2}-r^{*}(0,\cdot)^{2}\Big){\mathrm{d}}P_{n}$
$\ \ \ +\int-2\big(r_{0}(1,\cdot)-r_{0}(1,\cdot)\big){\mathrm{d}}P_{n}+\int-2\big(\widehat{r}_{n}(0,\cdot)-r^{*}(0,\cdot)\big){\mathrm{d}}P_{n}$
$\ \ \ +\int e_{0}(1\mid\cdot)\Big(r_{0}(1,\cdot)^{2}-r_{0}(1,\cdot)^{2}\Big){\mathrm{d}}P_{n}+\int e_{0}(0,\cdot)\Big(r_{0}(0,\cdot)^{2}-r_{0}(0,\cdot)^{2}\Big){\mathrm{d}}P_{n}$
$\ \ \ -\lambda J(\widehat{r})+\lambda J(r_{0}).$
Here, we used
$\int\mathbbm{1}[d=1]r_{0}(1,\cdot)^{2}{\mathrm{d}}P_{0}$ $=2\int\mathbbm{1}[d=1]r_{0}(1,\cdot)^{2}{\mathrm{d}}P_{0}-\int\mathbbm{1}[d=1]r_{0}(1,\cdot)^{2}{\mathrm{d}}P_{0}$
$=2\int r_{0}(1,\cdot){\mathrm{d}}P_{0}-\int\mathbbm{1}[d=1]r_{0}(1,\cdot)^{2}{\mathrm{d}}P_{0}.$
Appendix F Proof of Theorem 3.5
We show Theorem 3.5 by bounding
$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})$ (2)
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\int e_{0}(d,\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\Bigg)$
$\ \ \ +\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{n}+\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}\Bigg)$ (3)
in Proposition E.2.
Here, since $r_{0}\in{\mathcal{H}}$, it holds that $r^{*}=r_{0}$, which implies that
$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{n}+\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}\Bigg)=0.$
Therefore, we consider bounding the first sum
$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})+\int e_{0}(d,\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\Bigg).$
We can bound this term by using the empirical-process arguments.
F.1 Preliminary
The following proposition is originally presented in van de Geer (2000) and was rephrased by Kanamori et al. (2012) in a form that is convenient for the theoretical analysis in DRE.
Lemma F.1 (Lemma 5.13 in van de Geer (2000), Proposition 1 in Kanamori et al. (2012)).
Let $\mathcal{F}\subset L^{2}(P)$ be a function class and the map $I(f)$ be a complexity measure of $f\in\mathcal{F}$, where $I$ is a non-negative function on $\mathcal{F}$ and $I(f_{0})<\infty$ for a fixed $f_{0}\in\mathcal{F}$. We now define $\mathcal{F}_{M}=\{f\in\mathcal{F}:I(f)\leq M\}$ satisfying $\mathcal{F}=\bigcup_{M\geq 1}\mathcal{F}_{M}$. Suppose that there exist $c_{0}>0$ and $0<\gamma<2$ such that
$\sup_{f\in\mathcal{F}_{M}}\|f-f_{0}\|\leq c_{0}M,\ \sup_{\stackrel{{\scriptstyle f\in\mathcal{F}_{M}}}{{\|f-f_{0}\|_{L^{2}(P)}\leq\delta}}}\|f-f_{0}\|_{\infty}\leq c_{0}M,\quad\text{for all }\delta>0,$
and that $H_{B}(\delta,\mathcal{F}_{M},P)=O{M/\delta}^{\gamma}$. Then, we have
$\sup_{f\in\mathcal{F}}\frac{\left|\int(f-f_{0})d(P-P_{n})\right|}{D(f)}=O_{p}(1),\ (n\to\infty),$
where $D(f)$ is defined by
$D(f)=\max{\frac{\|f-f_{0}\|_{L^{2}(P)}^{1-\gamma/2}I(f)^{\gamma/2}}{\sqrt{n}}}{\frac{I(f)}{n^{2/(2+\gamma)}}}.$
F.2 Upper bound using the empirical-process arguments
Proposition F.2.
[From Lemma 2 in Kanamori et al. (2012)] Under the conditions of Theorem 3.5, for any $0<\gamma<2$, we have
$\left|\int\left(\widehat{r}(d,\cdot)-r^{*}(d,\cdot))({\mathrm{d}}(P_{0}-P_{n})\right)\right|$ $=O_{p}\left(\max\left\{\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left\|\widehat{r}(d,\cdot)\right\|^{\gamma/2}_{{\mathcal{H}}}}{\sqrt{n}},\frac{\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}}{n^{2/(2+\gamma)}}\right\}\right),$
$\left|\int\left(\widehat{r}^{2}(d,\cdot)-r^{*}(d,\cdot)^{2}\right)({\mathrm{d}}(P_{0}-P_{n}))\right|$ $=O_{p}\left(\max\left\{\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}},\frac{\left\|\widehat{r}(d,\cdot)\right\|^{2}_{{\mathcal{H}}}}{n^{2/(2+\gamma)}}\right\}\right),$
as $n\to\infty$.
F.3 Proof of Theorem 3.5
We prove Theorem 3.5 following the arguments in Kanamori et al. (2012).
Proof.
From Proposition E.2 and $r_{0}\in{\mathcal{R}}^{\mathrm{RKHS}}$, we have
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}+\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}$
$\leq\sum_{d\in\{1,0\}}\left|\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})\right|+\left|\int e_{0}(d,\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\right|+\lambda\|r_{0}\|^{2}_{{\mathcal{H}}}.$
From Proposition F.2, we have
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}+\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}$
$=\sum_{d\in\{1,0\}}O_{p}\left(\max\left\{\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}},\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right\}\right)+\lambda\|r_{0}\|^{2}_{{\mathcal{H}}}.$
We consider the following three possibilities:
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}+\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=O_{p}(\lambda),$ (4)
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}+\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right),$ (5)
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}+\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$ (6)
The above inequalities are analyzed as follows:
Case (4).
We have
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}=O_{p}(\lambda),$
$\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=O_{p}(\lambda).$
Therefore, we have $\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|_{P_{0}}=O_{p}(\lambda^{1/2})$ and $\|\widehat{r}\|_{{\mathcal{H}}}=O_{p}(1)$.
Case (5).
We have
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right),$
$\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right).$
From the first inequality, we have
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{n^{1/(2+\gamma)}}\right).$
By using this result, from the second inequality, we have
$\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}$ $=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right)$
$=\sum_{d\in\{1,0\}}O_{p}\left(\left(\frac{1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}}{n^{1/(2+\gamma)}}\right)^{1-\gamma/2}\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right)$
$=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$
This implies that
$\|\widehat{r}\|_{{\mathcal{H}}}$ $=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{\lambda^{1/2}n^{2/(2+\gamma)}}\right)=o_{p}(1).$
Therefore, the following inequity is obtained.
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{1}{n^{1/(2+\gamma)}}\right)=O_{p}(\lambda^{1/2}).$
Case 6.
We have
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right),$
$\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$
As well as the argument in (5), we have $\|\widehat{r}\|_{{\mathcal{H}}}=o_{p}(1)$. Therefore, we have
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{1}{n^{1/(2+\gamma)}}\right)=O_{p}(\lambda^{1/2}).$
Appendix G Proof of Theorem 3.8
Our proof procedure mainly follows those in Kato & Teshima (2021) and Zheng et al. (2022). In particular, we are inspired by the proof in Zheng et al. (2022).
We prove Theorem 3.8 by proving the following lemma:
Lemma G.1.
Suppose that Assumption 3.7 holds. For any $n\geq\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})$, there exists a constant $C>0$ depending on $(\mu,\sigma,M)$ such that for any $\gamma>0$, with probability at least $1-\exp(-\gamma)$, it holds that
$\Big\|\widehat{f}_{n}-f_{0}\Big\|_{2}\leq C\left(\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log(n)}{n}}+\big\|f^{*}-f_{0}\big\|_{2}+\sqrt{\frac{\gamma}{n}}\right).$
As shown in Zheng et al. (2022), we can bound $\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log(n)$ by specifying neural networks and obtain Theorem 3.8.
G.1 Proof of Lemma G.1
We prove Lemma G.1 by bounding (2) in Proposition E.2.
To bound (2), we show several auxiliary results. Define
$\widehat{{\mathcal{F}}}^{f^{*},u}$ $\coloneqq\{f\in{\mathcal{F}}^{\mathrm{FNN}}\colon\frac{1}{n}\sum^{n}_{i=1}(f(X_{i})-f^{*}(X_{i}))^{2}\leq u\},$
$\overline{{\mathcal{G}}}^{f^{*},u}$ $\coloneqq\Big\{(f-f^{*})\colon f\in\widehat{{\mathcal{F}}}^{f^{*},u}\Big\},$
$\kappa^{u}_{n}(u)$ $\coloneqq{\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},u}\right],$
$u^{\dagger}$ $\coloneqq\inf\big\{u\geq 0\colon\kappa^{u}_{n}(s)\leq s^{2}\ \ \ \forall s\geq u\big\}.$
Here, we show the following two lemmas:
Lemma G.2 (Corresponding to (26) in Zheng et al. (2022)).
Suppose that the conditions in Lemma G.1 hold. Then, for any $z>0$, with probability $1-\exp(-z)$ it holds that
$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,x)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{n}+\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}\Bigg)$
$\leq C\left(\|f^{*}(X)-f_{0}(D,X)\|^{2}_{2}+\|f^{*}(X)-f_{0}(X)\|_{2}\sqrt{\frac{z}{n}}+\frac{16Mz}{3n}\right).$
Lemma G.3 (Corresponding to (29) in Zheng et al. (2022)).
Suppose that the conditions in Lemma G.1 hold. If there exists $u_{0}>0$ such that
$\|\widehat{f}(X)-f^{*}(X)\|_{2}\leq u_{0},$
then it holds that
$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})+\int e_{0}(d\mid\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\Bigg)$
$\leq C\left({\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},r}\right]+u_{0}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right).$
Additionally, we use the following three propositions directly from Zheng et al. (2022).
Proposition G.4 (From (32) in Zheng et al. (2022)).
Let $u>0$ be a positive value such that
$\|f-f_{0}\|_{2}\leq u$
for all $f\in{\mathcal{F}}$. Then, for every $z>0$, with probability at least $1-2\exp(-z)$, it holds that
$\sqrt{\frac{1}{n}\sum^{n}_{i=1}\big(f(X_{i})-f_{0}(X_{i})\big)^{2}}\leq 2u.$
Proposition G.5 (Corresponding to (36) in Step 3 of Zheng et al. (2022)).
Suppose that the conditions in Lemma G.1 hold. Then, there exists a universal constant $C>0$ such that
$u^{\dagger}\leq CM\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log(n)}{n}}.$
Proposition G.6 (Upper bound of the Rademacher complexity).
Suppose that the conditions in Lemma G.1 hold. If $n\geq\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})$, $r_{0}\geq 1/n$, and $n\geq(2eM)^{2}$, we have
${\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},r}\right]\leq Cr_{0}\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log n}{n}}.$
Then, we prove Lemma G.1 as follows:
Proof of Lemma G.1.
If there exists $u_{0}>0$ such that
$\|\widehat{f}(X)-f^{*}(X)\|_{2}\leq u_{0},$
then from (2) and Lemmas G.2 and G.3, for every $z>0$, there exists a constant $C>0$ independent $n$ such that
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}$
$\leq C\left(\|f^{*}-f_{0}\|_{2}\sqrt{\frac{z}{n}}+\frac{16Mz}{3n}+u_{0}\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log n}{n}}+u_{0}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right).$ (7)
This result implies that if $\sqrt{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})}$, then there exists $n_{0}$ such that for all $n>n_{0}$, there exists $u_{1}<u_{0}$ such that
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}\leq u_{1}.$
For any $z>0$, define $\overline{u}$ as
$\overline{u}_{z}\geq\max\left\{\sqrt{\log(n)/n},4\sqrt{3}M\sqrt{z/n},u^{\dagger}\right\}.$
Define a subspace of ${\mathcal{F}}^{\mathrm{FNN}}$ as
${\mathcal{S}}^{\mathrm{FNN}}(f_{0},\overline{u}_{z}\coloneqq\big\{f\in{\mathcal{F}}^{\mathrm{FNN}}\colon\|f-f_{0}\|\leq\overline{u}_{z}\big\}.$
Define
$\ell\coloneqq\lfloor\log_{2}(2M/\sqrt{\log(n)/n})\rfloor.$
Using the definition of subspaces, we divide ${\mathcal{F}}^{\mathrm{FNN}}$ into the following $\ell+1$ subspaces:
$\overline{{\mathcal{S}}}^{\mathrm{FNN}}_{0}\coloneqq$ ${\mathcal{S}}^{\mathrm{FNN}}(f_{0},\overline{u}),$
$\overline{{\mathcal{S}}}^{\mathrm{FNN}}_{1}\coloneqq$ ${\mathcal{S}}^{\mathrm{FNN}}(f_{0},\overline{u})\backslash{\mathcal{S}}^{\mathrm{FNN}}(f_{0},\overline{u}),$
$\quad\vdots$
$\overline{{\mathcal{S}}}^{\mathrm{FNN}}_{\ell}\coloneqq$ ${\mathcal{S}}^{\mathrm{FNN}}(f_{0},2^{\ell}\overline{u})\backslash{\mathcal{S}}^{\mathrm{FNN}}(f_{0},2^{\ell-1}\overline{u}).$
Since $\overline{u}_{z}>u^{\dagger}$, from the definition of $u^{\dagger}$, we have
$\overline{u}^{2}_{z}\leq\kappa^{u}_{n}(\overline{u}).$
If there exists $j\leq\ell$ such that $\widehat{f}\in\overline{{\mathcal{S}}}^{\mathrm{FNN}}_{j}$, then from (G.1), for every $z>0$, with probability at least $1-8\exp(-z)$, there exists a constant $C>0$ independent of $n$ such that
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{2}$
$\leq C\left(2^{\ell-1}\overline{u}\left(\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log(n)}{n}}+\sqrt{\frac{z}{n}}\right)+\|f^{*}-f_{0}\|^{2}_{2}+\|f^{*}-f_{0}\|_{2}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right).$ (8)
Additionally, if
$C\left(\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log(n)}{n}}+\sqrt{\frac{z}{n}}\right)\leq\frac{1}{8}2^{j}\overline{u},$ (9)
$C\left(\|f^{*}-f_{0}\|^{2}_{2}+\|f^{*}-f_{0}\|_{2}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right)\leq\frac{1}{8}2^{2j}\overline{u}^{2}$ (10)
hold, then
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|_{2}\leq 2^{j-1}\overline{u}.$ (11)
Here, to obtain (11), we used $\overline{u}\geq\max\left\{\sqrt{\log(n)/n},4\sqrt{3}M\sqrt{z/n},u^{\dagger}\right\}$, (G.1), (9), and (10).
From Proposition G.5, it holds that
$u^{\dagger}\leq CM\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log(n)}{n}}.$
Therefore, we can choose $\overline{u}$ as
$\overline{u}\coloneqq C\left(\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log(n)}{n}}+\sqrt{\log(n)/n}+4\sqrt{3}M\sqrt{z/n}\right),$
where $C>0$ is a constant independent of $n$. ∎
G.2 Proof of Lemma G.2
From Proposition D.1, we have
$\int-2\Big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\Big){\mathrm{d}}P_{n}$
$\ \ \ \ \ \ \leq\int-2\Big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\Big){\mathrm{d}}P_{0}+\sqrt{2}C_{1}\|f^{*}(X)-f_{0}(X)\|\sqrt{\frac{z}{n}}+\frac{16C_{1}Mz}{3n},$
$\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}$
$\ \ \ \ \ \ \leq\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{0}+\sqrt{2}C_{2}\|f^{*}(X)-f_{0}(X)\|\sqrt{\frac{z}{n}}+\frac{16C_{2}Mz}{3n}.$
This is a direct consequence of Proposition D.1. Note that $r^{*}$ and $r_{0}$ are fixed, and it is enough to apply the standard law of large numbers; that is, we do not have to consider the uniform law of large numbers. However, we can still apply Proposition D.1, which is a general than the standard law of large numbers, with ignoring the Rademacher complexity part. Here, we also used
$\Big|r^{*}(d,x)-r_{0}(d,x)\Big|\leq\frac{1}{4}\Big|f^{*}(x)-f_{0}(x)\Big|,$
$\Big|r^{*}(d,x)^{2}-r_{0}(d,x)^{2}\Big|\leq\frac{1}{2}\Big|f^{*}(x)-f_{0}(x)\Big|.$
We have
$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{n}+\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}\Bigg)$
$\leq\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{0}+\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{0}$
$\ \ \ \ \ +\sqrt{2}C_{1}\|f^{*}-f_{0}\|\sqrt{\frac{z}{n}}+\frac{16C_{2}Mz}{3n}+\sqrt{2}C_{2}\|f^{*}-f_{0}\|\sqrt{\frac{z}{n}}+\frac{16C_{2}Mz}{3n}\Bigg)$
$\leq C\left(\|f^{*}-f_{0}\|^{2}_{2}+\|f^{*}-f_{0}\|\sqrt{\frac{z}{n}}+\frac{16CMz}{3n}\right).$
G.3 Proof of Lemma G.3
Let $g\coloneqq(f-f^{*})^{2}$. From the definition of FNNs, we have
$g\leq 4M^{2}$
Additionally, we assumed that $\|\widehat{f}-f^{*}\|_{2}\leq r_{0}$ holds. Then, it holds that $\mathrm{Var}_{P_{0}}(g)\leq 4M^{2}r^{2}$.
Here, we note that the followings hold for all $f$ ($r$):
$\Big|r(d,x)-r^{*}(d,x)\Big|\leq\frac{1}{4}\Big|f(x)-f^{*}(x)\Big|,$
$\Big|r(d,x)^{2}-r^{*}(d,x)^{2}\Big|\leq\frac{1}{2}\Big|f(x)-f^{*}(x)\Big|.$
Then, from Proposition D.1, for every $z>0$, with probability at least $1-\exp(-z)$, it holds that, for each $d\in\{1,0\}$ it holds that Then, it holds that
$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})+\int e_{0}(d\mid\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\Bigg)$
$\leq C\left({\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},r}\right]+r_{0}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right).$

### v2

@@
-Direct Bias-Correction Term Estimation for Propensity Scores and Average Treatment Effect Estimation
+Direct Bias-Correction Term Estimation for Average Treatment Effect Estimation
 Abstract
-This study considers the estimation of the average treatment effect (ATE).
-For ATE estimation, we estimate the propensity score through direct bias-correction term estimation.
-Let $\{(X_{i},D_{i},Y_{i})\}_{i=1}^{n}$ be the observations, where $X_{i}\in{\mathbb{R}}^{K}$ denotes $K$-dimensional covariates, $D_{i}\in\{0,1\}$ denotes a binary treatment assignment indicator, and $Y_{i}\in{\mathbb{R}}$ is an outcome.
-In ATE estimation, the bias-correction term $h_{0}(X_{i},D_{i})\coloneqq\frac{\mathbbm{1}[D_{i}=1]}{e_{0}(X_{i})}-\frac{\mathbbm{1}[D_{i}=0]}{1-e_{0}(X_{i})}$ plays an important role, where $e_{0}(X_{i})$ is the propensity score, the probability of being assigned treatment $1$.
-In this study, we propose estimating $h_{0}$ (or equivalently $e_{0}$) by directly minimizing the prediction error for $h_{0}$ without knowing $h_{0}$ in advance.
-After showing a basic result with least squares, we present a general framework for this direct bias-correction term estimation approach from the perspective of Bregman divergence minimization, which also generalizes the Riesz regression and covariate balancing.
+This study considers the estimation of the direct bias-correction term for estimating the average treatment effect (ATE).
+Let $\{(X_{i},D_{i},Y_{i})\}_{i=1}^{n}$ be the observations, where $X_{i}\in{\mathbb{R}}^{K}$ denotes $K$-dimensional covariates, $D_{i}\in\{0,1\}$ denotes a binary treatment assignment indicator, and $Y_{i}\in{\mathbb{R}}$ denotes an outcome.
+In ATE estimation, $h_{0}(D_{i},X_{i})\coloneqq\frac{\mathbbm{1}[D_{i}=1]}{e_{0}(X_{i})}-\frac{\mathbbm{1}[D_{i}=0]}{1-e_{0}(X_{i})}$ is called the bias-correction term, where $e_{0}(X_{i})$ is the propensity score.
+The bias-correction term is also referred to as the Riesz representer or clever covariates, depending on the literature, and plays an important role in construction of efficient ATE estimators.
+In this study, we propose estimating $h_{0}$ by directly minimizing the Bregman divergence between its model and $h_{0}$, which includes squared error and Kullback–Leibler divergence as special cases.
+Our proposed method is inspired by direct density ratio estimation methods and generalizes existing bias-correction term estimation methods, such as covariate balancing weights, Riesz regression, and nearest neighbor matching.
+Importantly, under specific choices of bias-correction term models and Bregman divergence, we can automatically ensure the covariate balancing property.
+Thus, our study provides a practical modeling and estimation approach through a generalization of existing methods.
 1 Introduction
@@
 Methods for estimating ATEs are typically designed to eliminate bias arising from treatment assignment and the estimation of nuisance parameters, aiming for (asymptotic) unbiasedness and efficiency.
-In ATE estimation, the propensity score, the probability of treatment assignment given covariates, plays a crucial role (Rosenbaum & Rubin, 1983).
-For example, the inverse probability weighting (IPW) estimator, also known as the Horvitz-Thompson estimator (Horvitz & Thompson, 1952), estimates the ATE by computing a weighted average of outcomes using the inverse of the propensity score.
-In the one-step bias correction method (van der Vaart, 2002), a bias-correction term depending on the propensity score is added to an initial estimator.
-In this study, we focus on the role of the propensity score as a bias-correction component and propose a novel method to directly estimate either the propensity score or the bias-correction term itself.
-The bias-correction term is fundamental to IPW-type estimators and the one-step bias correction method, and its accurate estimation can substantially improve ATE estimation.
-Importantly, our objective is not to estimate the propensity score, but rather the bias-correction term, in which the propensity score appears inversely and is weighted by the treatment indicator.
-Since estimating the propensity score itself is not the target, we hypothesize that directly estimating the bias-correction term will lead to improved performance.
-To this end, we propose estimating the bias-correction term by minimizing the empirical risk targeted for the true bias-correction term.
-The technical challenge is that the target variable remains unobserved, even when we directly estimate the bias-correction term or the propensity score.
-To address this issue, we employ techniques developed in the direct density-ratio estimation (DRE) literature (Sugiyama et al., 2012).
-In direct DRE, the goal is to minimize the empirical risk between the true density ratio and its model, even though the true density ratio is unknown.
-It is known that empirical risk minimization is feasible even without knowledge of the true propensity score.
-Since the inverse propensity score can be viewed as a density ratio, we can extend these existing methods to our setting.
-Our motivation is closely aligned with studies on covariate balancing (Imai & Ratkovic, 2013) and Riesz regression (Chernozhukov et al., 2022a), which also aim to improve ATE estimation by appropriately estimating the propensity score or the bias-correction term.
-Studies in covariate balancing focus on the balancing property of propensity score estimator and estimate them using the property.
-Chernozhukov et al.
-(2022a) proposes Riesz regression which represents the bias-correction term as the Riesz representer.
-Although the derivation process is different, we derive the objective function that is the same as Chernozhukov et al.
-(2022a) by using the DRE techniques.
-Further, we generalize our objective by using the Bregman divergence as well as DRE in Sugiyama et al.
-(2011).
-From this generalization, we connect our approach to the covariate balancing by showing the equivalence between our objective and empirical balancing (Chan et al., 2015).
 1.1 ATE estimators and bias correction
@@
 Let $e_{0}(X)=P_{0}(D=1\mid X)$ denote the probability of assigning treatment $1$ given covariates $X$, which is known as the propensity score.
-Throughout this study, we impose the following condition, commonly referred to as the overlap assumption:
+Throughout this study, we impose the following conditions, commonly referred to as the unconfoundedness and common support assumptions.
 Assumption 1.1.
-There exists a constant $C>0$ independent of $n$ such that $C<e_{0}(x)<1-C$ for all $x\in\mathcal{X}$.
+[…]
 When $e_{0}(x)$ is not constant, a distributional shift arises between the observed outcomes in the treatment and control groups, denoted by $\mathcal{G}_{1}$ and $\mathcal{G}_{0}$, respectively, where $\mathcal{G}_{d}\coloneqq\{i\in\{1,2,\dots,n\}\colon D_{i}=d\}$.
@@
 Bias-correction term. In both estimators, the term
-$h_{0}(D,X)\coloneqq h(D,X;1/e_{0})\coloneqq\frac{\mathbbm{1}[D=1]}{e_{0}(X)}-\frac{\mathbbm{1}[D=0]}{1-e_{0}(X)}$
+$h_{0}(D,X)\coloneqq h(D,X)\coloneqq\frac{\mathbbm{1}[D=1]}{e_{0}(X)}-\frac{\mathbbm{1}[D=0]}{1-e_{0}(X)}$
 is crucial.
@@
 A common approach is to estimate $e_{0}$ using logistic regression and then plug the resulting estimate $\widehat{e}^{{\mathrm{L}}}_{n}$ into $h$.
-Note that in automatic debiased machine learning, the term is also referred to as the Riesz representer (Chernozhukov et al., 2022b).
+Note that the bias-correction term is also referred to as the Riesz representer (Chernozhukov et al., 2021) or the clever covariates (van der Laan, 2006).
+We use the term bias-correction term because the Riesz representer is closely connected to the automatic debiased machine learning literature, and the clever covariates is closely connected to the targeted maximum likelihood estimation (TMLE) literature.
 For example, in a typical one-step bias correction, we first construct an ATE estimator as $\widehat{\tau}^{\mathrm{DM}}_{n}\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\widehat{\mu}_{n}(1,X)-\widehat{\mu}_{n}(0,X)\right)$, where $\widehat{\mu}_{n}$ is an estimator of $\mu_{0}$.
 This estimator is known as the direct method (DM) or naive plug-in estimator.
-To obtain an efficient estimator, we add the bias-correction term $\frac{1}{n}\sum^{n}_{i=1}h_{0}(X_{i},D_{i})$ to the first-stage DM estimator $\widehat{\tau}^{\mathrm{DM}}_{n}$, yielding the AIPW estimator.
-In this process, estimating the propensity score $e_{0}$ itself is a challenging problem, independent of ATE estimation.
-The well-known Vapnik principle states that “when solving a problem of interest, do not solve a more general problem as an intermediate step” (Vapnik, 1998).
-Following this principle, this study aims to estimate $h_{0}(D,X)$, or equivalently $e_{0}$, by directly minimizing the estimation error of $h_{0}(D,X)$.
-We emphasize that while $e_{0}$ is estimated in our approach, the primary estimation target is not $e_{0}$ but $h_{0}$.
+To obtain an efficient estimator, we add the bias-correction term $\frac{1}{n}\sum^{n}_{i=1}h_{0}(D_{i},X_{i})(Y_{i}-\widehat{\mu}_{n}(D_{i},X_{i}))$ to the first-stage DM estimator $\widehat{\tau}^{\mathrm{DM}}_{n}$, yielding the AIPW estimator.
+In this study, we propose a method to estimate the bias-correction term, also called the Riesz representer or the clever covariates.
+For example, we can estimate the bias-correction term by estimating the propensity score $e_{0}$ using the maximum likelihood estimation.
+However, our interest is not in propensity score estimation but in bias-correction term estimation.
+As the well-known Vapnik principle states, we should avoid such an intermediate problem and ideally aim to estimate the target objective in a more direct manner (Vapnik, 1998).
+Following this principle, this study considers estimating $h_{0}(D,X)$ by directly minimizing the estimation error for the true $h_{0}(D,X)$.
+The technical challenge is that the target objective $h_{0}$ is unknown.
+To address this issue, we employ techniques developed in the direct density-ratio estimation (DRE) literature (Sugiyama et al., 2012).
+In direct DRE, the goal is to minimize the empirical risk between the true density ratio and its model, even though the true density ratio is unknown.
+It is known that empirical risk minimization is feasible even without knowledge of the true propensity score.
+Since the inverse propensity score can be viewed as a density ratio, we can extend these existing methods to our setting.
+For causal inference researchers who are unfamiliar with DRE, we review the DRE literature in Appendix A.
+Our motivation is also closely aligned with studies on Riesz regression (Chernozhukov et al., 2021) and covariate balancing weights (Imai & Ratkovic, 2013b; Deville & Särndal, 1992), which also aim to estimate the bias-correction term in a direct manner.
+Studies in covariate balancing focus on the balancing property of propensity score estimator and estimate them using the property.
+Chernozhukov et al.
+(2021) proposes Riesz regression which represents the bias-correction term as the Riesz representer.
+Although the derivation process is different, we derive the objective function that is the same as Chernozhukov et al.
+(2021) by using the DRE techniques.
+Further, we generalize our objective by using the Bregman divergence as well as DRE in Sugiyama et al.
+(2011).
+From this generalization, we further connect our approach to the covariate balancing by showing the equivalence between our objective and empirical balancing through the duality arguments discussed in Zhao (2019) and Bruns-Smith et al.
+(2025).
 1.2 Our contributions
-Our first contribution is the proposal of a framework for direct bias-correction term estimation.
-We estimate the bias-correction term by directly minimizing the estimation error for the true bias-correction term $h_{0}$.
-To model the bias-correction term $h_{0}$, we define the inverse propensity score as $r_{0}(1,X)=\frac{1}{e_{0}(X)}$ and $r_{0}(0,X)=\frac{1}{1-e_{0}(X)}$.
-Let $\mathcal{R}$ be a model (hypothesis class) for $r_{0}$, and consider approximating $r_{0}$ using some $r\in\mathcal{R}$.
-We approximate $r_{0}$ by the minimizer $r^{*}\in\mathcal{R}$ of the mean squared error (MSE) ${\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}\right]$.
-Here, we estimate $h_{0}(D,X)=h(D,X;r_{0})$ using a model $h(D,X;r)$ that depends on $r$.
-We emphasize again that although we bypass the estimation of $r_{0}$, our primary estimation target is $h_{0}$, not $r_{0}$.
-Since this expected squared error involves the unknown function $r_{0}$, direct optimization is infeasible.
-However, we show that minimizing this expected squared error with respect to $r$ is equivalent to minimizing ${\mathbb{E}}\left[-2\left(r(1,X)+r(0,X)\right)+h(D,X;r)^{2}\right]$, which does not depend on the unknown function.
-That is, we establish the equivalence: $r^{*}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}{\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}\right]=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}{\mathbb{E}}\left[-2\left(r(1,X)+r(0,X)\right)+h(D,X;r)^{2}\right]$.
-The obtained squared error can then be approximated using a sample-based empirical risk function.
-Our second main contribution is the theoretical analysis of the estimator obtained via direct bias-correction term estimation.
+This study has the following four contributions: (i) a general framework for directly estimating the bias-correction term (also called the Riesz representer or clever covariates) via Bregman divergence minimization; (ii) our proposed framework includes Riesz regression in Chernozhukov et al.
+(2021) and the tailored loss in Zhao (2019) as special cases; (iii) under our framework, we show that there are appropriate choices of bias-correction term models and Bregman divergences under which covariate balancing is automatically realized as the dual of the Bregman divergence minimization problem (automatic covariate balancing); (iv) we provide a theoretical analysis of the estimator.
+Our first contribution is the proposal of a framework for direct bias-correction term estimation via Bregman divergence minimization.
+We estimate the bias-correction term by directly minimizing the estimation error of the true bias-correction function $h_{0}$, measured by the Bregman divergence, $\mathrm{BR}^{\dagger}_{g}\big(h_{0}\mid h\big)\coloneqq{\mathbb{E}}\Big[g\big(h_{0}(D,X)\big)-g\big(h(D,X)\big)-\partial g\big(h(d,X)\big)\big(h_{0}(D,X)-h(D,X)\big)\Big]$, where $g$ is a differentiable and strictly convex function.
+By changing $g$, we can measure the error using various metrics, such as the squared loss or KL divergence loss.
+Since the Bregman divergence involves the unknown function $h_{0}$, direct optimization is infeasible.
+To address this issue, we propose minimizing an alternative objective function, defined as $\mathrm{BR}_{g}\big(h\big)\coloneqq{\mathbb{E}}\Big[-g\big(h(D,X)\big)+\partial g\big(h(D,X)\big)h(D,X)-\partial g\big(h(1,X)\big)-\partial g\big(h(0,X)\big)\Big]$.
+Minimizing the original Bregman divergence $\mathrm{BR}^{\dagger}_{g}\big(h\big)$ is equivalent to minimizing $\mathrm{BR}_{g}\big(h\big)$, which does not depend on the unknown function.
+That is, we establish the equivalence: $h^{*}\coloneqq\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\mathrm{BR}^{\dagger}_{g}\big(h_{0}\mid h\big)=\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\mathrm{BR}_{g}\big(h\big)$.
+The resulting objective function can then be approximated using an empirical risk function.
+Our second contribution is the unification of existing literature.
+Our proposed Bregman divergence minimization objective includes Riesz regression in Chernozhukov et al.
+(2021) (when using the squared loss) and the tailored loss in Zhao (2019) (when using the KL divergence loss).
+Furthermore, our framework also integrates covariate balancing methods (Imai & Ratkovic, 2013a; Hainmueller, 2012; Zubizarreta, 2015; Chan et al., 2015; Wong & Chan, 2017).
+If we use linear models to approximate the bias-correction term and train the model with the squared loss (Riesz regression), the dual problem coincides with the optimization problem in stable balancing weights.
+If we model the bias-correction term via the propensity score with logistic models and train the model with the KL divergence loss (tailored loss), the dual problem becomes the same as the optimization problem in entropy balancing weights.
+Kato (2025a), a subsequent work of this study, refers to this property as automatic covariate balancing.
+See Table 1 in Section 2 and Figure 1 in Appendix.
+Our third main contribution is the theoretical analysis of the estimator obtained via direct bias-correction term estimation.
 Since we estimate $r_{0}$ using empirical risk minimization, we establish bounds on the estimation error using empirical process theory.
@@
 Using standard ATE estimation techniques, we demonstrate that our method yields a $\sqrt{n}$-consistent ATE estimator.
-Our third main contribution is the generalization of our framework.
-From the perspective of Bregman divergence minimization, we extend our framework to provide a more general methodology for direct bias-correction term estimation.
-Under this framework, various estimation strategies can be incorporated to enhance bias-correction term estimation.
-Our fourth contribution is the unification of the existing literature: Riesz regression and covariate balancing.
-If we use the squared loss function in the Bregman divergence minimization, we obtain the same loss as the one used in Riesz regression.
-If we use a different function in the Bregman divergence, we recover a special case of the tailored loss for covariate balancing proposed in Zhao (2019), which is also equivalent to empirical balancing (Chan et al., 2015; Hainmueller, 2012).
-2 Direct bias-correction term estimation
-In this study, we consider estimating $h_{0}$ by minimizing the empirical risk associated with the MSE ${\mathbb{E}}\left[\big(h_{0}(D,X)-h(D,X)\big)^{2}\right]$ over $h:\{1,0\}\times\mathcal{X}\to{\mathbb{R}}$.
-Here, since $h_{0}$ contains the unknown function $r_{0}$, this minimization problem is inherently intractable.
-However, we demonstrate that empirical risk minimization remains feasible without explicit knowledge of $r_{0}$.
-2.1 Least squares
-Rather than directly modeling $h_{0}$, we model either $r_{0}$ or $e_{0}$ within $h_{0}$.
-That is, we model $h_{0}$ via a model of $r_{0}$ as $h(D,X;r)$, while $h_{0}(D,X)$ is given as $h_{0}(D,X;r_{0})$.
-We then aim to estimate $r_{0}$ or $e_{0}$.
-Given a set $\mathcal{R}$ of functions $r\colon{\mathcal{X}}\to(1,\infty)$, we approximate the bias-correction term by solving
-$r^{*}\coloneqq\operatorname*{arg\,min}_{r\in\mathcal{R}}{\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}\right].$
-If $r_{0}\in\mathcal{R}$, then it follows that $r_{0}(\cdot,\cdot)=r^{*}(\cdot)$ and $h(\cdot,\cdot;r^{*})=h_{0}(\cdot,\cdot)$.
-This MSE is associated with the MSE between the ATE $\tau_{0}$ and some estimator using the bias-correction term.
-Since $e_{0}(x)\in(0,1)$, a model $r(x)$ takes a value in $(1,\infty)$.
-Example.
-For example, for $\widehat{\tau}^{\mathrm{IPW}}(r)\coloneqq\frac{1}{n}\sum^{n}_{i=1}h(D_{i},X_{i};r)Y_{i}$, we have ${\mathbb{E}}\left[\left(\tau_{0}-\widehat{\tau}^{\mathrm{IPW}}(r)\right)^{2}\right]={\mathbb{E}}\left[\left(\tau_{0}-\widetilde{\tau}^{\mathrm{IPW}}+\widetilde{\tau}^{\mathrm{IPW}}-\widehat{\tau}^{\mathrm{IPW}}(r)\right)^{2}\right]={\mathbb{E}}\left[\left(\tau_{0}-\widetilde{\tau}^{\mathrm{IPW}}\right)^{2}\right]+{\mathbb{E}}\left[\left(\widetilde{\tau}^{\mathrm{IPW}}-\widehat{\tau}^{\mathrm{IPW}}(r)\right)^{2}\right]$, where recall that we defined $\widetilde{\tau}^{\mathrm{IPW}}$ in the Introduction.
-Here, we have ${\mathbb{E}}\left[\left(\widetilde{\tau}^{\mathrm{IPW}}-\widehat{\tau}^{\mathrm{IPW}}(r)\right)^{2}\right]=\frac{1}{n}{\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}{\mathbb{E}}\left[Y^{2}\mid D,X\right]\right]$.
-Thus, the MSE ${\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}\right]$ is closely connected to the MSE ${\mathbb{E}}\left[\left(\tau_{0}-\widehat{\tau}^{\mathrm{IPW}}(r)\right)^{2}\right]$.
-Direct estimation without $r_{0}$.
-The essential problem in this approach is that the target $h_{0}$ or $r_{0}$ is unknown; therefore, this optimization looks infeasible.
-However, even if we do not know $h_{0}$, we can reformulate the optimization problem into an equivalent form that does not involve $h_{0}$ as
-$r^{*}=\operatorname*{arg\,min}_{r\in\mathcal{R}}{\mathbb{E}}\Big[-2r(1,X)-2r(0,X)+\mathbbm{1}[D=1]r(1,X)^{2}+\mathbbm{1}[D=0]r(0,X)^{2}\Big].$
-This reformulation is derived as follows: $\min_{r\in\mathcal{R}}{\mathbb{E}}\left[\big(h(D,X;r_{0})-h(D,X;r)\big)^{2}\right]=\min_{r\in\mathcal{R}}{\mathbb{E}}\Big[h(D,X;r_{0})^{2}-2h(D,X;r_{0})h(D,X;r)+h(D,X;r)^{2}\Big]=\min_{r\in\mathcal{R}}{\mathbb{E}}\left[-2h(D,X;r_{0})h(D,X;r)+h(D,X;r)^{2}\right]=\min_{r\in\mathcal{R}}{\mathbb{E}}\left[-2\left(r(1,X)+r(0,X)\right)+h(D,X;r)^{2}\right]$.
-From the second to the third line, we omit the term $h_{0}(D,X)^{2}$ since it does not affect the optimization.
-From the third to the fourth line, we use the following for the first term: ${\mathbb{E}}\left[h(D,X;r_{0})h(D,X)\right]={\mathbb{E}}\Bigg[\left(\frac{\mathbbm{1}[D=1]}{e_{0}(X)}-\frac{\mathbbm{1}[D=0]}{1-e_{0}(X)}\right)\Big(\mathbbm{1}[D=1]r(1,X)-\mathbbm{1}[D=0]r(0,X)\Big)\Bigg]={\mathbb{E}}\left[\frac{\mathbbm{1}[D=1]}{e_{0}(X)}r(1,X)+\frac{\mathbbm{1}[D=0]}{1-e_{0}(X)}r(0,X)\right]={\mathbb{E}}\left[\frac{e_{0}(X)}{e_{0}(X)}r(1,X)+\frac{1-e_{0}(X)}{1-e_{0}(X)}r(0,X)\right]={\mathbb{E}}\left[r(1,X)+r(0,X)\right]$.
-Thus, surprisingly, we demonstrate that the least squares estimate for the unknown true bias-correction term $h_{0}$ can be defined by an objective function that does not explicitly include $h_{0}$ (or the unknown $r_{0}$) itself.
+As a side product of our contributions, we find that we can import various existing results from the DRE literature.
+Since Riesz regression is essentially the same as LSIF, various results about convergence rate analysis and optimization methods have already been established.
+For example, Kanamori et al.
+(2012) shows the convergence rate when using a reproducing kernel hilbert space (RKHS) for the density ratio, or equivalently the bias-correction term.
+Kato & Teshima (2021) shows the rate when using neural networks, which has been further refined in Zheng et al.
+(2022).
+Rhodes et al.
+(2020) and Kato & Teshima (2021) point out the overfitting problem characteristic of DRE estimation and propose techniques to avoid the problem.
+Lin et al.
+(2023) finds that nearest neighbor matching can be interpreted as density ratio estimation, and it can also be interpreted as a special case of LSIF or Riesz regression (See Appendix H).
+These findings not only help deepen our understanding of Riesz regression, but also prevent unnecessary reinvention.
+For example, the covariate adaption method proposed in Chernozhukov et al.
+(2025) uses Riesz regression, but it is essentially the same as covariate adaption with a density ratio estimated via LSIF (Kanamori et al., 2009), except for the regression adjustment.
+While Chernozhukov et al.
+(2022a) proposes neural networks and random forests for Riesz regression, the techniques for estimating the density ratio have also been proposed in the DRE literature (Kanamori et al., 2012; Abe & Sugiyama, 2019; Rhodes et al., 2020; Kato & Teshima, 2021).
+2 Bias-correction term estimation via Bregman divergence minimization
+In this study, we consider estimating $h_{0}$ by minimizing the empirical risk associated with the Bregman divergence between $h_{0}$ and its estimator $h\colon\{1,0\}\times\mathcal{X}\to{\mathbb{R}}$.
+2.1 Population Bregman divergence minimization
+Let $g\colon{\mathbb{R}}\to{\mathbb{R}}$ be a differentiable and strictly convex function.
+Given $d\in\{1,0\}$, we define the Bregman divergence between $h_{0}$ and $h$ as $\mathrm{br}^{\dagger}_{g}\big(h_{0}(d,x)\mid h(d,x)\big)\coloneqq g\big(h_{0}(d,x)\big)-g\big(h(d,x)\big)-\partial g\big(h(d,x)\big)\big(h_{0}(d,x)-h(d,x)\big)$, where $\partial g$ denotes the derivative of $g$.
+Then, we define the average Bregman divergence as $\mathrm{BR}^{\dagger}_{g}\big(h_{0}\mid h\big)\coloneqq{\mathbb{E}}\Big[g\big(h_{0}(D,X)\big)-g\big(h(D,X)\big)-\partial g\big(h(d,X)\big)\big(h_{0}(D,X)-h(D,X)\big)\Big]$.
+Then, we estimate $h_{0}$ by $h^{*}=\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\mathrm{BR}^{\dagger}_{g}\big(h_{0}\mid h\big)$.
+By dropping the term that is irrelevant to learning, we have
+$h^{*}=\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\mathrm{BR}_{g}\big(h\big),$
+$\text{where}\quad\mathrm{BR}_{g}\big(h\big)\coloneqq{\mathbb{E}}\Big[-g\big(h(D,X)\big)+\partial g\big(h(D,X)\big)h(D,X)-\partial g\big(h(1,X)\big)+\partial g\big(h(0,X)\big)\Big].$
+This can be shown as follows:
+$h^{*}=\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\left[\mathbbm{1}[D=d]\left(g(h_{0}(d,X))-g(h(d,X))-\partial g(h(d,X))\Big(h_{0}(d,X)-h(d,X)\Big)\right)\right]$
+$=\operatorname*{arg\,min}_{r\in{\mathcal{H}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(-g(h(d,X))-\partial g(h(d,x))\Big(h_{0}(d,X)-h(d,X)\Big)\right)\Big]$
+$=\operatorname*{arg\,min}_{r\in{\mathcal{H}}}\sum_{d\in\{1,0\}}\left({\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(-g(h(d,X))+\partial g(h(d,X))h(d,X)\right)\Big]-{\mathbb{E}}\Big[\mathbbm{1}[D=d]\partial g(h(d,x))h_{0}(d,X)\Big]\right)$
+$=\operatorname*{arg\,min}_{r\in{\mathcal{H}}}\left\{{\mathbb{E}}\Big[\left(-g(h(D,X))+\partial g(h(D,X))h(d,X)\right)\Big]-{\mathbb{E}}\Big[\partial g(h(1,X))\Big]+{\mathbb{E}}\Big[\partial g(h(0,X))\Big]\right\}.$
+Here, we dropped terms irrelevant to the optimization and used ${\mathbb{E}}[\mathbbm{1}[D=1]h_{0}(1,X)\mid X]={\mathbb{E}}[e_{0}(X)h_{0}(1,X)\mid X]=1$ and ${\mathbb{E}}[\mathbbm{1}[D=0]h_{0}(0,X)\mid X]=-1$.
+Thus, surprisingly, we demonstrate that the least squares estimate for the unknown true bias-correction term $h_{0}$ can be defined by an objective function that does not explicitly include $h_{0}$ itself.
 As discussed in the following subsection, this objective function can be easily approximated using observations.
-2.2 Empirical risk minimization
-We then estimate $r_{0}$ by solving the following empirical risk minimization:
-$\widehat{r}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\widehat{{\mathcal{L}}}_{n}(r)+\lambda J(r),$
-where the empirical risk is given as
-$\widehat{{\mathcal{L}}}_{n}(r)\coloneqq\frac{1}{n}\sum^{n}_{i=1}\Big(-2r(1,X_{i})-2r(0,X_{i})+\mathbbm{1}[D_{i}=1]r(1,X_{i})^{2}+\mathbbm{1}[D_{i}=0]r(0,X_{i})^{2}\Big),$
-and $J\colon{\mathcal{R}}\to{\mathbb{R}}^{+}$ is a regularization term with a penalty coefficient $\lambda>0$.
-This type of estimation method is referred to as least-squares importance fitting (LSIF, Kanamori et al., 2009) in the literature on density-ratio estimation (DRE).
+2.2 Empirical Bregman divergence minimization
+Then, we estimate the bias-correction term $h_{0}$ by minimizing an empirical Bregman divergence as
+$\widehat{h}_{n}\coloneqq\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\widehat{\mathrm{BR}}_{g}\big(h\big)+\lambda J(h),$
+where $J(h)$ is some regularization function and
+$\widehat{\mathrm{BR}}_{g}(h)\coloneqq\frac{1}{n}\sum^{n}_{i=1}\Big(-g(h(D_{i},X_{i}))+\partial g(h(D_{i},X_{i}))h(D_{i},X_{i})-\partial g(h(1,X_{i})+\partial g(h(0,X_{i}))\Big).$
+2.3 Losses for the bias-correction term estimation
+By changing $g$, we can obtain various loss functions for estimating the bias-correction term, as shown in the subsequent subsections.
+In particular, if we use the squared loss in the Bregman divergence, we obtain Riesz regression in Chernozhukov et al.
+(2021), which is originally called Least-Squares Importance Fitting (LSIF) in the DRE literature Kanamori et al.
+(2009).
+Note that kernel mean matching by Gretton et al.
+(2009) is also the same as, or a variant of, LSIF.
+If we use the KL divergence, we obtain the tailored loss in Zhao (2019), which is originally called KLIEP in the DRE literature Sugiyama et al.
+(2008).
+Furthermore, as we discuss in Section 3, if we use linear models for $h_{0}$ and train them with the squared loss, the covariate balancing property is automatically obtained, as shown in Bruns-Smith et al.
+(2025).
+If we model $h_{0}$ using the propensity score $e_{0}$ approximated via logistic models and train it with the tailored loss, the covariate balancing property is automatically obtained, as shown in Zhao (2019).
+We demonstrate the correspondence of the existing methods in Table 1.
+Also see Figure 1 in Appendix for the relationship among bias-correction term estimation via Bregman divergence minimization, density ratio estimation, and covariate balancing, summarized in Kato (2025a) and Kato (2025c).
+Table 1: Correspondence among DRE methods and bias-correction term estimation methods (BCE).
+DRE method BCE method $g(t)$
+LSIF (Kanamori et al., 2009) Riesz regression (Chernozhukov et al., 2021) $(t-1)^{2}/2$
+Kernel Mean Matching (Gretton et al., 2009) Stable balancing weights (Zubizarreta, 2015)
+UKL (Nguyen et al., 2010) Tailored loss (Zhao, 2019) $t\log(t)-t$
+KLIEP (Sugiyama et al., 2008) Entropy balancing weights (Hainmueller, 2012)
+Binary KL divergence $t\log(t)-(1+t)\log(1+t)$
+PULogLoss (Kato et al., 2019) $C\log\left(1-t\right)$
+$Ct\left(\log\left(t\right)-\log\left(1-t\right)\right)$ for $0<t<1$
+2.4 Squared loss
+Our least squares method for direct bias-correction term estimation can be obtained by using a squared loss $g^{\text{SL}}(h)=(h-1)^{2}$.
+By substituting this function into the Bregman divergence, we formulate the estimation problem as $h^{*}\coloneqq\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\mathrm{BR}_{g^{\text{SL}}}\big(h\big)$, where
+$\mathrm{BR}_{g^{\text{SL}}}\big(h\big)={\mathbb{E}}\Big[-2\big(h(1,X)-h(0,X)\big)+h(D,X)^{2}\Big].$
+Then, we estimate the bias-correction term as $\widehat{h}_{n}\coloneqq\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\widehat{\mathrm{BR}}_{g^{\mathrm{SL}}}\big(h\big)+\lambda J(h)$, where $\widehat{\mathrm{BR}}_{g^{\mathrm{SL}}}(h)=\frac{1}{n}\sum^{n}_{i=1}\big(-2\big(h(1,X_{i})-h(0,X_{i})\big)+h(D_{i},X_{i})^{2}\big)$.
+This objective function is the same as the one used in Chernozhukov et al.
+(2021).
+This type of estimation method is referred to as LSIF in density-ratio estimation (Kanamori et al., 2009).
+2.5 KL divergence loss
+Consider $g^{\mathrm{KL}}(h)=|h|\log|h|-|h|$, which is a convex function.
+By substituting this function into the Bregman divergence, we formulate the estimation problem as $h^{*}\coloneqq\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\mathrm{BR}_{g^{\mathrm{KL}}}\big(h\big)$, where
+$\mathrm{BR}_{g^{\mathrm{KL}}}\big(h\big)\coloneqq{\mathbb{E}}\Big[|h(D_{i},X_{i})|-\log(|h(1,X)|)-\log(|h(0,X)|)\Big].$
+Then, we estimate the bias-correction term as $\widehat{h}_{n}\coloneqq\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\widehat{\mathrm{BR}}_{g^{\mathrm{KL}}}\big(h\big)+\lambda J(h)$, where $\widehat{\mathrm{BR}}_{g^{\mathrm{KL}}}(h)=\frac{1}{n}\sum^{n}_{i=1}\big(|h(D_{i},X_{i})|-\log\big(|h(1,X_{i})|\big)-\log\big(|h(0,X_{i})|\big)\big)$.
+This estimation method corresponds to unnormalized Kullback–Leibler (UKL) minimization in DRE (Nguyen et al., 2010), which generalizes the KL importance estimation procedure (KLIEP).
+Also see Appendix B.
+2.6 Tailored loss (a variant of the KL divergence loss)
+Next, as a variant of the KL divergence loss, we propose the tailored loss.
+Let us redefine a model $\mathcal{H}$} as a set of functions $h(1,\cdot)\colon{\mathcal{X}}\to(1,\infty)$ and $h(0,\cdot)\colon{\mathcal{X}}\to(-1,-\infty)$; that is, we restrict the space of $h$.
+This restriction is justified from the form of $h_{0}$ and the common support assumption.
+Let us consider $g^{\text{TL}}(h)=(|h|-1)\log\left(|h|-1\right)-|h|$.
+By substituting this function, we obtain
+$\mathrm{BR}_{g^{\text{TL}}}\big(h\big)\coloneqq{\mathbb{E}}\Big[\log\left(|h(D,X)|-1\right)+|h(D,X)|-\log\left(|h(1,X)|-1\right)-\log\left(|h(0,X)|-1\right)\Big].$
+Note that it holds that $\mathrm{BR}_{g^{\text{TL}}}\big(h\big)\coloneqq{\mathbb{E}}\Big[-\mathbbm{1}[D=0]\log\left(|h(1,X)|-1\right)-\mathbbm{1}[D=1]\log\left(|h(0,X)|-1\right)+\mathbbm{1}[D=1]h(1,X)-\mathbbm{1}[D=0]h(0,X)\Big]$.
+Then, we estimate the bias-correction term as $\widehat{h}_{n}\coloneqq\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\widehat{\mathrm{BR}}_{g^{\text{TL}}}\big(h\big)$, where the empirical Bregman divergence becomes $\widehat{\mathrm{BR}}_{g^{\text{TL}}}(h)=\frac{1}{n}\sum^{n}_{i=1}\big(\mathbbm{1}[D_{i}=0]\log\big(|h(1,X_{i})|-1\big)+\mathbbm{1}[D_{i}=1]\log\big(|h(0,X_{i})|-1\big)+\mathbbm{1}[D_{i}=1]|h(1,X_{i})|-\mathbbm{1}[D_{i}=0]|h(0,X_{i})|\big)$.
+3 Automatic covariate balancing
+Under specific choices of Riesz regression models and Bregman divergence, we can automatically enforce the covariate balancing property.
+The key tool is the duality relationship between the Bregman divergence minimization problem and the covariate balancing optimization problem.
+This result is shown in Kato (2025a), and we introduce the result for reference.
+3.1 Linear models and squared loss
+Consider a linear model
+$h_{\beta}(D,X)=\Phi(D,X)^{\top}\beta,$
+where $\Phi\colon\{1,0\}\times{\mathcal{X}}\to{\mathbb{R}}^{p}$ is a basis function.
+For this model, using the squared loss (Riesz regression) automatically achieves covariate balancing, as discussed in Bruns-Smith et al.
+(2025).
+Specifically, under linear models, by duality, this MSE minimization problem is equivalent to solving
+$\min_{w\in{\mathbb{R}}^{n}}\|w\|^{2}_{2}\quad\text{s.t.}\ \ \sum^{n}_{i=1}w_{i}\Phi(D_{i},X_{i})-\left(\sum^{n}_{i=1}\Big(\Phi(1,X_{i})-\Phi(0,X_{i})\Big)\right)=\bm{0}_{p},$
+where $\bm{0}_{p}$ is the $p$-dimensional zero vector. This optimization problem matches that used to obtain stable weights (Zubizarreta, 2015).
+It enforces the covariate balancing condition $\sum^{n}_{i=1}\widehat{w}_{i}\Phi(D_{i},X_{i})-\left(\sum^{n}_{i=1}\Big(\Phi(1,X_{i})-\Phi(0,X_{i})\Big)\right)=\bm{0}_{p}$, where $\widehat{w}_{i}=\Phi(D_{i},X_{i})^{\top}\widehat{\beta}$.
+Another advantage of using linear models is that we can write the entire ATE estimation with a single linear model, as shown by Bruns-Smith et al. (2025).
+3.2 Logistic models and tailored loss
+We can model the Riesz representer by modeling the propensity score as
+$h_{\beta}(D,X)=\mathbbm{1}[D=1]r_{\beta}(1,X)-\mathbbm{1}[D=0]r_{\beta}(0,X),$
+where $r_{\beta}(1,X)=\frac{1}{e_{\beta}(X)}$, $r_{\beta}(0,X)=\frac{1}{1-e_{\beta}(X)}$, $e_{\beta}(X)\coloneqq\frac{1}{1+\exp\big(-\beta^{\top}\Phi(X)\big)}$, and $\Phi\colon{\mathcal{X}}\to{\mathbb{R}}^{p}$ is a basis function.
+Note that we do not include $D$, unlike the basis function used in linear models.
+For this model, if we use the KL-divergence–flavored convex function defined in Section 2.6, which corresponds to the tailored loss in Zhao (2019), we automatically achieve covariate balancing.
+Define $\widehat{\beta}\coloneqq\operatorname*{arg\,min}_{\beta}\frac{1}{n}\sum^{n}_{i=1}\sum_{d\in\{1,0\}}\Big(\mathbbm{1}[D_{i}=d]\left(-\log\left(\frac{1}{r_{\beta}(d,X_{i})-1}\right)+r_{\beta}(d,X_{i})\right)\Big)$, and denote $r_{\widehat{\beta}}$ by $\widehat{r}$.
+Under logistic models, by duality, the KL divergence-flavored loss is equivalent to solving
+$\min_{w\in(1,\infty)^{n}}\sum^{n}_{i=1}(w_{i}-1)\log(w_{i}-1)\quad\text{s.t.}\ \ \left(\sum^{n}_{i=1}\Big(\mathbbm{1}[D_{i}=1]w_{i}\Phi(X_{i})-\mathbbm{1}[D_{i}=0]w_{i}\Phi(X_{i})\Big)\right)=\bm{0}_{p}.$
+This optimization problem matches that used in entropy balancing (Hainmueller, 2012).
+Note that this objective function is derived from $\widehat{\mathrm{BR}}_{g^{\text{TL}}}(h)$ when we use the logistic model specified in this section.
+As a result, we obtain $\sum^{n}_{i=1}\Big(\mathbbm{1}[D_{i}=1]\widehat{w}_{i}\Phi(X_{i})-\mathbbm{1}[D_{i}=0]\widehat{w}_{i}\Phi(X_{i})\Big)=\bm{0}_{p}$, where $\widehat{w}_{i}=\widehat{r}(X_{i}).$
+This model has the advantage that we can use a basis function $\Phi(X)$ independent of $D$.
+Moreover, it naturally achieves covariate balance in the sense that the covariate distributions match between the treated and control groups.
+Additionally, it allows us to automatically impose nonnegativity on $h(1,X)$ and h(0,X)-h(0,X), which may be violated in linear models.
+Note that $h_{0}(1,X)=\frac{1}{e(X)}$ and $h_{0}(1,X)=\frac{1}{1-e(X)}$.
+3.3 Comparison
+We first discuss the advantages of using logistic models over linear models.
+One benefit of using logistic models is that we can simplify the basis function by making it independent of $D$.
+Furthermore, we can express covariate balancing in a clearer form as $\sum^{n}_{i=1}\Big(\mathbbm{1}[D_{i}=1]\widehat{w}_{i}\Phi(X_{i})-\mathbbm{1}[D_{i}=0]\widehat{w}_{i}\Phi(X_{i})\Big)=\bm{0}_{p}$, while under linear models, $\sum^{n}_{i=1}\widehat{w}_{i}\Phi(D_{i},X_{i})-\left(\sum^{n}_{i=1}\Big(\Phi(1,X_{i})-\Phi(0,X_{i})\Big)\right)=\bm{0}_{p}$ is attained, but it is somewhat harder to interpret.
+Moreover, using logistic models incorporates more information about the form of the bias-correction term, which includes the inverse propensity function.
+Logistic models also naturally impose restrictions such that $h(1,X)\in(1,\infty)$ and $h(0,X)\in(-\infty,-1)$ under the common support assumption.
+In contrast, if we use linear models, we can express the entire ATE estimator with a single linear model, as shown in Bruns-Smith et al.
+(2025).
+Furthermore, we can obtain the estimator of the bias-correction term as a closed-form solution.
+In addition, as discussed in Kato (2025b), a subsequent work of this study.
+nearest neighbor matching is also an instance of linear models trained via Riesz regression (squared loss).
+We introduce the result in Appendix H for reference.
+Ultimately, there is no clear dominance between the use of linear and logistic models.
+Moreover, we can also use more complex models, such as random forests and neural networks.
+The choice of model should be made based on the data and application, and once the model is selected, we can determine appropriate specifications that ensure covariate balancing automatically.
 4 Estimation error analysis
-This section provides an estimation error analysis for $r_{0}$ estimated by the direct bias-correction term estimation method.
-For simplicity, throughout the analysis, we use the following model for $e_{0}$: $e(x)=\frac{1}{1+\exp(-f(x))}$, where $f\colon{\mathcal{X}}\to{\mathbb{R}}$ is a function belonging to a set $\mathcal{F}$}.
-We use various models for $\mathcal{F}$}, including linear models, RKHS, and neural networks.
-Then, the model of $r_{0}$ is given as $r(1,x)=1+\exp(-f(x)),\quad r(0,x)=\exp(f(x))$.
-Let $f_{0}$ be the true function for $e_{0}$; that is, $e_{0}(X)=\frac{1}{1+\exp(-f_{0}(x))}$.
-3.1 Linear models
-First, we consider the case where $f_{0}$ belongs to a linear model.
-There exists $\theta_{0}\in\Theta^{K}$ such that $e_{0}(x)=\frac{1}{1+\exp\big(-x^{\top}\theta_{0}\big)}$, where $\Theta$ is a compact parameter space.
-Under this assumption, we have $r_{0}(1,X)=1+\exp\big(-x^{\top}\theta_{0}\big)$ and $r_{0}(0,X)=1+\exp\big(x^{\top}\theta_{0}\big)$.
-Let us use $r_{\theta}(1,X)=1+\exp\big(-x^{\top}\theta\big)$ and $r_{\theta}(0,X)=1+\exp\big(x^{\top}\theta\big)$ as a model for $r_{0}$.
-Then, we define the estimator as $\widehat{r}^{\mathrm{Lin}}_{n}=r_{\widehat{\theta}_{n}}$, where $\widehat{\theta}_{n}\coloneqq\operatorname*{arg\,min}_{\theta\in\Theta}\widehat{{\mathcal{L}}}_{n}(r_{\theta})$.
-Here, we set $\lambda=0$.
-We show the asymptotic normality of this estimator.
-Assumption 3.1.
-We assume the following: (i) $\theta_{0}$ is in the interior of $\Theta$; (ii) $\frac{\partial}{\partial\theta}\widehat{{\mathcal{L}}}_{n}(r_{\theta})$ is twice continuously differentiable on some neighborhood $\mathcal{M}$} of $\theta_{0}$ (with probability one); (iii) $\sqrt{n}\frac{\partial}{\partial\theta}\widehat{{\mathcal{L}}}_{n}(r_{\theta})\xrightarrow{d}{\mathcal{N}}(0,\Sigma_{0})$; (iv) for any sequence $\tilde{\theta}_{n}\xrightarrow{p}\theta_{0}$, it holds that $\frac{\partial^{2}}{\partial\theta\partial\theta^{T}}\widehat{{\mathcal{L}}}_{n}(r_{\theta})-B_{0}\xrightarrow{p}0$, for some non-stochastic $K\times K$ matrix $B_{0}$ that is nonsingular.
-Then, from the asymptotic theory of extremum estimators, we obtain the following result:
-Lemma 3.2.
-Suppose that Assumptions 1.1 and 3.1 hold. Then, it holds that $\sqrt{n}\left(\widehat{\theta}_{n}-\theta_{0}\right)\xrightarrow{{\mathrm{d}}}{\mathcal{N}}(0,B_{0}^{-1}\Sigma_{0}B_{0}^{-1})$.
-From the Taylor expansion of $r_{\theta}(x)$ around $\theta=\theta_{0}$, Lemma 3.2 immediately yields the following result:
-Theorem $3.$
-Suppose that Assumptions 1.1 and 3.1 hold.
-Then, it holds that $\sqrt{n}\left(\widehat{r}^{\mathrm{Lin}}_{n}(x)-r_{0}(x)\right)\xrightarrow{{\mathrm{d}}}{\mathcal{N}}(0,\kappa(x))$, where $\kappa(x)\coloneqq\left(\frac{\partial}{\partial\theta}r_{\theta_{0}}(x)\right)^{\top}B_{0}^{-1}\Sigma_{0}B_{0}^{-1}\left(\frac{\partial}{\partial\theta}r_{\theta_{0}}(x)\right)$.
+This section provides an estimation error analysis for $h_{0}$ estimated by the direct bias-correction term estimation method.
+We can use various models for $\mathcal{H}$}, including RKHS and neural networks.
+4.1 Model
+We define a model of the bias-correction term $h_{0}$ by $h(D,X)=\zeta^{-1}\circ f(D,X)=\zeta^{-1}(f(D,X))$, where $\zeta$ is a continuously differentiable and globally Lipschitz link function, and $f$ is some basic model.
+For example, if we use linear model for the bias-correction term $h_{0}$, we can write $h(D,X)=\Phi(D,X)^{\top}\beta$, where $\zeta$ is the identity function, $f(D,X)=\Phi(D,X)^{\top}\beta$, $\Phi$ is some basis function and $\beta$ is the corresponding parameter.
+If we use logistic model for the bias-correction term $h_{0}$, we can use logistic link for $\zeta$, and $f(D,X)=\Phi(X)^{\top}\beta$.
 4.2 RKHS
-Next, we investigate the case with RKHS regression.
-Let ${\mathcal{R}}^{\mathrm{RKHS}}$ be a class of RKHS functions, and define the estimator as $\widehat{r}^{\mathrm{RKHS}}_{n}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}^{\mathrm{RKHS}}}\widehat{{\mathcal{L}}}_{n}(r)+\lambda\|r\|^{2}_{{\mathcal{H}}}$, where $\|\cdot\|^{2}_{{\mathcal{H}}}$ is the RKHS norm.
-We analyze the estimation error by employing the results in Kanamori et al.
+First, we investigate the case with RKHS regression.
+Let ${\mathcal{F}}^{\mathrm{RKHS}}$ be a class of RKHS functions, and define $\widehat{f}^{\mathrm{RKHS}}_{n}\coloneqq\operatorname*{arg\,min}_{f\in{\mathcal{F}}^{\mathrm{RKHS}}}\widehat{{\mathcal{L}}}_{n}(\zeta^{-1}\circ f)+\lambda\|f\|^{2}_{{\mathcal{F}}}$, where $\|\cdot\|^{2}_{{\mathcal{F}}}$ is the RKHS norm.
+Then, we define an estimator as $h^{\mathrm{RKHS}}\coloneqq\zeta^{-1}\circ\widehat{f}^{\mathrm{RKHS}}_{n}$ We analyze the estimation error by employing the results in Kanamori et al.
 (2012), which study RKHS-based LSIF in DRE.
-We define the following localized class of RKHS functions as a technical device: ${\mathcal{R}}^{\text{RKHS}}_{M}\coloneqq\big\{r\in{\mathcal{R}}^{\text{RKHS}}\colon I(r)\leq M\big\}$.
+We define the following localized class of RKHS functions as a technical device: ${\mathcal{F}}^{\text{RKHS}}_{M}\coloneqq\big\{f\in{\mathcal{F}}^{\text{RKHS}}\colon I(f)\leq M\big\}$ for some norm $I(f)$ of $f$.
+We also define ${\mathcal{H}}^{\text{RKHS}}\coloneqq\left\{\zeta^{-1}\circ f\colon f\in{\mathcal{F}}^{\mathrm{RKHS}}\right\}$.
 We then make the following assumption using this localized class.
 Assumption 4.1.
-There exist constants $0<\gamma<2$, $0\leq\beta\leq 1$, $c_{0}>0$, and $A>0$ such that for all $M\geq 1$, it holds that $H_{B}(\delta,{\mathcal{R}}^{\text{RKHS}}_{M},P_{0})\leq A\left(\frac{M}{\delta}\right)^{\gamma}$, where $H_{B}(\delta,{\mathcal{R}}^{\text{RKHS}}_{M},P_{0})$ is the bracketing entropy with radius $\delta>0$ for the function class ${\mathcal{R}}^{\text{RKHS}}_{M}$ and the distribution $P_{0}$.
+There exist constants $0<\gamma<2$, $0\leq\beta\leq 1$, $c_{0}>0$, and $A>0$ such that for all $M\geq 1$, it holds that $H_{B}(\delta,{\mathcal{F}}^{\text{RKHS}}_{M},P_{0})\leq A\left(\frac{M}{\delta}\right)^{\gamma}$, where $H_{B}(\delta,{\mathcal{F}}^{\text{RKHS}}_{M},P_{0})$ is the bracketing entropy with radius $\delta>0$ for the function class ${\mathcal{F}}^{\text{RKHS}}_{M}$ and the distribution $P_{0}$.
 For the details of the definition of the bracketing entropy, see Appendix F and Definition 2.2 in van de Geer (2000).
@@
 Theorem 4.1 ($L_{2}$-norm estimation error bound).
-Suppose that Assumptions 1.1 and 3.4 hold.
+Suppose that $g$ is $\mu$-strongly convex and there exist constant $C>0$ such that $|g^{\prime\prime}(t)|\leq C\quad\forall t\in{\mathbb{R}}$.
+Assume also that $\zeta^{-1}(0)$ is finite.
+Suppose that Assumptions 1.1 and 4.1 hold.
 Set the regularization parameter $\lambda=\lambda_{n}$ so that $\lim_{n\to\infty}\lambda_{n}=0$ and $\lambda^{-1}_{n}=O(n^{1-\delta})$ ($n\to\infty$).
-If $r_{0}\in{\mathcal{R}}^{\mathrm{RKHS}}$, then we have $\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}=\left\|e_{0}(X)\left(\widehat{r}^{\mathrm{RKHS}}_{n}(1,X)-r_{0}(1,X)\right)\right\|^{2}_{P_{0}}+\left\|\big(1-e_{0}(X)\big)\left(\widehat{r}^{\mathrm{RKHS}}_{n}(0,X)-r_{0}(0,X)\right)\right\|^{2}_{P_{0}}=O_{P_{0}}\left(\lambda^{1/2}\right)$.
+If $h_{0}\in{\mathcal{H}}^{\mathrm{RKHS}}$, then we have $\left\|\widehat{h}^{\mathrm{RKHS}}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}=O_{P_{0}}\left(\lambda^{1/2}\right)$.
 The proof is provided in Appendix F, following the approach of Kanamori et al. (2012). The parameter $\gamma$ is determined by the function class to which $f_{0}$ belongs.
 4.3 Neural networks
-This section provides an estimation error analysis when we use neural networks for $\mathcal{R}$}. Our analysis is mostly based on Kato & Teshima (2021) and Zheng et al. (2022).
-Feedforward neural networks (FNNs). We define FNNs as follows:
-Definition 3.6 (Feedforward neural networks. From Zheng et al. (2022)).
+Second, we provide an estimation error analysis when we use neural networks for $\mathcal{H}$}.
+Our analysis is mostly based on Kato & Teshima (2021) and Zheng et al.
+(2022).
+We define Feedforward neural networks (FNNs) as follows:
+Definition 4.1 (FNNs. From Zheng et al. (2022)).
 Let $\mathcal{D}$}, $\mathcal{W}$}, $\mathcal{U}$}, and ${\mathcal{S}}\in(0,\infty)$ be parameters that can depend on $n$.
-Let ${\mathcal{F}}^{\mathrm{FNN}}\coloneqq{\mathcal{F}}^{\mathrm{FNN}}_{M,{\mathcal{D}},{\mathcal{W}},{\mathcal{U}},{\mathcal{S}}}$ be a class of ReLU-activated FNNs $r_{\theta}(1,x)=1/e_{\theta}(x)$ and $r_{\theta}(0,x)=1/(1-e_{\theta}(x))$, where $e_{\theta}\colon{\mathbb{R}}^{K}\to{\mathbb{R}}$ with parameter $\theta$, depth $\mathcal{D}$}, width $\mathcal{W}$}, size $\mathcal{S}$}, number of neurons $\mathcal{U}$}, satisfies the following conditions: (i) the number of hidden layers is $\mathcal{D}$}; (ii) the maximum width of the hidden layers is $\mathcal{W}$}; (iii) the number of neurons in $e_{\theta}$ is $\mathcal{U}$}; (iv) the total number of parameters in $e_{\theta}$ is $\mathcal{S}$}.
-Let $\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})$ be the pseudo-dimension of ${\mathcal{F}}^{\mathrm{FNN}}$. For the definition, see Anthony & Bartlett (1999) and Definition 3 in Zheng et al. (2022).
-We model $r_{0}$ by $r(1,X)=\frac{1}{e(x)}$ and $r(0,X)=\frac{1}{1-e(x)}$, where $e(x)=\frac{1}{1+\exp(-f(x))}$ for $f\in{\mathcal{F}}^{\mathrm{FNN}}$.
-Let us denote such an estimator by $r_{f}$ for $f\in{\mathcal{F}}^{\mathrm{FNN}}$.
-Let $f_{0}$ denote the true function such that $r_{f_{0}}=1+\exp(-f_{0}(x))$ holds.
-Then, the estimator is denoted as $\widehat{f}^{\mathrm{FNN}}_{n}\coloneqq\operatorname*{arg\,min}_{f\in{\mathcal{F}}^{\mathrm{FNN}}}\widehat{{\mathcal{L}}}_{n}(r_{f})$.
-Estimation error analysis. For the estimator, we can prove an estimation error bound. Let us make the following assumption.
+Let ${\mathcal{F}}^{\mathrm{FNN}}\coloneqq{\mathcal{F}}^{\mathrm{FNN}}_{M,{\mathcal{D}},{\mathcal{W}},{\mathcal{U}},{\mathcal{S}}}$ be a class of ReLU-activated FNNs with parameter $\theta$, depth $\mathcal{D}$}, width $\mathcal{W}$}, size $\mathcal{S}$}, number of neurons $\mathcal{U}$}, satisfies the following conditions: (i) the number of hidden layers is $\mathcal{D}$}; (ii) the maximum width of the hidden layers is $\mathcal{W}$}; (iii) the number of neurons in $e_{\theta}$ is $\mathcal{U}$}; (iv) the total number of parameters in $e_{\theta}$ is $\mathcal{S}$}.
+For the model ${\mathcal{F}}^{\mathrm{FNN}}$, we define $\widehat{f}^{\mathrm{FNN}}_{n}\coloneqq\operatorname*{arg\,min}_{f\in{\mathcal{F}}^{\mathrm{FNN}}}\widehat{{\mathcal{L}}}_{n}(\zeta^{-1}\circ f)$.
+Then, we define an estimator as $\widehat{h}^{\mathrm{FNN}}_{n}\coloneqq\zeta^{-1}\circ\widehat{f}^{\mathrm{FNN}}_{n}$.
+For the estimator, we can prove an estimation error bound. Let us make the following assumption.
 Assumption 4.2.
 There exists a constant $0<M<\infty$ such that $\|f_{0}\|_{\infty}<M$, and $\|f\|_{\infty}\leq M$ for any $f\in{\mathcal{F}}^{\mathrm{FNN}}$.
+Let $\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})$ be the pseudo-dimension of ${\mathcal{F}}^{\mathrm{FNN}}$.
+For the definition, see Anthony & Bartlett (1999) and Definition 3 in Zheng et al.
+(2022).
 Then, we prove the following estimation error bound:
 Theorem 4.2 (Estimation error bound for neural networks).
+Suppose that $g$ is $\mu$-strongly convex and there exist constant $C>0$ such that $|g^{\prime\prime}(t)|\leq M\quad\forall t\in{\mathbb{R}}$.
+Assume also that $\zeta^{-1}(0)$ is finite.
 Suppose that Assumption 4.2 holds.
-Also assume $r_{0}\in\Sigma(\beta,M,[0,1]^{d})$ with $\beta=k+a$, where $k\in{\mathbb{N}}^{+}$ and $a\in(0,1]$, and ${\mathcal{F}}^{\mathrm{FNN}}$ has width $\mathcal{W}$} and depth $\mathcal{D}$} such that ${\mathcal{W}}=38\big(\lfloor\beta\rfloor+1\big)^{2}d^{\lfloor\beta\rfloor+1}$ and ${\mathcal{D}}=21\big(\lfloor\beta\rfloor+1\big)^{2}\lceil n^{\frac{d}{2(d+2\beta)}}\log_{2}\left(8n^{\frac{d}{2(d+2\beta)}}\right)\rceil$.
-Then, for $M\geq 1$ and $n\leq\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})$, it holds that $\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}=\left\|e_{0}(X)\left(\widehat{r}^{\mathrm{FNN}}_{n}(1,X)-r_{0}(1,X)\right)\right\|^{2}_{P_{0}}+\left\|(1-e_{0}(X))\left(\widehat{r}^{\mathrm{FNN}}_{n}(0,X)-r_{0}(0,X)\right)\right\|^{2}_{P_{0}}=C_{0}\big(\lfloor\beta\rfloor+1\big)^{9}d^{2\lfloor\beta\rfloor+(\beta\land 3)}n^{-\frac{2\beta}{d+2\beta}}\log^{3}n$, where $C_{0}>0$ is a constant independent of $n$.
+For $f_{0}$ such that $h_{0}=\zeta^{-1}\circ f_{0}$, also assume $f_{0}\in\Sigma(\beta,M,[0,1]^{d})$ with $\beta=k+a$, where $k\in{\mathbb{N}}^{+}$ and $a\in(0,1]$, and ${\mathcal{F}}^{\mathrm{FNN}}$ has width $\mathcal{W}$} and depth $\mathcal{D}$} such that ${\mathcal{W}}=38\big(\lfloor\beta\rfloor+1\big)^{2}d^{\lfloor\beta\rfloor+1}$ and ${\mathcal{D}}=21\big(\lfloor\beta\rfloor+1\big)^{2}\lceil n^{\frac{d}{2(d+2\beta)}}\log_{2}\left(8n^{\frac{d}{2(d+2\beta)}}\right)\rceil$.
+Then, for $M\geq 1$ and $n\leq\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})$, it holds that $\left\|\widehat{h}^{\mathrm{FNN}}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}=C_{0}\big(\lfloor\beta\rfloor+1\big)^{9}d^{2\lfloor\beta\rfloor+(\beta\land 3)}n^{-\frac{2\beta}{d+2\beta}}\log^{3}n$, where $C_{0}>0$ is a constant independent of $n$.
 The proof is provided in Appendix G, following the approach of Zheng et al. (2022). This result directly implies the minimax optimality of the proposed method when $f_{0}$ belongs to a Hölder class.
@@
 We analyze the AIPW estimator with an estimated propensity score.
-Recall that the AIPW estimator is defined as $\widetilde{\tau}^{\mathrm{AIPW}}_{n}=\frac{1}{n}\sum^{n}_{i=1}\Big(h(X_{i},D_{i};\widehat{r}_{n})\left(Y_{i}-\widehat{\mu}_{n}(D_{i},X_{i})\right)+\widehat{\mu}_{n}(1,X_{i})-\widehat{\mu}_{n}(0,X_{i})\Big)$, which is also called the DR estimator.
+Recall that the AIPW estimator is defined as $\widetilde{\tau}^{\mathrm{AIPW}}_{n}=\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{h}_{n}(D_{i},X_{i})\left(Y_{i}-\widehat{\mu}_{n}(D_{i},X_{i})\right)+\widehat{\mu}_{n}(1,X_{i})-\widehat{\mu}_{n}(0,X_{i})\Big)$, which is also called the DR estimator.
 We first make the following assumption.
-Assumption 4.1 (Donsker condition).
-The hypothesis classes $\mathcal{R}$} and $\mathcal{M}$} belong to the Donsker class.
-For example, this assumption holds when the bracketing entropy is finite.
+Assumption 5.1 (Donsker condition or cross fitting).
+Either of the followings holds: (i) the hypothesis classes $\mathcal{H}$} and $\mathcal{M}$} belong to the Donsker class, or (ii) $\widehat{\mu}_{n}$ and $\widehat{h}_{n}$ are estimated via cross fitting.
+For example, the Donsker condition holds when the bracketing entropy of $\mathcal{H}$} is finite.
 In contrast, it is violated in high-dimensional regression or series regression settings where the model complexity diverges as $n\to\infty$.
@@
 However, if these quantities grow with the sample size, the assumption is no longer valid.
-Even if this assumption does not hold, we can still establish asymptotic normality by employing sample splitting (Klaassen, 1987).
-There are various ways to implement sample splitting, and one of the most well-known is cross-fitting, used in double machine learning (DML Chernozhukov et al., 2018, ,).
+Even if the Donsker condition does not hold, we can still establish asymptotic normality by employing sample splitting (Klaassen, 1987).
+There are various ways to implement sample splitting, and one of the most well-known is cross-fitting, used in double machine learning (DML, Chernozhukov et al., 2018).
 In DML, the dataset is split into several folds, and the nuisance parameters are estimated using only a subset of the folds.
-This ensures that in $h(X_{i},D_{i};\widehat{r}_{n})\left(Y_{i}-\widehat{\mu}_{n}(D_{i},X_{i})\right)+\widehat{\mu}_{n}(1,X_{i})-\widehat{\mu}_{n}(0,X_{i})$, the observations $(X_{i},D_{i},Y_{i})$ are not used to construct $\widehat{\mu}_{n}$ and $\widehat{r}_{n}$.
+This ensures that in $\widehat{h}_{n}(D_{i},X_{i})\left(Y_{i}-\widehat{\mu}_{n}(D_{i},X_{i})\right)+\widehat{\mu}_{n}(1,X_{i})-\widehat{\mu}_{n}(0,X_{i})$, the observations $(X_{i},D_{i},Y_{i})$ are not used to construct $\widehat{\mu}_{n}$ and $\widehat{r}_{n}$.
 For more details, see Chernozhukov et al.
@@
 Assumption 5.2 (Convergence rate).
-$\big\|\widehat{r}-r_{0}\big\|_{2}=o_{p}(1)$, $\big\|\widehat{\mu}-\mu_{0}\big\|_{2}=o_{p}(1)$, and $\big\|\widehat{r}-r_{0}\big\|_{2}\big\|\widehat{\mu}-\mu_{0}\big\|_{2}=o_{p}(1/\sqrt{n})$.
+$\big\|\widehat{h}-h_{0}\big\|_{2}=o_{p}(1)$, $\big\|\widehat{\mu}-\mu_{0}\big\|_{2}=o_{p}(1)$, and $\big\|\widehat{h}-h_{0}\big\|_{2}\big\|\widehat{\mu}-\mu_{0}\big\|_{2}=o_{p}(1/\sqrt{n})$.
 Under these assumptions, we show the asymptotic normality of $\widetilde{\tau}^{\mathrm{AIPW}}_{n}$. We omit the proof. For details, see Schuler & van der Laan (2024), for example.
 Theorem 5.1 (Asymptotic normality).
-Suppose that Assumptions 1.1, and 4.14.2 hold.
+Suppose that Assumptions 1.1, and 5.15.2 hold.
 Then, the AIPW estimator converges in distribution to a normal distribution as $\sqrt{n}\Big(\widetilde{\tau}^{\mathrm{AIPW}}_{n}-\tau_{0}\Big)\xrightarrow{{\mathrm{d}}}{\mathcal{N}}(0,V^{*})$, where $V^{*}$ is the efficiency bound defined as $V^{*}\coloneqq{\mathbb{E}}\left[\frac{\sigma^{2}(1,X)}{e_{0}(X)}+\frac{\sigma^{2}(0,X)}{1-e_{0}(X)}+\big(\tau_{0}(X)-\tau_{0}\big)^{2}\right]$ and $\tau_{0}(X)\coloneqq{\mathbb{E}}[Y(1)-Y(0)\mid X]$.
-Here, $V^{*}$ is the efficiency bound given by the variance of the efficient influence function; that is, $V^{*}={\mathbb{E}}\left[\Psi^{*}(X,D,Y)^{2}\right]$ holds (van der Vaart, 1998).
-Thus, this estimator is efficient.
-5 Generalization via the Bregman divergence minimization
-We can further generalize our direct bias-correction term estimation from the viewpoint of Bregman divergence minimization and point out the connection to Riesz regression and covariate balancing.
-5.1 Riesz regression and covariate balancing
-Our resulting objective function is the same as the one used in Riesz regression (Chernozhukov et al., 2022a, 2024).
-Although the motivation and derivation differ, we point out that Riesz regression can be interpreted as a specific instance of DRE.
-Building on this perspective, we generalize the objective function using Bregman divergence, following Sugiyama et al.
-(2011).
-Bregman divergence is a measure of discrepancy between two points, defined in terms of a strictly convex function (Bregman, 1967).
-In the context of DRE, Sugiyama et al.
-(2011) demonstrates that various existing methods can be formulated as Bregman divergence minimization problems.
-Inspired by this idea, we also extend the direct estimation method for the bias-correction term within the Bregman divergence minimization framework.
-This generalization further allows us to derive the empirical balancing method proposed in Chan et al.
-(2015), which aims to achieve covariate balance and known as an instance of tailored loss function (Zhao, 2019).
-These results suggest that Riesz regression and covariate balancing can be unified under the DRE framework.
-5.2 Bregman divergence minimization
-Let $\mathcal{G}$} be a set of functions $g\colon{\mathbb{R}}\to{\mathbb{R}}$ that is differentiable and strictly convex.
-Given $d\in\{1,0\}$, we define the Bregman divergence between $r(d,\cdot),r(d,\cdot)\colon{\mathcal{X}}\to(1,\infty)$ as
-$\mathrm{br}^{\dagger}_{g}\big(r_{0}(d,x)\mid r(d,x)\big)\coloneqq g(r_{0}(d,x))-g(r(d,x))-\partial g(r(d,x))(r_{0}(d,x)-r(d,x)),$
-where $\partial g$ denotes the derivative of $g$. Then, we define the average Bregman divergence as
-$\mathrm{BR}^{\dagger}_{g}\big(r_{0}\mid r\big)\coloneqq\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(g(r_{0}(d,X))-g(r(d,X))-\partial g(r(d,x))\Big(r_{0}(d,X)-r(d,X)\Big)\right)\Big].$
-We estimate $r_{0}$ by $r^{*}=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}^{\dagger}_{g}\big(r_{0}\mid r\big)$. By dropping the term that is irrelevant to learning, we have
-$r^{*}=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}_{g}\big(r\big),$
-where $\mathrm{BR}_{g}\big(r\big)\coloneqq\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\Big(-g(r(d,X))+\partial g(r(d,X))r(d,X)\Big)-\partial g(r(d,X))\Big]$. For the derivation, see Appendix B.
-Then, we estimate the bias-correction term $h_{0}$ (or $r_{0}$) by minimizing an empirical Bregman divergence as
-$\widehat{r}_{n}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\widehat{\mathrm{BR}}_{g}\big(r\big)+\lambda J(r),$
-where $\widehat{\mathrm{BR}}_{g}(r)\coloneqq\sum_{d\in\{1,0\}}\frac{1}{n}\sum^{n}_{i=1}\left(\mathbbm{1}[D_{i}=d]\Big(-g(r(d,X_{i}))+\partial g(r(d,X_{i}))r(d,X_{i})\Big)-\partial g(r(d,X_{i}))\right).$
-5.3 Least squares
-Our least squares method for direct bias-correction term estimation can be obtained by using $g(r)=r^{2}$.
-Under this choice of $g$, we obtain $\mathrm{BR}_{g}\big(r\big)=\sum_{d\in\{1,0\}}{\mathbb{E}}\left[-2r(d,X)+\mathbbm{1}[D=d]r(d,X)^{2}\right]$.
-This gives an objective that is the same as the one used in Chernozhukov et al.
-(2022a).
-5.4 Constrained maximum likelihood estimation
-By using different choices of $g$, we can derive various objective functions for direct bias-correction term estimation.
-As a specific case of Bregman divergence minimization, we introduce constrained maximum likelihood estimation.
-Consider $g^{\mathrm{L}}(r)=r\log r-r$, which is a convex function. By substituting this function into the Bregman divergence, we obtain
-$r^{*}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}_{g^{\mathrm{L}}}\big(r\big).$ (1)
-where $\mathrm{BR}_{g^{\mathrm{L}}}\big(r\big)\coloneqq{\mathbb{E}}\big[-\log(r(1,X))-\log(r(0,X))+\mathbbm{1}[D_{i}=1]r(1,X_{i})+\mathbbm{1}[D_{i}=0]r(0,X_{i})\big]$.
-Then, we estimate the bias-correction term as $\widehat{r}_{n}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\widehat{\mathrm{BR}}_{g^{\mathrm{L}}}\big(r\big)+\lambda J(r)$, where $\widehat{\mathrm{BR}}_{g^{\mathrm{L}}}(r)=\frac{1}{n}\sum^{n}_{i=1}\big(-\log\big(r(1,X_{i})\big)-\log\big(r(0,X_{i})\big)+\mathbbm{1}[D_{i}=1]r(1,X_{i})+\mathbbm{1}[D_{i}=0]r(0,X_{i})\big)$.
-This estimation method corresponds to unnormalized Kullback–Leibler (UKL) minimization in DRE (Sugiyama et al., 2011), which generalizes the KL importance estimation procedure (KLIEP).
-Note that solving (1) is equal to
-$r^{*}=\operatorname*{arg\,max}_{r\in{\mathcal{R}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\left[\log r(d,X)\right]\quad\mathrm{s.t.}~~~{\mathbb{E}}\big[\mathbbm{1}[D=1]r(1,X_{i})\big]={\mathbb{E}}\big[\mathbbm{1}[D=0]r(0,X_{i})\big]=1.$
-This technique is known as Silverman’s trick (Silverman, 1982).
-For details, see Theorem 3.3 in Kato et al.
-(2023).
-We can replace the expected values with the sample means and define the estimation problem as $\widehat{r}_{n}=\operatorname*{arg\,max}_{r\in{\mathcal{R}}}\frac{1}{n}\sum^{n}_{i=1}\sum_{d\in\{1,0\}}\log r(d,X_{i})\quad\mathrm{s.t.}~~~\frac{1}{n}\sum^{n}_{i=1}\mathbbm{1}[D_{i}=1]r(1,X_{i})=\frac{1}{n}\sum^{n}_{i=1}\mathbbm{1}[D_{i}=0]r(0,X_{i})=1$.
-5.5 Empirical balancing
-Next, we derive empirical balancing as a special case of Bregman divergence minimization (Chan et al., 2015).
-Empirical balancing is known as a specific form of covariate balancing and can be derived from a tailored loss function (Zhao, 2019).
-Let us consider $g^{\mathrm{E}}(r)=(r-1)\log\left(r-1\right)-r$, which is also convex for $r\in(1,\infty)$ and its derivative is $\partial g^{\mathrm{E}}(r)=\log\left(r-1\right)$.
-By substituting this function, we obtain $\mathrm{BR}_{g^{\mathrm{E}}}\big(r\big)\coloneqq\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\Big(\log\left(r-1\right)+r(d,X)\Big)-\log\left(r-1\right)\Big]$.
-Note that it holds that $\mathrm{BR}_{g^{\mathrm{E}}}\big(r\big)=$
-${\mathbb{E}}\Big[-\mathbbm{1}[D=0]\log\left(r(1,X)-1\right)-\mathbbm{1}[D=1]\log\left(r(0,X)-1\right)+\mathbbm{1}[D=1]r(1,X)+\mathbbm{1}[D=0]r(0,X)\Big].$
-Then, we estimate the bias-correction term as $\widehat{r}_{n}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\widehat{\mathrm{BR}}_{g^{\mathrm{E}}}\big(r\big)$, where the empirical Bregman divergence becomes $\widehat{\mathrm{BR}}_{g^{\mathrm{E}}}(r)=\frac{1}{n}\sum^{n}_{i=1}\big(\mathbbm{1}[D_{i}=0]\log\big(r(1,X_{i})-1\big)+\mathbbm{1}[D_{i}=1]\log\big(r(0,X_{i})-1\big)+\mathbbm{1}[D_{i}=1]r(1,X_{i})+\mathbbm{1}[D_{i}=0]r(0,X_{i})\big)$.
-If we model $r$ as $r(1,x)=1/e(x)$ and r(0,x)=1/(1e(x))r(0,x)=1/(1-e(x)), we have r(1,x)1=1/(r(0,x)1)r(1,x)-1=1/(r(0,x)-1).
-Therefore, we have $\widehat{\mathrm{BR}}_{g^{\mathrm{E}}}(r)=$
-$\frac{1}{n}\sum^{n}_{i=1}\left(\mathbbm{1}[D_{i}=1]\left(-\log\left(\frac{1}{r(1,X_{i})-1}\right)+r(1,X_{i})\right)+\mathbbm{1}[D_{i}=0]\left(-\log\left(\frac{1}{r(0,X_{i})-1}\right)+r(0,X_{i})\right)\right).$
-This objective function is equivalent to the one with the tailored loss proposed in Zhao (2019). From this objective, we can also derive empirical balancing (Chan et al., 2015).
-6 Discussion and related work
-6.1 Related work
-The estimation of the bias-correction term or the propensity score has been a core interest in causal inference.
-The bias-correction term can be interpreted as a gradient in a one-step estimator (van der Vaart, 2002).
-This idea was refined by van der Laan (2006) as targeted maximum likelihood estimation (TMLE).
-Chernozhukov et al.
-(2024) also explore a related topic from the viewpoint of Riesz representers and double machine learning (DML) (Chernozhukov et al., 2018).
-They propose automatic DML along with its implementation (Chernozhukov et al., 2022a).
-This topic has been addressed from various perspectives.
-One such approach focuses on estimating the propensity score by matching the distributions of the treatment and control groups (Chan et al., 2015; Deville & Särndal, 1992; Graham et al., 2012; Bryan S.
-Graham & Egel, 2016; Hellerstein & Imbens, 1999).
-For example, Imai & Ratkovic (2013) introduces the Covariate Balancing Propensity Score (CBPS), which estimates propensity scores by explicitly balancing covariate means.
-In parallel with, or subsequent to, Imai & Ratkovic (2013), several methods related to covariate balancing have been proposed (Hainmueller, 2012; Zubizarreta, 2015; Athey et al., 2018).
-From a methodological perspective, our study is inspired by DRE (Sugiyama et al., 2012). We review the literature of DRE in Appendix A.
+Here, $V^{*}$ matches the efficiency bound given as the variance of the efficient influence function (van der Vaart, 1998). Thus, this estimator is efficient.
 Table 2: Experimental results. We report the empirical MSE and Bias of each method.
-Data Dimension DM DBC (LS) DBC (UKL) Logistic CBPS RieszNet DM
+Data Dimension DM DBC (LS) DBC (KL) MLE CBPS RieszNet DM
 Three-layer perceptron Dragonnet Linear model
@@
 Model 2 $K=3$ MSE 0.048 0.343 0.033 0.819 0.037 2.838 0.045 1.848 0.044 0.030 0.034 0.051 2.866
-Bias $K=3$ -0.009 -0.275 -0.011 -0.382 -0.010 -0.403 -0.011 -0.781 -0.012 -0.022 -0.020 -0.057 -0.214
+$K=3$ Bias -0.009 -0.275 -0.011 -0.382 -0.010 -0.403 -0.011 -0.781 -0.012 -0.022 -0.020 -0.057 -0.214
 $K=3$ MSE 0.517 2.006 0.474 2.980 0.477 6.517 0.507 3.816 0.512 0.407 0.446 0.424 7.482
@@
 6 Simulation studies
-We assess the empirical performance of our method through simulation studies, evaluating ATE estimation error across a range of scenarios.
-We compare our approach with CBPS (Imai & Ratkovic, 2013) and RieszNet (Chernozhukov et al., 2022a).
-Because our least squares is equivalent to Resz regression, we include RieszNet primarily as a numerical check of equivalence, noting architectural differences.
+We assess the performance of our method through simulation studies, evaluating ATE estimation error.
+We denote our direct bias-correction term estimation methods as DBC (LS) when using the squared loss, and DBC (TL) when using the tailored loss.
+We compare our approach with ATE estimators using propensity score estimated by maximum likelihood estimation (MLE), CBPS (Imai & Ratkovic, 2013a), and RieszNet (Chernozhukov et al., 2022a).
+Because our DBC (LS) is equivalent to Resz regression, we include RieszNet primarily as a numerical check of equivalence, noting architectural differences.
+In this section, for simplicity, we do not apply cross-fitting.
+We also conduct experiments in Appendices I and J using synthetic and semi-synthetic data, respectively, in which we apply cross-fitting.
 We consider two different dimensions for $X$, setting $K=3$ and $K=10$, and two different outcome models.
@@
 In Model 2, the outcome is generated as $Y_{i}=X^{\top}_{i}\beta+\left(X^{\top}_{i}\beta\right)^{2}+3\sin(X_{i,1})+1.1+\tau_{0}D_{i}+\varepsilon_{i}$.
-For propensity score estimation, we employ a three-layer neural network with an Exponential Linear Unit (ELU) activation function for each hidden layer ($100$ nodes per layer).
+We model $h_{0}$ by modeling $e_{0}$.
+To model $e_{0}$, we use a three-layer neural network with an Exponential Linear Unit (ELU) activation function for each hidden layer ($100$ nodes per layer).
 The final output layer applies a sigmoid function to ensure that the estimated propensity scores remain in $(0,1)$.
@@
 For RieszNet, we adopt the DragonNet architecture proposed in Shi et al.
-(2019), following the original implementation by Chernozhukov et al.
-(2022a).
-For each propensity score estimation method, including ours, we compute both the IPW and AIPW estimators using the estimated scores.
+(2019), following Chernozhukov et al.
+(2022b).
+For each method, including ours, we compute both the IPW and AIPW estimators using the estimated scores.
 Additionally, we include the direct method (DM) estimator with neural networks for comparison.
@@
 Finally, we present numerical experiments that demonstrate the effectiveness of our approach in practical scenarios, including importance weighting and outlier detection.
-Appendix B Bregman divergence
-Here, we show the equivalence between
-$r^{*}=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}^{\dagger}_{g}\big(r_{0}\mid r\big),$
-where $\mathrm{BR}^{\dagger}_{g}\big(r_{0}\mid r\big)\coloneqq\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(g(r_{0}(d,X))-g(r(d,X))-\partial g(r(d,x))\Big(r_{0}(d,X)-r(d,X)\Big)\right)\Big]$, and
-$r^{*}=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}_{g}\big(r\big),$
-where $\mathrm{BR}_{g}\big(r\big)=\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\Big(-g(r(d,X))+\partial g(r(d,X))r(d,X)\Big)-\partial g(r(d,X))\Big]$.
-This can be shown as follows:
-$r^{*}=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}^{\dagger}_{g}\big(r_{0}\mid r\big)$
-$=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(g(r_{0}(d,X))-g(r(d,X))-\partial g(r(d,x))\Big(r_{0}(d,X)-r(d,X)\Big)\right)\Big]$
-$=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(-g(r(d,X))-\partial g(r(d,x))\Big(r_{0}(d,X)-r(d,X)\Big)\right)\Big]$
-$=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(-g(r(d,X))-\partial g(r(d,x))r(d,X)\right)\Big]-{\mathbb{E}}\Big[\mathbbm{1}[D=d]\partial g(r(d,x))r_{0}(d,X)\Big]$
-$=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(-g(r(d,X))-\partial g(r(d,x))r(d,X)\right)\Big]-{\mathbb{E}}\Big[\partial g(r(d,x))\Big].$
-Here, we used
-${\mathbb{E}}[\mathbbm{1}[D=1]r_{0}(1,X)\mid X]={\mathbb{E}}[e_{0}(X)r_{0}(1,X)\mid X]=1.$
-We can show the same result for the case with $D=0$.
+Appendix B Silverman’s trick
+Note that minimization of the Bregman divergence with the KL divergence loss is equal to
+$r^{*}=\operatorname*{arg\,max}_{r\in{\mathcal{R}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\left[\log r(d,X)\right]\quad\mathrm{s.t.}~~~{\mathbb{E}}\big[\mathbbm{1}[D=1]r(1,X_{i})\big]={\mathbb{E}}\big[\mathbbm{1}[D=0]r(0,X_{i})\big]=1.$
+This technique is known as Silverman’s trick (Silverman, 1982).
+For details, see Theorem 3.3 in Kato et al.
+(2023).
+We can replace the expected values with the sample means and define the estimation problem as $\widehat{r}_{n}=\operatorname*{arg\,max}_{r\in{\mathcal{R}}}\frac{1}{n}\sum^{n}_{i=1}\sum_{d\in\{1,0\}}\log r(d,X_{i})\quad\mathrm{s.t.}~~~\frac{1}{n}\sum^{n}_{i=1}\mathbbm{1}[D_{i}=1]r(1,X_{i})=\frac{1}{n}\sum^{n}_{i=1}\mathbbm{1}[D_{i}=0]r(0,X_{i})=1$.
 Appendix C Estimation of the average treatment effect for the treated (ATT)
@@
 Thus, the bias-correction term for ATT estimation is given as
-$\widetilde{h}_{0}(D,X,e_{0},\pi_{0})\coloneqq\frac{\mathbbm{1}[D=1]}{\pi_{0}}-\frac{e_{0}(X)\mathbbm{1}[D=0]}{\pi_{0}(1-e_{0}(X))},$
+$\widetilde{h}_{0}(D,X)\coloneqq\frac{\mathbbm{1}[D=1]}{\pi_{0}}-\frac{e_{0}(X)\mathbbm{1}[D=0]}{\pi_{0}(1-e_{0}(X))},$
 where $\pi_{0}={\mathbb{E}}[\mathbbm{1}[D=1]]$.
-Let $w_{0}(x)\coloneqq\frac{e_{0}(X)\mathbbm{1}[D=0]}{(1-e_{0}(X))}$. Then, we denote the bias-correction term as
-$\widetilde{h}_{0}(D,X,w_{0},\pi_{0})\coloneqq\frac{\mathbbm{1}[D=1]}{\pi_{0}}-\frac{w_{0}(X)\mathbbm{1}[D=0]}{\pi_{0}}.$
+Let $w_{0}(x)\coloneqq\frac{e_{0}(X)}{(1-e_{0}(X))}$. Then, we denote the bias-correction term as
+$\widetilde{h}_{0}(D,X)\coloneqq\frac{\mathbbm{1}[D=1]}{\pi_{0}}-\frac{w_{0}(X)\mathbbm{1}[D=0]}{\pi_{0}}.$
 Let $\mathcal{W}$} be a set of functions $w\colon{\mathcal{X}}\to{\mathbb{R}}_{+}$. Then, we define the following least squares:
@@
 Note that we use $\pi_{0}$ itself. We can show that this least squares is equivalent to
-$w^{*}=\operatorname*{arg\,min}_{r\in\mathcal{R}}\left\{-2{\mathbb{E}}_{1}\left[w(X)\right]+{\mathbb{E}}\left[w(X)^{2}\mathbbm{1}[D=0]\right]\right\},$
+$w^{*}=\operatorname*{arg\,min}_{r\in\mathcal{R}}\left\{-2{\mathbb{E}}\left[w(X)\right]+{\mathbb{E}}\left[w(X)^{2}\mathbbm{1}[D=0]\right]\right\},$
 where ${\mathbb{E}}_{1}$ is expectation over the treated group ($p(x\mid d=1)$). The empirical version of this risk is given as
@@
 $={\mathbb{E}}\left[e_{0}(X)w(X)/\pi_{0}\right]$
-$=\int\frac{1}{\pi_{0}}e_{0}(x)w(x)\pi_{0}(x){\mathrm{d}}x$
+$=\int\frac{1}{\pi_{0}}e_{0}(x)w(x)p_{0}(x){\mathrm{d}}x$
 $=\int\frac{1}{\pi_{0}}\frac{\pi_{0}p_{0}(x\mid d=1)}{p_{0}(x)}w(x)p_{0}(x){\mathrm{d}}x$
@@
 Let $\sigma_{1},\dots,\sigma_{n}$ be $n$ independent Rademacher random variables; that is, independent random variables for which $P(\sigma_{i}=1)=P(\sigma_{i}=-1)=1/2$. Let us define
-$\mathfrak{R}_{n}f\coloneqq\frac{1}{n}\sum^{n}_{i=1}\sigma_{i}f(X_{i}).$
+$\mathfrak{R}_{n}f\coloneqq\frac{1}{n}\sum^{n}_{i=1}\sigma_{i}f(W_{i}).$
 Additionally, given a class $\mathcal{F}$}, we define
@@
 Let $\mathcal{F}$} be a class of functions that map $\mathcal{X}$} into $[a,b]$. For $f\in{\mathcal{F}}$, let us define
-$Pf\coloneqq{\mathbb{E}}[f(X)],$
-$P_{n}f\coloneqq\frac{1}{n}\sum^{n}_{i=1}f(X_{i}).$
+$Pf\coloneqq{\mathbb{E}}[f(W)],$
+$P_{n}f\coloneqq\frac{1}{n}\sum^{n}_{i=1}f(W_{i}).$
 We introduce the following result about the Rademacher complexity.
@@
 Let $\mathcal{F}$} be a class of functions that map $\mathcal{X}$} into $[a,b]$.
-Assume that there is some $r>0$ such that for every $f\in{\mathcal{F}}$, $\mathrm{Var}(f(X))\leq r$.
+Assume that there is some $r>0$ such that for every $f\in{\mathcal{F}}$, $\mathrm{Var}(f(W))\leq r$.
 Then, for every $z>0$, with probability at least $1-\exp(-z)$, it holds that
@@
 Appendix E Basic inequalities
+E.1 Strong convexity
+Lemma E.1 ($L_{2}$ distance bound from Lemma 4 in Kato & Teshima (2021)).
+If $\inf_{h\in(-\infty),\infty}g^{\prime\prime}(h)>0$, then there exists $\mu>0$ such that for all $h\in{\mathcal{H}}$,
+$\|h-h_{0}\|^{2}_{2}\leq\frac{2}{\mu}\Big(\mathrm{BR}_{g}(h)-\mathrm{BR}_{g}(h_{0})\Big)$
+holds.
+From the strong convexity and Lemma E.1, we have
+$\frac{\mu}{2}\|\widehat{h}_{n}-h_{0}\|^{2}_{2}\leq\mathrm{BR}_{g}(\widehat{h}_{n})-\mathrm{BR}_{g}\big(h_{0}\big).$
 Recall that we have defined an estimator $\widehat{r}$ as follows:
-$\widehat{r}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\widehat{{\mathcal{L}}}_{n}(r)+\lambda J(r),$
-where $J(r)$ is some regularization term.
-Throughout the proof, we use the following basic inequalities that hold for $\widehat{r}$.
+$\widehat{h}\coloneqq\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\widehat{{\mathcal{L}}}_{n}(h)+\lambda J(h),$
+where $J\left(h\right)$ is some regularization term.
+E.2 Preliminary
 Proposition E.2.
 The estimator $\widehat{r}$ satisfies the following inequality:
-[…]
-$\leq\frac{1}{n}\sum^{n}_{i=1}\Big(-2r_{0}(1,X_{i})-2r_{0}(0,X_{i})+\mathbbm{1}[D_{i}=1]r_{0}(1,X_{i})^{2}+\mathbbm{1}[D_{i}=0]r_{0}(0,X_{i})^{2}\Big)+\lambda J(r_{0}).$
-For a function $f\colon{\mathcal{X}}\to{\mathbb{R}}$ and $X$ following $P$, let us denote the sample mean and expectation as
-$\int f{\mathrm{d}}P_{n}$ $\coloneqq\int f(\cdot){\mathrm{d}}P_{n}\coloneqq\frac{1}{n}\sum^{n}_{i=1}f(X_{i}),$
-$\int f{\mathrm{d}}P$ $\coloneqq\int f(\cdot){\mathrm{d}}P\coloneqq{\mathbb{E}}[f(X)].$
-Following this notation, we also express this inequality as
-$\int-2\big(\widehat{r}_{n}(1,\cdot)-r_{0}(1,\cdot)\big){\mathrm{d}}P_{n}+\int-2\big(\widehat{r}_{n}(0,\cdot)-r_{0}(0,\cdot)\big){\mathrm{d}}P_{n}$
-$\ \ \ +\int\mathbbm{1}[D_{i}=1]\Big(\widehat{r}_{n}(1,\cdot)^{2}-r_{0}(1,\cdot)^{2}\Big){\mathrm{d}}P_{n}+\int\mathbbm{1}[D_{i}=0]\Big(\widehat{r}_{n}(0,\cdot)^{2}-r_{0}(0,X)^{2}\Big){\mathrm{d}}P_{n}$
-$\ \ \ +\lambda J(\widehat{r})-\lambda J(r_{0})\leq 0.$
+$\widehat{\mathrm{BR}}_{g}(\widehat{h})+\lambda J(\widehat{h})\leq\widehat{\mathrm{BR}}_{g}(h^{*})+\lambda J(h^{*}),$
+where recall that
+$\widehat{\mathrm{BR}}_{g}(h)\coloneqq\frac{1}{n}\sum^{n}_{i=1}\Big(-g(h(D_{i},X_{i}))+\partial g(h(D_{i},X_{i}))h(D_{i},X_{i})-\partial g(h(1,X_{i}))-\partial g(h(0,X_{i}))\Big).$
+Let $Z\in{\mathcal{Z}}$ be a random variable with a space $\mathcal{Z}$}, and $\{Z_{i}\}^{n}_{i=1}$ be its realizations.
+For a function $f\colon{\mathcal{Z}}\to{\mathbb{R}}$ and $X$ following $P$, let us denote the sample mean as
+$\widehat{{\mathbb{E}}}[f(Z)]\coloneqq\frac{1}{n}\sum^{n}_{i=1}f(Z_{i}).$
+We also denote $\widehat{{\mathbb{E}}}[f(Z)]-{\mathbb{E}}[f(Z)]=(\widehat{{\mathbb{E}}}-{\mathbb{E}})f(Z)$
+E.3 Risk bound
+Recall that
+$\widehat{\mathrm{BR}}_{g}(h)=\frac{1}{n}\sum^{n}_{i=1}\Big(-g(h(D_{i},X_{i}))+\partial g(h(D_{i},X_{i}))h(D_{i},X_{i})-\partial g(h(1,X_{i}))-\partial g(h(0,X_{i}))\Big).$
+Let us define
+$L(h,D,X)\coloneqq-g(h(D,X))+\partial g(h(D,X))h(D,X)-\partial g(h(1,X))-\partial g(h(0,X)),$
+and we can write
+$\widehat{\mathrm{BR}}_{g}(h)=\widehat{{\mathbb{E}}}\big[L(h,D,X)\big]$
+Then, from Proposition E.2, we have
+$\widehat{{\mathbb{E}}}\big[L(h^{*},D,X)\big]-\widehat{{\mathbb{E}}}\big[L(\widehat{h}_{n},D,X)\big]+\lambda J(\widehat{h})-\lambda J(h^{*})\geq 0.$
+Throughout the proof, we use the following basic inequalities that hold for $\widehat{h}$.
 Proposition E.3.
 The estimator $\widehat{r}$ satisfies the following inequality:
-$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}$
-$\leq\sum_{d\in\{1,0\}}\Bigg(\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})$
-$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\int e_{0}(d,\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\Bigg)$
-$\ \ \ +\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{n}+\int e_{0}(d\mid X)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}\Bigg)$
+$\frac{\mu}{2}\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}$
 […]
@@
 Proof.
-Let $e_{0}(1\mid x)=e_{0}(x)$ and $e_{0}(0\mid x)=1-e_{0}(x)$. Then, the following holds for the $L_{2}$ estimation error:
-$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}$
-$=\left\|\mathbbm{1}[D=1]\big(\widehat{r}(1,X)-r_{0}(1,X)\big)-\mathbbm{1}[D=0]\big(\widehat{r}(0,X)-r_{0}(0,X)\big)\right\|^{2}_{P_{0}}$
-$=\left\|e_{0}(1\mid\cdot)\big(\widehat{r}(1,X)-r_{0}(1,X)\big)\right\|^{2}_{P_{0}}+\left\|e_{0}(1\mid 0)\big(\widehat{r}(0,X)-r_{0}(0,X)\big)\right\|^{2}_{P_{0}}$
-$=\int-2\big(\widehat{r}_{n}(1,\cdot)-r_{0}(1,\cdot)\big){\mathrm{d}}P_{0}+\int-2\big(\widehat{r}_{n}(0,\cdot)-r_{0}(0,\cdot)\big){\mathrm{d}}P_{0}$
-$\ \ \ +\int e_{0}(1\mid\cdot)\Big(\widehat{r}_{n}(1,\cdot)^{2}-r_{0}(1,\cdot)^{2}\Big){\mathrm{d}}P_{0}+\int e_{0}(0,\cdot)\Big(\widehat{r}_{n}(0,\cdot)^{2}-r_{0}(0,\cdot)^{2}\Big){\mathrm{d}}P_{0}.$
+From the strong convexity and Lemma E.1, we have
+$\frac{\mu}{2}\|\widehat{h}_{n}-h_{0}\|^{2}_{2}\leq\mathrm{BR}_{g}(\widehat{h}_{n})-\mathrm{BR}_{g}\big(h_{0}\big)={\mathbb{E}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right].$
 From Proposition E.2, we have
-$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}$
-$\leq\int-2\big(\widehat{r}_{n}(1,\cdot)-r_{0}(1,\cdot)\big){\mathrm{d}}P_{0}+\int-2\big(\widehat{r}_{n}(0,\cdot)-r_{0}(0,\cdot)\big){\mathrm{d}}P_{0}$
-$\ \ \ +\int e_{0}(1\mid\cdot)\Big(\widehat{r}_{n}(1,\cdot)^{2}-r_{0}(1,\cdot)^{2}\Big){\mathrm{d}}P_{0}+\int e_{0}(0,\cdot)\Big(\widehat{r}_{n}(0,\cdot)^{2}-r_{0}(0,\cdot)^{2}\Big){\mathrm{d}}P_{0}$
-$\ \ \ -\int-2\big(\widehat{r}_{n}(1,\cdot)-r^{*}(1,\cdot)\big){\mathrm{d}}P_{n}-\int-2\big(\widehat{r}_{n}(0,\cdot)-r^{*}(0,\cdot)\big){\mathrm{d}}P_{n}$
-$\ \ \ -\int e_{0}(1\mid\cdot)\Big(\widehat{r}_{n}(1,\cdot)^{2}-r^{*}(1,\cdot)^{2}\Big){\mathrm{d}}P_{n}-\int e_{0}(0,\cdot)\Big(\widehat{r}_{n}(0,\cdot)^{2}-r^{*}(0,\cdot)^{2}\Big){\mathrm{d}}P_{n}$
-$\ \ \ -\lambda J(\widehat{r})+\lambda J(r_{0})$
-$=\int-2\big(\widehat{r}_{n}(1,\cdot)-r_{0}(1,\cdot)\big){\mathrm{d}}P_{0}+\int-2\big(\widehat{r}_{n}(0,\cdot)-r_{0}(0,\cdot)\big){\mathrm{d}}P_{0}$
-$\ \ \ +\int e_{0}(1\mid\cdot)\Big(\widehat{r}_{n}(1,\cdot)^{2}-r_{0}(1,\cdot)^{2}\Big){\mathrm{d}}P_{0}+\int e_{0}(0,\cdot)\Big(\widehat{r}_{n}(0,\cdot)^{2}-r_{0}(0,\cdot)^{2}\Big){\mathrm{d}}P_{0}$
-$\ \ \ -\int-2\big(\widehat{r}_{n}(1,\cdot)-r^{*}(1,\cdot)\big){\mathrm{d}}P_{n}-\int-2\big(\widehat{r}_{n}(0,\cdot)-r^{*}(0,\cdot)\big){\mathrm{d}}P_{n}$
-$\ \ \ -\int e_{0}(1\mid\cdot)\Big(\widehat{r}_{n}(1,\cdot)^{2}-r^{*}(1,\cdot)^{2}\Big){\mathrm{d}}P_{n}-\int e_{0}(0,\cdot)\Big(\widehat{r}_{n}(0,\cdot)^{2}-r^{*}(0,\cdot)^{2}\Big){\mathrm{d}}P_{n}$
-$\ \ \ +\int-2\big(r_{0}(1,\cdot)-r_{0}(1,\cdot)\big){\mathrm{d}}P_{n}+\int-2\big(\widehat{r}_{n}(0,\cdot)-r^{*}(0,\cdot)\big){\mathrm{d}}P_{n}$
-$\ \ \ +\int e_{0}(1\mid\cdot)\Big(r_{0}(1,\cdot)^{2}-r_{0}(1,\cdot)^{2}\Big){\mathrm{d}}P_{n}+\int e_{0}(0,\cdot)\Big(r_{0}(0,\cdot)^{2}-r_{0}(0,\cdot)^{2}\Big){\mathrm{d}}P_{n}$
-$\ \ \ -\lambda J(\widehat{r})+\lambda J(r_{0}).$
-Here, we used
-$\int\mathbbm{1}[d=1]r_{0}(1,\cdot)^{2}{\mathrm{d}}P_{0}$ $=2\int\mathbbm{1}[d=1]r_{0}(1,\cdot)^{2}{\mathrm{d}}P_{0}-\int\mathbbm{1}[d=1]r_{0}(1,\cdot)^{2}{\mathrm{d}}P_{0}$
-$=2\int r_{0}(1,\cdot){\mathrm{d}}P_{0}-\int\mathbbm{1}[d=1]r_{0}(1,\cdot)^{2}{\mathrm{d}}P_{0}.$
+$\frac{\mu}{2}\left\|\widehat{h}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}$
+$\leq{\mathbb{E}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
+$={\mathbb{E}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
+$\ \ \ \ \ \ -\widehat{{\mathbb{E}}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
+$\ \ \ \ \ \ +\widehat{{\mathbb{E}}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
+$\leq{\mathbb{E}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
+$\ \ \ \ \ \ -\widehat{{\mathbb{E}}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
+$\ \ \ \ \ \ +\widehat{{\mathbb{E}}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
+$\ \ \ \ \ \ -\widehat{{\mathbb{E}}}\left[L(\widehat{h}_{n},D,X)-L(h^{*},D,X)\right]+\lambda J(\widehat{h})-\lambda J(h_{0}).$
 Appendix F Proof of Theorem 4.1
 We show Theorem 4.1 by bounding
-$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})$ (2)
-$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\int e_{0}(d,\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\Bigg)$
-$\ \ \ +\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{n}+\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}\Bigg)$ (3)
-in Proposition E.2.
-Here, since $r_{0}\in{\mathcal{H}}$, it holds that $r^{*}=r_{0}$, which implies that
-$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{n}+\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}\Bigg)=0.$
-Therefore, we consider bounding the first sum
-$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})+\int e_{0}(d,\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\Bigg).$
-We can bound this term by using the empirical-process arguments.
+$\left({\mathbb{E}}-\widehat{{\mathbb{E}}}\right)\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right],$ (1)
+in Proposition E.3. We can bound this term by using the empirical-process arguments.
+Note that since $h_{0}\in{\mathcal{H}}$, it holds that $h^{*}=h_{0}$, which implies that
 F.1 Preliminary
-The following proposition is originally presented in van de Geer (2000) and was rephrased by Kanamori et al. (2012) in a form that is convenient for the theoretical analysis in DRE.
-Lemma F.1 (Lemma 5.13 in van de Geer (2000), Proposition 1 in Kanamori et al. (2012)).
+We introduce the following propositions from van de Geer (2000), Kanamori et al. (2012) and Kato & Teshima (2021).
+Definition F.1 (Derived function class and bracketing entropy (from Definition 4 in Kato & Teshima (2021))).
+Given a real-valued function class $\mathcal{F}$}, define $\ell\circ{\mathcal{F}}:=\{\ell\circ f\colon f\in{\mathcal{F}}\}$.
+By extension, we define $I:\ell\circ{\mathcal{H}}\to[1,\infty)$ by $I(\ell\circ h)=I(h)$ and $\ell\circ{\mathcal{H}}_{M}:=\{\ell\circ h:h\in{\mathcal{H}}_{M}\}$.
+Note that, as a result, $\ell\circ{\mathcal{H}}_{M}$ coincides with $\{\ell\circ h\in\ell\circ{\mathcal{H}}:I(\ell\circ h)\leq M\}$.
+Proposition F.1.
+Let $\ell:{\mathbb{R}}\to{\mathbb{R}}$ be a $v$-Lipschitz continuous function.
+Let $H_{B}\big(\delta,{\mathcal{F}},\|\cdot\|_{L_{2}(P_{0})}\big)$ denote the bracketing entropy of $\mathcal{F}$} with respect to a distribution $P$.
+Then, for any distribution $P$, any $\gamma>0$, any $M\geq 1$, and any $\delta>0$, we have
+$H_{B}\big(\delta,\ell\circ{\mathcal{H}},\|\cdot\|_{L_{2}(P_{0})}\big)$ $\leq\frac{(s+1)(2v)^{\gamma}}{\gamma}\left(\frac{M}{\delta}\right)^{\gamma}.$
+Moreover, there exists $M>0$ such that for any $M\geq 1$ and any distribution $P$,
+$\sup_{\ell\circ h\in\ell\circ{\mathcal{H}}_{M}}\|\ell\circ h-\ell\circ h^{*}\|_{L_{2}(P_{0})}$ $\leq c_{0}vM,$
+$\sup_{\stackrel{{\scriptstyle\ell\circ h\in\ell\circ{\mathcal{H}}_{M}}}{{\|\ell\circ h-\ell\circ h^{*}\|_{L_{2}(P_{0})}\leq\delta}}}\|\ell\circ h-\ell\circ h^{*}\|_{\infty}$ $\leq c_{0}vM,\quad\text{for all }\delta>0.$
+Proposition F.2 (Lemma 5.13 in van de Geer (2000), Proposition 1 in Kanamori et al. (2012)).
 Let ${\mathcal{F}}\subset L^{2}(P)$ be a function class and the map $I(f)$ be a complexity measure of $f\in{\mathcal{F}}$, where $I$ is a non-negative function on $\mathcal{F}$} and $I(f_{0})<\infty$ for a fixed $f_{0}\in{\mathcal{F}}$.
@@
 $\sup_{f\in{\mathcal{F}}_{M}}\|f-f_{0}\|\leq c_{0}M,\ \sup_{\stackrel{{\scriptstyle f\in{\mathcal{F}}_{M}}}{{\|f-f_{0}\|_{L^{2}(P)}\leq\delta}}}\|f-f_{0}\|_{\infty}\leq c_{0}M,\quad\text{for all }\delta>0,$
-and that $H_{B}(\delta,\mathcal{F}_{M},P)=O{M/\delta}^{\gamma}$. Then, we have
+and that $H_{B}(\delta,{\mathcal{F}}_{M},P)=O\left((M/\delta)^{\gamma}\right)$. Then, we have
 $\sup_{f\in{\mathcal{F}}}\frac{\left|\int(f-f_{0})d(P-P_{n})\right|}{D(f)}=O_{p}(1),\ (n\to\infty),$
@@
 $D(f)=\max{\frac{\|f-f_{0}\|_{L^{2}(P)}^{1-\gamma/2}I(f)^{\gamma/2}}{\sqrt{n}}}{\frac{I(f)}{n^{2/(2+\gamma)}}}.$
+Proposition F.3.
+Let $g\colon{\mathcal{K}}\to{\mathbb{R}}$ be twice continuously differentiable and strictly convex for the space $\mathcal{K}$} of $h_{0}$, and suppose that there exists $M>0$ such that
+$|g^{\prime\prime}(t)|\leq M\quad\text{for all }t\in{\mathbb{R}}.$
+Let $\zeta^{-1}\colon{\mathbb{R}}\to{\mathbb{R}}$ be continuously differentiable and globally Lipschitz, that is, there exists $L_{\zeta}>0$ such that
+$|\zeta^{-1}(s)-\zeta^{-1}(t)|\leq L_{\zeta}|s-t|\quad\text{for all }s,t\in{\mathbb{R}}.$
+Assume also that $\zeta^{-1}(0)$ is finite, and define
+$a_{0}:=|\zeta^{-1}(0)|,\qquad a_{1}:=L_{\zeta},$
+so that
+$|\zeta^{-1}(u)|\leq a_{0}+a_{1}|u|\quad\text{for all }u\in{\mathbb{R}}.$
+Let $h$ be a bounded real-valued function on the domain of $(D,X)$, and write
+$\|h\|_{\infty}:=\sup_{d,x}|h(d,x)|.$
+Let $L$ be a linear functional acting on bounded functions, such that for some constant $C_{L}>0$,
+$|L(f)|\leq C_{L}\bigl(1+\|f\|_{\infty}\bigr)\quad\text{for all bounded }f.$
+Define
+$L(\zeta^{-1}\circ f)$ $=g\bigl(\zeta^{-1}\circ f(D,X)\bigr)+\partial g\bigl(\zeta^{-1}\circ f(D,X)\bigr)\,\zeta^{-1}\circ h(D,X)$
+$\quad-\partial g\bigl(\zeta^{-1}\circ f(1,X)\bigr)-\partial g\bigl(\zeta^{-1}\circ f(0,X)\bigr).$
+Then there exists a constant $C>0$ (depending only on $g$, $\zeta^{-1}$ and $C_{L}$) such that
+$|L(\zeta^{-1}\circ f)|\leq C\bigl(1+\|f\|_{\infty}^{2}\bigr).$
 F.2 Upper bound using the empirical-process arguments
+From Propositions F.1F.3, we obtain the following result.
 Proposition F.4.
-[From Lemma 2 in Kanamori et al. (2012)] Under the conditions of Theorem 3.5, for any $0<\gamma<2$, we have
-$\left|\int\left(\widehat{r}(d,\cdot)-r^{*}(d,\cdot))({\mathrm{d}}(P_{0}-P_{n})\right)\right|$ $=O_{p}\left(\max\left\{\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left\|\widehat{r}(d,\cdot)\right\|^{\gamma/2}_{{\mathcal{H}}}}{\sqrt{n}},\frac{\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}}{n^{2/(2+\gamma)}}\right\}\right),$
-$\left|\int\left(\widehat{r}^{2}(d,\cdot)-r^{*}(d,\cdot)^{2}\right)({\mathrm{d}}(P_{0}-P_{n}))\right|$ $=O_{p}\left(\max\left\{\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}},\frac{\left\|\widehat{r}(d,\cdot)\right\|^{2}_{{\mathcal{H}}}}{n^{2/(2+\gamma)}}\right\}\right),$
+Under the conditions of Theorem 4.1, for any $0<\gamma<2$, we have
+$d\left({\mathbb{E}}-\widehat{{\mathbb{E}}}\right)\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
+$=O_{p}\left(\max\left\{\frac{\|\widehat{h}_{n}-h^{*}\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{h}_{n}\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}},\frac{\left(1+\left\|\widehat{h}_{n}\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right\}\right),$
 as $n\to\infty$.
@@
 Proof.
-From Proposition E.2 and $r_{0}\in{\mathcal{R}}^{\mathrm{RKHS}}$, we have
-$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}+\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}$
-$\leq\sum_{d\in\{1,0\}}\left|\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})\right|+\left|\int e_{0}(d,\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\right|+\lambda\|r_{0}\|^{2}_{{\mathcal{H}}}.$
+From Proposition E.3 and $h_{0}\in{\mathcal{H}}^{\mathrm{RKHS}}$, we have
+$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{h}\|^{2}_{{\mathcal{H}}}$
+$\leq\left({\mathbb{E}}-\widehat{{\mathbb{E}}}\right)\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]+\lambda\|f_{0}\|^{2}_{{\mathcal{H}}}.$
 From Proposition F.4, we have
-$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}+\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}$
-$=\sum_{d\in\{1,0\}}O_{p}\left(\max\left\{\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}},\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right\}\right)+\lambda\|r_{0}\|^{2}_{{\mathcal{H}}}.$
+$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{f}_{n}\|^{2}_{{\mathcal{H}}}$
+$=\O _{p}\left(\max\left\{\frac{\|\widehat{h}-h_{0}\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}},\frac{\left(1+\left\|\widehat{h}\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right\}\right)+\lambda\|r_{0}\|^{2}_{{\mathcal{H}}}.$
 We consider the following three possibilities:
-$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}+\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=O_{p}(\lambda),$ (4)
-$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}+\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right),$ (5)
-$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}+\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$ (6)
+$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{f}_{n}\|^{2}_{{\mathcal{H}}}=O_{p}(\lambda),$ (2)
+$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{f}_{n}\|^{2}_{{\mathcal{H}}}=O_{p}\left(\frac{\|\widehat{f}-f_{0}|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right),$ (3)
+$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{f}_{n}\|^{2}_{{\mathcal{H}}}=O_{p}\left(\frac{\left(1+\left\|\widehat{f}\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$ (4)
 The above inequalities are analyzed as follows:
@@
 We have
-$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}=O_{p}(\lambda),$
-$\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=O_{p}(\lambda).$
-Therefore, we have $\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|_{P_{0}}=O_{p}(\lambda^{1/2})$ and $\|\widehat{r}\|_{{\mathcal{H}}}=O_{p}(1)$.
+$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}=O_{p}(\lambda),$
+$\lambda\|\widehat{f}_{n}\|^{2}_{{\mathcal{H}}}=O_{p}(\lambda).$
+Therefore, we have $\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|_{P_{0}}=O_{p}(\lambda^{1/2})$ and $\|\widehat{r}\|_{{\mathcal{H}}}=O_{p}(1)$.
 Case (3).
 We have
-$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right),$
-$\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right).$
+$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}=O_{p}\left(\frac{\|\widehat{f}_{n}-f_{0})\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}_{n}\right\|_{{\mathcal{F}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right),$
+$\lambda\|\widehat{f}_{n}\|^{2}_{{\mathcal{H}}}=O_{p}\left(\frac{\|\widehat{f}_{n}-f_{0})\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}_{n}\right\|_{{\mathcal{F}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right).$
 From the first inequality, we have
-$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{n^{1/(2+\gamma)}}\right).$
+$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{f}_{n}\right\|_{{\mathcal{F}}}\right)^{1+\gamma/2}}{n^{1/(2+\gamma)}}\right).$
 By using this result, from the second inequality, we have
-$\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}$ $=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\|\widehat{r}(d,\cdot)-r^{*}(d,\cdot)\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right)$
-$=\sum_{d\in\{1,0\}}O_{p}\left(\left(\frac{1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}}{n^{1/(2+\gamma)}}\right)^{1-\gamma/2}\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right)$
-$=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$
+$\lambda\|\widehat{f}_{n}\|^{2}_{{\mathcal{H}}}$ $=O_{p}\left(\frac{\|\widehat{f}_{n}-f_{0})\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}_{n}\right\|_{{\mathcal{F}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right)$
+$=O_{p}\left(\left(\frac{1+\left\|\widehat{f}_{n}\right\|_{{\mathcal{F}}}}{n^{1/(2+\gamma)}}\right)^{1-\gamma/2}\frac{\left(1+\left\|\widehat{f}_{n}\right\|_{{\mathcal{F}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right)$
+$=O_{p}\left(\frac{\left(1+\left\|\widehat{f}_{n}\right\|_{{\mathcal{F}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$
 This implies that
-$\|\widehat{r}\|_{{\mathcal{H}}}$ $=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{\lambda^{1/2}n^{2/(2+\gamma)}}\right)=o_{p}(1).$
+$\|\widehat{f}\|_{{\mathcal{H}}}$ $=O_{p}\left(\frac{\left(1+\left\|\widehat{f}_{n}\right\|_{{\mathcal{F}}}\right)^{2}}{\lambda^{1/2}n^{2/(2+\gamma)}}\right)=o_{p}(1).$
 Therefore, the following inequity is obtained.
-$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{1}{n^{1/(2+\gamma)}}\right)=O_{p}(\lambda^{1/2}).$
-Case 6.
+$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|_{P_{0}}=O_{p}\left(\frac{1}{n^{1/(2+\gamma)}}\right)=O_{p}(\lambda^{1/2}).$
+Case 4.
 We have
-$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right),$
-$\lambda\|\widehat{r}\|^{2}_{{\mathcal{H}}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{r}(d,\cdot)\right\|_{{\mathcal{H}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$
+$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}=O_{p}\left(\frac{\left(1+\left\|\widehat{f}_{n}\right\|_{{\mathcal{F}}}\right)^{2}}{n^{2/(2+\gamma)}}\right),$
+$\lambda\|\widehat{f}_{n}\|^{2}_{{\mathcal{H}}}=O_{p}\left(\frac{\left(1+\left\|\widehat{f}_{n}\right\|_{{\mathcal{F}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$
 As well as the argument in (3), we have $\|\widehat{r}\|_{{\mathcal{H}}}=o_{p}(1)$. Therefore, we have
-$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{1}{n^{1/(2+\gamma)}}\right)=O_{p}(\lambda^{1/2}).$
+$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|_{P_{0}}=O_{p}\left(\frac{1}{n^{1/(2+\gamma)}}\right)=O_{p}(\lambda^{1/2}).$
 Appendix G Proof of Theorem 4.2
@@
 To bound (1), we show several auxiliary results. Define
-$\widehat{{\mathcal{F}}}^{f^{*},u}$ $\coloneqq\{f\in{\mathcal{F}}^{\mathrm{FNN}}\colon\frac{1}{n}\sum^{n}_{i=1}(f(X_{i})-f^{*}(X_{i}))^{2}\leq u\},$
+$\widehat{{\mathcal{F}}}^{f^{*},u}$ $\coloneqq\{f\in{\mathcal{F}}^{\mathrm{FNN}}\colon\frac{1}{n}\sum^{n}_{i=1}(f(D_{i},X_{i})-f^{*}(D_{i},X_{i}))^{2}\leq u\},$
 $\overline{{\mathcal{G}}}^{f^{*},u}$ $\coloneqq\Big\{(f-f^{*})\colon f\in\widehat{{\mathcal{F}}}^{f^{*},u}\Big\},$
@@
 Suppose that the conditions in Lemma G.1 hold. Then, for any $z>0$, with probability $1-\exp(-z)$ it holds that
-$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,x)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{n}+\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}\Bigg)$
-$\leq C\left(\|f^{*}(X)-f_{0}(D,X)\|^{2}_{2}+\|f^{*}(X)-f_{0}(X)\|_{2}\sqrt{\frac{z}{n}}+\frac{16Mz}{3n}\right).$
+$\widehat{{\mathbb{E}}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
+$\leq C\left(\|f^{*}(D,X)-f_{0}(D,X)\|^{2}_{2}+\|f^{*}(D,X)-f_{0}(D,X)\|_{2}\sqrt{\frac{z}{n}}+\frac{16Mz}{3n}\right).$
 Lemma G.3 (Corresponding to (29) in Zheng et al. (2022)).
 Suppose that the conditions in Lemma G.1 hold. If there exists $u_{0}>0$ such that
-$\|\widehat{f}(X)-f^{*}(X)\|_{2}\leq u_{0},$
+$\|\widehat{f}(D,X)-f^{*}(D,X)\|_{2}\leq u_{0},$
 then it holds that
-$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})+\int e_{0}(d\mid\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\Bigg)$
-$\leq C\left({\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},r}\right]+u_{0}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right).$
+$\left({\mathbb{E}}-\widehat{{\mathbb{E}}}\right)\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
+$\leq C\left({\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},u_{0}}\right]+u_{0}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right).$
 Additionally, we use the following three propositions directly from Zheng et al. (2022).
@@
 Proposition G.6 (Upper bound of the Rademacher complexity).
-Suppose that the conditions in Lemma G.1 hold. If $n\geq\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})$, $r_{0}\geq 1/n$, and $n\geq(2eM)^{2}$, we have
-${\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},r}\right]\leq Cr_{0}\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log n}{n}}.$
+Suppose that the conditions in Lemma G.1 hold. If $n\geq\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})$, $u_{0}\geq 1/n$, and $n\geq(2eM)^{2}$, we have
+${\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},u_{0}}\right]\leq Cr_{0}\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log n}{n}}.$
 Then, we prove Lemma G.1 as follows:
@@
 then from (1) and Lemmas G.2 and G.3, for every $z>0$, there exists a constant $C>0$ independent $n$ such that
-$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}$
+$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}$
 $\leq C\left(\|f^{*}-f_{0}\|_{2}\sqrt{\frac{z}{n}}+\frac{16Mz}{3n}+u_{0}\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log n}{n}}+u_{0}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right).$ (5)
 This result implies that if $\sqrt{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})}$, then there exists $n_{0}$ such that for all $n>n_{0}$, there exists $u_{1}<u_{0}$ such that
-$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}\leq u_{1}.$
+$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}\leq u_{1}.$
 For any $z>0$, define $\overline{u}$ as
@@
 If there exists $j\leq\ell$ such that $\widehat{f}\in\overline{{\mathcal{S}}}^{\mathrm{FNN}}_{j}$, then from (G.1), for every $z>0$, with probability at least $1-8\exp(-z)$, there exists a constant $C>0$ independent of $n$ such that
-$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{2}$
+$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{2}$
 $\leq C\left(2^{\ell-1}\overline{u}\left(\sqrt{\frac{\mathrm{Pdim}({\mathcal{F}}^{\mathrm{FNN}})\log(n)}{n}}+\sqrt{\frac{z}{n}}\right)+\|f^{*}-f_{0}\|^{2}_{2}+\|f^{*}-f_{0}\|_{2}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right).$ (6)
@@
 hold, then
-$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|_{2}\leq 2^{j-1}\overline{u}.$ (11)
+$\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|_{2}\leq 2^{j-1}\overline{u}.$ (9)
 Here, to obtain (9), we used $\overline{u}\geq\max\left\{\sqrt{\log(n)/n},4\sqrt{3}M\sqrt{z/n},u^{\dagger}\right\}$, (G.1), (7), and (8).
@@
 From Proposition D.1, we have
-$\int-2\Big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\Big){\mathrm{d}}P_{n}$
-$\ \ \ \ \ \ \leq\int-2\Big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\Big){\mathrm{d}}P_{0}+\sqrt{2}C_{1}\|f^{*}(X)-f_{0}(X)\|\sqrt{\frac{z}{n}}+\frac{16C_{1}Mz}{3n},$
-$\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}$
-$\ \ \ \ \ \ \leq\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{0}+\sqrt{2}C_{2}\|f^{*}(X)-f_{0}(X)\|\sqrt{\frac{z}{n}}+\frac{16C_{2}Mz}{3n}.$
+$\widehat{{\mathbb{E}}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
+$\leq{\mathbb{E}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]+\sqrt{2}C\|f^{*}(X)-f_{0}(X)\|\sqrt{\frac{z}{n}}+\frac{16C_{1}Mz}{3n}.$
 This is a direct consequence of Proposition D.1.
-Note that $r^{*}$ and $r_{0}$ are fixed, and it is enough to apply the standard law of large numbers; that is, we do not have to consider the uniform law of large numbers.
+Note that $h^{*}$ and $h_{0}$ are fixed, and it is enough to apply the standard law of large numbers; that is, we do not have to consider the uniform law of large numbers.
 However, we can still apply Proposition D.1, which is a general than the standard law of large numbers, with ignoring the Rademacher complexity part.
-Here, we also used
-$\Big|r^{*}(d,x)-r_{0}(d,x)\Big|\leq\frac{1}{4}\Big|f^{*}(x)-f_{0}(x)\Big|,$
-$\Big|r^{*}(d,x)^{2}-r_{0}(d,x)^{2}\Big|\leq\frac{1}{2}\Big|f^{*}(x)-f_{0}(x)\Big|.$
 We have
-$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{n}+\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}\Bigg)$
-$\leq\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{0}+\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{0}$
-$\ \ \ \ \ +\sqrt{2}C_{1}\|f^{*}-f_{0}\|\sqrt{\frac{z}{n}}+\frac{16C_{2}Mz}{3n}+\sqrt{2}C_{2}\|f^{*}-f_{0}\|\sqrt{\frac{z}{n}}+\frac{16C_{2}Mz}{3n}\Bigg)$
+$\widehat{{\mathbb{E}}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
+$\leq{\mathbb{E}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
+$\ \ \ \ \ +\sqrt{2}C_{1}\|f^{*}-f_{0}\|\sqrt{\frac{z}{n}}+\frac{16C_{2}Mz}{3n}+\sqrt{2}C_{2}\|f^{*}-f_{0}\|\sqrt{\frac{z}{n}}+\frac{16C_{2}Mz}{3n}$
 $\leq C\left(\|f^{*}-f_{0}\|^{2}_{2}+\|f^{*}-f_{0}\|\sqrt{\frac{z}{n}}+\frac{16CMz}{3n}\right).$
@@
 $g\leq 4M^{2}$
-Additionally, we assumed that $\|\widehat{f}-f^{*}\|_{2}\leq r_{0}$ holds. Then, it holds that $\mathrm{Var}_{P_{0}}(g)\leq 4M^{2}r^{2}$.
+Additionally, we assumed that $\|\widehat{f}-f^{*}\|_{2}\leq u_{0}$ holds. Then, it holds that $\mathrm{Var}_{P_{0}}(g)\leq 4M^{2}u^{2}_{0}$.
 Here, we note that the followings hold for all $f$ ($r$):
-$\Big|r(d,x)-r^{*}(d,x)\Big|\leq\frac{1}{4}\Big|f(x)-f^{*}(x)\Big|,$
-$\Big|r(d,x)^{2}-r^{*}(d,x)^{2}\Big|\leq\frac{1}{2}\Big|f(x)-f^{*}(x)\Big|.$
-Then, from Proposition D.1, for every $z>0$, with probability at least $1-\exp(-z)$, it holds that, for each $d\in\{1,0\}$ it holds that Then, it holds that
-$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})+\int e_{0}(d\mid\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\Bigg)$
-$\leq C\left({\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},r}\right]+r_{0}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right).$
+$L(h)-L(h^{*})\leq C\Big|f(d,x)-f^{*}(d,x)\Big|,$
+where $C>0$ is some constant
+Then, from Proposition D.1, for every $z>0$, with probability at least $1-\exp(-z)$, it holds that
+$\left({\mathbb{E}}-\widehat{{\mathbb{E}}}\right)\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
+$\leq C\left({\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},u_{0}}\right]+r_{0}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right).$
+Appendix H Nearest neighbor matching
+In this section, we show that nearest neighbor (NN) matching for the ATE can be interpreted as a special case of our direct bias-correction term estimation with the squared loss, that is, Riesz regression or LSIF.
+This result is shown in Kato (2025b), a subsequent work of this study.
+The key step is to express the ATE bias-correction term $h_{0}(D,X)$ in terms of density ratios with respect to the marginal covariate distribution and then to approximate these density ratios via nearest neighbor cells, following the density-ratio interpretation in Lin et al.
+(2023).
+H.1 ATE bias-correction term and density ratios
+Let $p_{X}$ denote the marginal density of $X$ and $p_{X\mid D=d}$ the conditional density of $X$ given $D=d$.
+Let $\pi_{1}\coloneqq P_{0}(D=1)$ and $\pi_{0}\coloneqq P_{0}(D=0)=1-\pi_{1}$.
+By Bayes’ rule,
+$p_{X\mid D=d}(x)=\frac{p_{X}(x)P_{0}(D=d\mid X=x)}{P_{0}(D=d)}=\frac{p_{X}(x)e_{0}(x)^{d}(1-e_{0}(x))^{1-d}}{\pi_{d}},$
+where $\pi_{d}=P_{0}(D=d)$ and $e_{0}(x)=P_{0}(D=1\mid X=x)$.
+Define the density ratios with respect to the marginal distribution of $X$ by
+$r_{1}(x)\coloneqq\frac{p_{X}(x)}{p_{X\mid D=1}(x)},\qquad r_{0}(x)\coloneqq\frac{p_{X}(x)}{p_{X\mid D=0}(x)}.$
+From the expression above,
+$r_{1}(x)=\frac{\pi_{1}}{e_{0}(x)},\qquad r_{0}(x)=\frac{\pi_{0}}{1-e_{0}(x)}.$
+Therefore, the ATE bias-correction term
+$h_{0}(D,X)=\frac{\mathbbm{1}[D=1]}{e_{0}(X)}-\frac{\mathbbm{1}[D=0]}{1-e_{0}(X)}$
+can be written in terms of $r_{1}$ and $r_{0}$ as
+$h_{0}(D,X)$ $=\mathbbm{1}[D=1]\frac{r_{1}(X)}{\pi_{1}}-\mathbbm{1}[D=0]\frac{r_{0}(X)}{\pi_{0}}.$ (10)
+Thus, estimating $h_{0}$ is equivalent to estimating the pair $(r_{1},r_{0})$, the density ratios between the marginal covariate distribution and the treated and control covariate distributions.
+H.2 Squared loss objective and decomposition into two LSIF problems
+Recall that when we choose the squared loss $g^{\mathrm{SL}}(h)=(h-1)^{2}$, the population Bregman divergence objective for $h$ is
+$\mathrm{BR}_{g^{\mathrm{SL}}}(h)={\mathbb{E}}\Big[-2\big(h(1,X)-h(0,X)\big)+h(D,X)^{2}\Big].$
+Consider the parameterization
+$h(D,X)=\mathbbm{1}[D=1]\frac{r_{1}(X)}{\pi_{1}}-\mathbbm{1}[D=0]\frac{r_{0}(X)}{\pi_{0}},$
+with $r_{1},r_{0}$ defined above. Substituting this into $\mathrm{BR}_{g^{\mathrm{SL}}}(h)$ and using the law of total expectation, we obtain
+$\mathrm{BR}_{g^{\mathrm{SL}}}(h)$ $=C-2{\mathbb{E}}\Big[\frac{r_{1}(X)}{\pi_{1}}+\frac{r_{0}(X)}{\pi_{0}}\Big]+{\mathbb{E}}\Big[h(D,X)^{2}\Big],$ (11)
+where $C$ is a constant independent of $(r_{1},r_{0})$. The last term can be decomposed as
+${\mathbb{E}}\big[h(D,X)^{2}\big]=\pi_{1}{\mathbb{E}}\Big[\bigg(\frac{r_{1}(X)}{\pi_{1}}\bigg)^{2}\Bigm|D=1\Big]+\pi_{0}{\mathbb{E}}\Big[\bigg(\frac{r_{0}(X)}{\pi_{0}}\bigg)^{2}\Bigm|D=0\Big].$
+Rewriting (11) in terms of expectations with respect to $p_{X}$ and $p_{X\mid D=d}$ and dropping constants gives
+$\mathrm{BR}_{g^{\mathrm{SL}}}(h)$ $\coloneqq-2{\mathbb{E}}_{X}\left[r_{1}(X)\right]+{\mathbb{E}}_{X\mid D=1}\left[r_{1}(X)^{2}\right]-2{\mathbb{E}}_{X}\left[r_{0}(X)\right]+{\mathbb{E}}_{X\mid D=0}\left[r_{0}(X)^{2}\right].$ (12)
+Hence minimizing $\mathrm{BR}_{g^{\mathrm{SL}}}(h)$ over $(r_{1},r_{0})$ is equivalent to solving two independent LSIF-type problems
+$r_{1}^{*}$ $=\operatorname*{arg\,min}_{r_{1}}\left\{-2{\mathbb{E}}_{X}[r_{1}(X)]+{\mathbb{E}}_{X\mid D=1}[r_{1}(X)^{2}]\right\},$
+$r_{0}^{*}$ $=\operatorname*{arg\,min}_{r_{0}}\left\{-2{\mathbb{E}}_{X}[r_{0}(X)]+{\mathbb{E}}_{X\mid D=0}[r_{0}(X)^{2}]\right\},$
+and then plugging $(r_{1}^{*},r_{0}^{*})$ into (10).
+At the sample level, with ${\mathcal{G}}_{1}$ and ${\mathcal{G}}_{0}$ defined as in the Introduction, the empirical LSIF objectives are
+$\widehat{J}_{1}(r_{1})$ $\coloneqq-\frac{2}{n}\sum_{i=1}^{n}r_{1}(X_{i})+\frac{1}{|{\mathcal{G}}_{1}|}\sum_{i\in{\mathcal{G}}_{1}}r_{1}(X_{i})^{2},$ (13)
+$\widehat{J}_{0}(r_{0})$ $\coloneqq-\frac{2}{n}\sum_{i=1}^{n}r_{0}(X_{i})+\frac{1}{|{\mathcal{G}}_{0}|}\sum_{i\in{\mathcal{G}}_{0}}r_{0}(X_{i})^{2}.$ (14)
+Minimizing $\widehat{J}_{1}$ and $\widehat{J}_{0}$ and then using (10) yields an LSIF (Riesz regression) estimator of the ATE bias-correction term $h_{0}$.
+H.3 Nearest-neighbor partition and histogram model
+To connect this LSIF formulation to nearest neighbor matching, we now choose a simple histogram-type model for $(r_{1},r_{0})$ based on nearest neighbor cells.
+Let us consider the $M$-nearest neighbor partition induced by the sample $\{X_{i}\}_{i=1}^{n}$.
+For each treated unit $i\in{\mathcal{G}}_{1}$, let $N_{M}^{(0)}(i)\subset{\mathcal{G}}_{0}$ denote the set of $M$ nearest control units to $X_{i}$.
+Similarly, for each control unit $j\in{\mathcal{G}}_{0}$, let $N_{M}^{(1)}(j)\subset{\mathcal{G}}_{1}$ denote the set of $M$ nearest treated units to $X_{j}$.
+We define the neighbor counts
+$K_{M}^{(1)}(k)\coloneqq\big|\{i\in{\mathcal{G}}_{1}\colon k\in N_{M}^{(0)}(i)\}\big|,\qquad K_{M}^{(0)}(k)\coloneqq\big|\{j\in{\mathcal{G}}_{0}\colon k\in N_{M}^{(1)}(j)\}\big|.$
+Thus $K_{M}^{(1)}(k)$ counts how often unit $k$ is selected as a control neighbor of treated units, and $K_{M}^{(0)}(k)$ counts how often it is selected as a treated neighbor of control units.
+The total numbers of neighbor links are
+$\sum_{k=1}^{n}K_{M}^{(1)}(k)=M|{\mathcal{G}}_{1}|,\qquad\sum_{k=1}^{n}K_{M}^{(0)}(k)=M|{\mathcal{G}}_{0}|.$
+We now approximate each density ratio $r_{d}$ by a histogram that is constant on the Voronoi cells induced by the sample:
+$r_{d}(x)=\sum_{k=1}^{n}\theta^{(d)}_{k}\psi_{k}(x),$
+where $\{\psi_{k}\}_{k=1}^{n}$ is the partition of $\mathcal{X}$} such that $\psi_{k}(x)=1$ if $x$ lies in the cell associated with $X_{k}$ and $\psi_{k}(x)=0$ otherwise.
+Approximating the integrals in (13) and (14) by assigning each observation $X_{i}$ to the nearest cell, the empirical objectives become (up to constants)
+$\widehat{J}_{1}(\theta^{(1)})$ $\approx-\frac{2}{n}\sum_{k=1}^{n}K_{M}^{(X)}(k)\,\theta^{(1)}_{k}+\frac{1}{|{\mathcal{G}}_{1}|}\sum_{k=1}^{n}\mathbbm{1}[k\in{\mathcal{G}}_{1}]\big(\theta^{(1)}_{k}\big)^{2},$ (15)
+$\widehat{J}_{0}(\theta^{(0)})$ $\approx-\frac{2}{n}\sum_{k=1}^{n}K_{M}^{(X)}(k)\,\theta^{(0)}_{k}+\frac{1}{|{\mathcal{G}}_{0}|}\sum_{k=1}^{n}\mathbbm{1}[k\in{\mathcal{G}}_{0}]\big(\theta^{(0)}_{k}\big)^{2},$ (16)
+where $K_{M}^{(X)}(k)$ denotes the number of times $X_{k}$ is selected as a nearest neighbor when we run the $M$-NN search over the whole sample $\{X_{i}\}_{i=1}^{n}$.
+[^2] For a detailed derivation of this approximation, see the analysis of histogram LSIF in Lin et al.
+(2023).
+Minimizing the quadratic objectives (15) and (16) with respect to each $\theta^{(d)}_{k}$ yields the closed-form solutions
+$\theta^{(1)*}_{k}\propto K_{M}^{(X)}(k)\mathbbm{1}[k\in{\mathcal{G}}_{1}],\qquad\theta^{(0)*}_{k}\propto K_{M}^{(X)}(k)\mathbbm{1}[k\in{\mathcal{G}}_{0}].$
+Therefore, up to a common normalization constant,
+$r_{1}(X_{k})\propto K_{M}^{(X)}(k)\mathbbm{1}[k\in{\mathcal{G}}_{1}],\qquad r_{0}(X_{k})\propto K_{M}^{(X)}(k)\mathbbm{1}[k\in{\mathcal{G}}_{0}].$
+Substituting these expressions into (10) gives
+$h^{\mathrm{NN}}(D_{k},X_{k})$ $=(2D_{k}-1)\Big(1+\frac{K_{M}^{(X)}(k)}{M}\Big)\times c_{n},$ (17)
+for some sample-size dependent normalization constant $c_{n}$.
+Equation (17) coincides, up to normalization, with the nearest-neighbor based bias-correction weights derived in Lin et al.
+(2023) for the ATE.
+H.4 Nearest neighbor matching as Riesz regression
+Using the bias-correction term $h^{\mathrm{NN}}$ in (17), the corresponding IPW-type ATE estimator becomes
+$\widehat{\tau}^{\mathrm{NN}}=\frac{1}{n}\sum_{k=1}^{n}h^{\mathrm{NN}}(D_{k},X_{k})Y_{k},$
+which can be expanded to the familiar $M$-nearest neighbor matching form
+$\widehat{\tau}^{\mathrm{NN}}=\frac{1}{n}\sum_{i\in{\mathcal{G}}_{1}}\Big(Y_{i}-\frac{1}{M}\sum_{j\in N_{M}^{(0)}(i)}Y_{j}\Big)-\frac{1}{n}\sum_{j\in{\mathcal{G}}_{0}}\Big(Y_{j}-\frac{1}{M}\sum_{i\in N_{M}^{(1)}(j)}Y_{i}\Big),$
+that is, a two-sided nearest neighbor matching estimator for the ATE that matches treated units to control units and control units to treated units.
+Therefore, nearest neighbor matching for the ATE is obtained by minimizing the squared-loss Bregman divergence within a nearest-neighbor histogram model for the density ratios $(r_{1},r_{0})$ and then plugging the resulting estimator into the bias-correction term $h(D,X)$.
+In other words, nearest neighbor matching is a special case of Riesz regression (LSIF) with a particular choice of feature dictionary based on nearest neighbor cells.
+This formally justifies the statement in the main text that nearest neighbor matching can be interpreted as a direct bias-correction term estimator obtained from our squared-loss Bregman divergence framework.
+Table 3: Results of additional simulation studies.
+CR denotes the coverage ratio of 9595% confidence intervals; that is, values close to $0.95$ are better.
+DM denotes the direct method, which is independent of the direct bias-correction term estimation methods; therefore, in theory, the results of the DM estimator should not differ across DBC (LS), DBC (KL), and DBC (TL).
+Since we compute the DM estimator when constructing the AIPW estimator in each of DBC (LS), DBC (KL), and DBC (TL), we also report the DM estimator results for reference.
+True DBC (LS) DBC (KL) DBC (TL)
+DM IPW AIPW DM IPW AIPW DM IPW AIPW DM IPW AIPW
+MSE 0.00 1.10 0.01 0.30 0.59 0.11 0.30 0.41 0.08 0.31 0.36 0.09
+CR 1.00 0.92 0.97 0.17 0.97 0.87 0.11 0.97 0.88 0.11 0.92 0.87
+Appendix I Additional simulation studies
+In this section, we conduct additional simulation studies to more closely examine the finite sample behavior of our direct bias-correction approach under different choices of Bregman divergence.
+We focus on the three representative losses introduced in Section 2: the squared loss corresponding to Riesz regression (denoted by DBC (LS)), the KL divergence loss (DBC (KL)), and the tailored loss (DBC (TL)).
+We refer to our method collectively as the direct bias-correction (DBC) approach.
+Unlike the simulation design in Section 2 (Simulation studies), here we explicitly use cross fitting in the sense of Assumption 5.1.
+This setting illustrates how our framework can be combined with modern high-capacity models without requiring the Donsker assumption.
+I.1 Design and implementation
+We consider the same basic ATE setting as in the previous simulations.
+The covariates are three dimensional, $K=3$, and we fix the sample size at $n=3000$.
+In each Monte Carlo replication, we generate covariates $X_{i}\in{\mathbb{R}}^{3}$ from a multivariate normal distribution $\mathcal{N}(0,I_{3})$, and construct a nonlinear propensity score model with polynomial and interaction terms, as in the main simulation study.
+Treatment assignments $D_{i}$ are then sampled from the resulting Bernoulli distribution with success probability $e_{0}(X_{i})$.
+The outcome $Y_{i}$ is generated from a nonlinear regression model that includes both squared terms and a nonlinear transformation, with the true ATE fixed at $\tau_{0}=5.0$.
+The noise term is standard normal.
+This design yields a moderately complex but smooth data generating process for both the propensity score and the conditional outcome.
+To evaluate the efficiency and coverage properties of the estimators, we construct an oracle benchmark that uses the true nuisance functions.
+For each replication, we compute the infeasible DM, IPW, and AIPW estimators based on the true propensity score and the true conditional expectations of $Y(d)$, and we use their corresponding influence functions to form oracle 9595% confidence intervals.
+The performance of these oracle estimators is summarized in the “True” columns of Table 3.
+For our proposed DBC estimators, we estimate the bias-correction term $h_{0}(D,X)$ using one hidden layer neural networks.
+In all cases, we use fully connected networks with a single hidden layer of $100$ nodes.
+For DBC (LS), we employ the squared loss objective associated with Riesz regression.
+For DBC (KL) and DBC (TL), we use the KL divergence loss and the tailored loss introduced in Section 2.6, respectively.
+The conditional outcome regression $\mu_{0}(d,X)$ for the DM and AIPW estimators is also modeled by a neural network with one hidden layer and $100$ nodes.
+In DBC (LS), we model $h_{0}$ directly using a neural network with one hidden layer consisting of 100 nodes.
+In DBC (KL), DBC (TL), and MLE, we model $h_{0}$ by estimating the propensity score using a neural network with one hidden layer consisting of 100 nodes.
+To avoid relying on the Donsker condition, all nuisance functions (the bias-correction term and the outcome regression) are estimated with two-fold cross fitting.
+Specifically, in each replication, we split the sample into two folds, estimate the nuisance functions on one fold, evaluate the corresponding scores on the other fold, and then swap the roles of the folds.
+The final estimators are obtained by aggregating the two cross-fitted folds.
+For each loss (LS, KL, TL), we report three estimators:
+  • the direct method (DM), which depends only on the outcome regression;
+  • the IPW estimator, constructed using the estimated bias-correction term;
+  • the AIPW estimator, which combines both the estimated bias-correction term and the outcome regression.
+Note that the DM estimator is theoretically independent of the specific loss used to estimate the bias-correction term.
+In practice, we recompute the DM estimator within each DBC (LS), DBC (KL), and DBC (TL) run to construct the AIPW estimator, and we report the resulting DM performance for reference.
+Small differences among the DM columns therefore reflect only Monte Carlo variation.
+We repeat the experiment $100$ times.
+For each method and each estimator (DM, IPW, AIPW), we compute the empirical mean squared error (MSE) of the ATE estimate and the empirical coverage ratio (CR) of the nominal 9595% confidence interval, defined as the fraction of replications in which the interval contains the true effect $\tau_{0}$.
+The results are summarized in Table 3.
+I.2 Results
+Table 3 reports the MSE and coverage ratio for the oracle estimators (True) and for the three DBC variants.
+The oracle AIPW estimator achieves a very small MSE (approximately $0.01$) and a coverage ratio close to the nominal level ($0.97$), as expected.
+The oracle IPW estimator exhibits a larger MSE (around $1.10$) and slightly conservative coverage ($0.92$).
+The oracle DM estimator is unbiased by construction, hence its MSE is essentially zero and its coverage ratio is close to one.
+For the feasible DBC estimators, the DM columns are nearly identical across DBC (LS), DBC (KL), and DBC (TL), with MSE around $0.30$ and poor coverage (CR between $0.11$ and $0.17$).
+This behavior reflects the well known fact that the plug in DM estimator is not debiased and is not suitable for inference in this design, even when the outcome model is reasonably flexible.
+The IPW estimators based on our direct bias-correction term exhibit substantially reduced MSE relative to the oracle IPW benchmark that uses the true propensity score.
+Sucha a “paradox” is reporeted and analyzed in existing studies, such as Hirano et al.
+(2003) and Henmi & Eguchi (2004).
+Under DBC (LS), the IPW MSE is about $0.59$, while DBC (KL) and DBC (TL) further reduce it to approximately $0.41$ and $0.36$, respectively.
+The coverage ratios for IPW are close to the nominal level for all three losses (around $0.97$ for DBC (LS) and DBC (KL), and $0.92$ for DBC (TL)).
+These results indicate that direct estimation of the bias-correction term can improve both efficiency and coverage for IPW, and that the KL and tailored losses provide modest gains over the squared loss in this setting.
+The AIPW estimators exhibit the best overall performance.
+All three DBC variants achieve small MSEs, with values around $0.11$ for DBC (LS), $0.08$ for DBC (KL), and $0.09$ for DBC (TL), which are close to the oracle AIPW MSE of $0.01$.
+The coverage ratios of the AIPW estimators are slightly below the nominal level (between $0.87$ and $0.88$) but still reasonably close, especially given the moderate number of Monte Carlo replications.
+The differences among the three losses are minor, with DBC (KL) and DBC (TL) showing a slight advantage in terms of MSE.
+Overall, these additional experiments support our theoretical findings.
+First, they confirm that direct estimation of the bias-correction term via Bregman divergence minimization yields ATE estimators that are close to the oracle benchmark when combined with cross fitting.
+Second, they show that the choice of Bregman divergence (squared loss, KL loss, or tailored loss) has only a modest impact on the performance of the AIPW estimator, while the KL and tailored losses can provide small efficiency gains in some cases.
+Third, they illustrate that our framework can be implemented with flexible neural network models and cross fitting, without relying on the Donsker condition.
+Table 4: MSE and coverage ratio (CR) of ATE estimators in the semi-synthetic IHDP experiment.
+We report the mean squared error (MSE) and the empirical coverage ratio (CR) of nominal 9595% confidence intervals over 1000 replications for the direct method (DM), inverse probability weighting (IPW), and augmented IPW (AIPW) estimators.
+Nuisance functions are estimated either by a neural network with one hidden layer of size 100 or by an RKHS regression with 100 Gaussian basis functions.
+The columns correspond to different variants of the direct bias-correction (DBC) approach based on least squares (LS), Kullback–Leibler (KL), truncated likelihood (TL), and maximum likelihood (MLE) criteria.
+Neural network RKHS
+DBC (LS) DBC (LS) DBC (TL) DBC (MLE) DBC (LS) DBC (LS) DBC (TL) DBC (MLE)
+DM IPW AIPW DM IPW AIPW DM IPW AIPW DM IPW AIPW DM IPW AIPW DM IPW AIPW DM IPW AIPW DM IPW AIPW
+MSE 1.52 6.82 0.31 1.57 9.42 0.44 1.55 2.84 0.32 1.58 3.00 0.43 19.98 3.56 19.97 3.50 1.91 4.58 2.59 1.78 4.45 2.48 1.22 2.32
+CR 0.03 0.41 1.00 0.06 0.08 1.00 0.03 0.73 0.94 0.01 0.61 0.90 0.00 0.00 0.00 0.34 0.91 0.82 0.48 0.93 0.88 0.39 0.81 0.84
+Appendix J Experiments with semi-synthetic datasets
+We next evaluate the proposed estimators on a semi-synthetic benchmark based on the Infant Health and Development Program (IHDP) data, following Chernozhukov et al.
+(2022a).
+The IHDP was a randomized trial that investigated the effect of an early childhood intervention on subsequent developmental and health outcomes.
+Following the standard setting “A” implemented in the npci package, we generate 1000 semi-synthetic datasets, each consisting of $n=747$ observations with a binary treatment $T$, an outcome $Y$, and $p=25$ continuous and binary covariates $X$.
+The estimand of interest is the average treatment effect (ATE) of the intervention on $Y$.
+For each semi-synthetic dataset we compute three ATE estimators: the direct method (DM), the inverse probability weighting (IPW) estimator, and the augmented IPW (AIPW) estimator.
+All estimators use our direct bias-correction (DBC) approach for estimating the Riesz representer or density ratio.
+We consider several variants of DBC based on different divergence criteria, including least squares (LS), Kullback–Leibler (KL), truncated likelihood (TL), and maximum likelihood (MLE).
+The nuisance functions are estimated either by a feedforward neural network or by a reproducing kernel Hilbert space (RKHS) regression.
+The neural network has a single hidden layer with 100 units and is trained for 100 epochs.
+For the RKHS learner we use 100 Gaussian basis functions; the bandwidth of the Gaussian kernel as well as the ridge regularization parameter are chosen by cross validation.
+To assess estimation accuracy and uncertainty quantification, we report the mean squared error (MSE) of each ATE estimator and the empirical coverage ratio (CR) of nominal 9595% Wald-type confidence intervals across the 1000 replications.
+Here, CR is defined as the proportion of replications in which the confidence interval contains the true ATE, so values close to $0.95$ indicate well calibrated intervals.
+The results are summarized in Table 4.
+Overall, when neural networks are used for nuisance estimation, the AIPW estimator combined with our DBC schemes achieves substantially smaller MSE than the corresponding DM and IPW estimators, while its CR is close to one, indicating slightly conservative but reliable inference.
+The DM estimator exhibits noticeable bias and severe undercoverage, and the IPW estimator can be unstable, especially for some DBC variants.
+When RKHS learners are employed, the IPW estimator performs relatively well in terms of both MSE and CR, whereas the DM and AIPW estimators are more sensitive to the choice of DBC method and can suffer from larger MSE or poor coverage.
+These findings suggest that, in this IHDP benchmark, DBC-based AIPW with neural network nuisance learners provides the most accurate and well calibrated ATE estimates.